A Complete Bibliography of the Journal of Econometrics (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

(p, d, q), d = 0 [De 80]. 1 [BR85, De 80]. AR(1) [Tho87]. χ^{2} [BJY84]. F [BJY84, UCS84]. k [DS84, SS83]. M [Dun87, PP86]. R^{2} [Cra87]. t [BR85, RT80b].

* [AG84, PMWA83].

-class [DS84, SS83]. -estimation [Dun87]. -estimators [PP86]. -ratios [UCS84].

1 [Ano81g]. 17 [Ano83c]. 1980 [Ano80w, Ano80x, Ano80y, Ano80t, Ano80q, Ano80u, Ano80r, Ano80v, Ano80s, Ano81-30, Ano80g]. 1980-3 [Ano80g]. 1981 [Ano81-29, Ano81-28, Ano81-27, Ano81z, Ano81w, Ano81y, Ano81v, Ano81x, Ano81u, Ano81g]. 1981-1 [Ano81g]. 1982 [Ano81c, Ano81s, Ano81-31, Ano82p, Ano82n, Ano82r, Ano82q, Ano82o, Ano82l, Ano82m, Ano82k]. 1983 [Ano81b, Ano83i, Ano83s, Ano83r, Ano83t, Ano83p, Ano83n, Ano83o, Ano83q]. 1984 [Ano84m, Ano84l, Ano84h, Ano84k, Ano84i, Ano84j]. 1985 [Ano85q, Ano85y, Ano85t, Ano85r, Ano85p, Ano85w, Ano85x, Ano85u, Ano85s, Ano85v]. 1986 [Ano86s, Ano86t, Ano86u, Ano86n, Ano86o, Ano86r, Ano86q, Ano86p]. 1987 [Ano87o, Ano87n, Ano87q, Ano87l, Ano87p, Ano87m]. 1988 [Ano88r, Ano88o, Ano88t, Ano88u, Ano88q, Ano88s, Ano88p]. 1989 [Ano89s, Ano89p, Ano89r, Ano89t, Ano89v, Ano89o, Ano89u, Ano89q, Ano89n].

252 [Ano83c]. 2SLS [Kni85].

3 [Ano80g]. 37 [Xie11]. 3rd [Ano81c].

40 [Kra13].

8th [Ano81c].

absence [AL84]. absolute [Pow84]. accelerated [Lee84]. acceleration [Gew88]. accuracy [LY89, Pal81]. Acknowledgement [Ano80a, Ano80b, Ano81d, Ano82a, Ano83a, Ano85a, Ano85b, Ano85c, Ano86b, Ano87a, Ano88a]. across [HS89b]. adaptation [HT84]. Adaptive [New88, PP86]. adjustment [Cec86, HS81a, LWD87, Mar80, May89, MO86, Pie80, Sto80]. adjustments [BF86b, GP83]. adoption [DR89]. affect [BW87, PMWA83]. after [CGW87, Gol81, OT80, TO86]. against [Kin83b]. agent [Sin85]. Aggregate [Bat81, HC86, BYBM87, CH86, Har82, Koh82, Nic85]. aggregated [Bak83, PN82]. aggregates [Bar80a, Bar80b, BHW86, CN80, Off80, TG80, TSF80]. aggregates-comment [CN80, Off80]. aggregates-reply [Bar80b]. Aggregation [Ant86, Bar80a, BHW86, Gra80, PS85b, Wei84]. aggregation-theoretic [BHW86]. agricultural [BYBM87]. aid [CR82]. airline [SGJ86]. Airlines [Sic85]. Algorithms [BY84, Bun88, CDG88, Wam80, WE83]. Allocative [SGJ86, KD82, Kum87, SL80, Zie83]. Alternative [BBR88, FM81, HR85, MT81, Shi81, WE83, Bal81a, BN84, Cam87, CCH84b, DM83, GMD83, Har82, Kin81, MWD83, Mar80, PN86, Ros85, SL84, Sin85, Sto80]. Alternatives [ATG83, AH89, BM87, Kie85b, Rog86]. ambulatory [OS80]. Amemiya [Ame83a, Bal83]. American [Ano81c, Ano81b, Bre81]. among [Har82, RT80a, Sha85]. amounts [NN85a]. analog [Agb88]. Analysis [Lee82a, Ano80z, BH84, BHL89, BJ87, CDJ85, CC80, CCG^{+}85, CI87, De 83, DS83, Dor81, Duf82, EY85b, FM83, GM85, GWK89, Gre83b, Han87b, Han87a, HM88, HS84, HNS88, Hol82, Jar87, Kad85, Kie82, KVT85, Law87, Lee81, Lo86, LP88, Lub85a, MaC82, Mag82, Man85, Man86, MR81a, Mon83, Mun81, Pic89, PS85a, RS88, Shi88, Sic85, TS82, Tsu82, VW89, Var85, VZ85, YN81, ZR84, Zel88a, ZBV88, vDK80, CCSH84]. Analytic [GMD83]. Analyzing [HM83, FH82]. Annals [Ano80g, Ano81g]. Announcement [Ano80c, Ano80d, Ano80e, Ano81e, Ano81f, Ano82b, Ano83b, Ano84a, Ano84b, Ano86c, Ano86d, Ano87b, Ano88b, Ano88c, Ano88d, Ano89a]. Announcements [Ano80f]. Antithetic [Gew88]. Appliance [HT84]. appliances [Har82]. Application [Sat85, AL84, AT87, Bar80a, BR85, Bel81, Bie87, BH88, BW82, CR82, Dun87, EGH89, ED80, GS81, HK85, HS89a, HW81, Jar87, Kha89, Law87, Lo86, MI80, SN86, SJ86, Sic85, SW80, Sla89, Tri80, TWI86]. applications [And88, Dor81, EY85b, FM83, Lan85a, ZM85]. Applied [Ano80g, Ano81g, Lan85b]. approach [AT87, BJ82, BS82, Das85, DF83, Dun87, Eng82, GK84, GS81, HS81a, Kle88a, LV87, LH81, Maa86, MV86, Mor86, RT80a, Thu87, TV81]. approaches [CP87, Pol84]. appropriateness [PR81]. Approximate [Dun80, AH89, CM89, PH83]. approximation [LM88, ZH88]. Approximations [Dur81, Str83, Pal81, Tse82]. April [Ano80w, Ano81-29, Ano81-30, Ano82p, Ano83s, Ano85q, Ano86s]. arbitrage [SW88]. arbitrary [Kie80, Sch83]. arch [Gew89, GWK89]. ARCH-quantile [GWK89]. ARIMA [De 80, Van81]. arising [Kro88]. ARMA [Cam86, Lüt84, Mon83]. Armax [Bie87, Bai80]. Aspects [Sin82, Das83, Jen80, Lub85b]. Assessing [BCH81, Bel81, Bel82, HHCS^{+}81]. assessment [GD85, Kin81, Sla86]. asset [HW81, MR81b]. associated [Bal80, FM81]. association [Ano81-32]. assumption [Lee84, Smi87]. assumptions [Lee83b, New87b]. Asymptotic [Cha86, Cha87, FMKT82, GM81b, Han89, Man85, MT84, Sai89, ATG83, BCC81, DS85, Han85, Lüt89]. attainable [Dur88a, Fie88]. attributes [Maa89, TSF80]. August [Ano81c, Ano85r, Ano80x, Ano81-28, Ano82n, Ano83r, Ano86t]. Austria [Ano83i]. author [Ano89w]. auto [Cha83]. autocorrelated [FG83, GD80, GB84, Mae80, Now83, Oht82, PM80]. Autocorrelation [KG84, Kin89, Sha85]. autocorrelations [De 80, DR85]. autocorrrelation [Kin89]. automobile [Bre81, TT83]. Autoreg [HS80]. autoregressions [HS89a, Oha88]. autoregressive [AN80, Bol86, DG83, FH80, HS80, HE88, HMP88, Kin83b, Kin84, Kin85, LT80, MP89b, NS85, NS88a, PF85, PT84, Sar81, ST85, Str83, TM84, VR87]. auxiliary [Mit85]. average [AN80, Bal80, Dav81, Kin83b, MD83, PH83, Pie80]. Avoiding [Zam84]. award [Ano81t].

Balestra [PH83]. band [BH82]. bank [BHL89]. bankruptcies [Lo86]. based [BDFW85, GWK89, Haj87, Kie85b, Man86, McC89, McF87, MV86, Mue81, Now89, NS89]. Bayes [BB84, CTJ88, Roe84]. Bayesian [AP83, BH84, BB84, BS82, CDJ85, Dia82, GM85, Gew88, GD80, Hol82, Kad85, KVT85, LH81, Lub85a, Mon83, NS88b, Oht82, PS85a, Pol85, RS88, SP84, TS82, Tsu82, TT83, ZR84, ZM85, Zel88a, ZBV88, ZH89, vDK80]. be [NN85a]. behavior [MP82, NN85a, Var85, Web85]. behavioral [Ano85e]. Bell [SN86]. Benefit [CCSH84]. benefits [GK84]. Bernard [Ano83c]. best [LV87, MT89]. between [AJR83, BHW86, BDFW85, GNW89, MDR87, SW80, SBN89, TO86]. between-group [GNW89]. bias [BK85, CR82, Duf88, Gal81, Gre83b, Lee82a, NN85b, PF85]. biased [Tre84]. binary [Man86, MT89]. binding [Ran87, WW83]. bisection [GD85]. Bivariate [BM87, Poi80b, SP84]. blacks [WO81]. board [Ano80j, Ano82e, Ano85g, Ano86h, Ano86i, Ano87g, Ano88h, Ano89d, Ano89e, Ano80i, Ano80k, Ano81j, Ano81k, Ano81l, Ano82d, Ano82f, Ano83d, Ano83e, Ano84c, Ano84d, Ano85f, Ano85h, Ano85i, Ano86g, Ano87e, Ano87f, Ano88g, Ano88i, Ano89f]. Book [Ano80g, Ano81g]. Bootstrapping [SS89]. both [PN82]. bounded [Kra81]. bounded-influence [Kra81]. Bounding [Kle88b, KP87]. Bounds [Kin83a, Kro88, Han85, MR81a]. Box [AP81, HW81, Kha89, Poi80a]. building [MR81b]. business [PA84].

cake [NPM80]. calculating [Han85]. Calculation [ZH88]. calculations [Han89]. call [Ano81a]. calls [PMWA83]. can [NN85a]. Canadian [HH84]. capacity [BF86b, Hul86, Mor86]. capital [ED80, Mee80, RS87]. care [BW87, GLS87, OS80]. Carlo [BR85, CG85, Gew88, GP83, GS86, LM89, MDR87, Nic85, Oha88, OSW80, Paa84, Sla86, ZBV88, vDK80]. cars [BCH81]. case [CCSH84, HM86, HIHI81, Shi81, TSF80, YJM84, Zam81]. categorical [Bar83, FM83, Mut83]. Causal [Poi88, SI88, GS88, Zel88b]. Causality [Bas88, HIHI81, Zel88b, CL82, EW83, GMD83, Gra88, HS89a, Lüt82, PT84, SW89b, SV88, Tjø81]. causation [Sky88]. caused [BK85]. Censored [Pow86, Che85, CI87, Dun86b, Fer86, Hor86, Law87, Paa84, Pow84, SM83a]. censoring [Cha86]. census [HH84]. central [BJY84, Xie88, Xie11]. certain [BJY84]. change [Ant86, BS82, Dia82, Hul86, Sla89]. change-points [BS82]. changes [Bla89, Sal82]. characteristic [VMD89]. characteristics [MT86, PMWA83]. characterizations [EW83]. Charging [PMWA83]. check [Hum81]. Chi [And88, Xie88, Xie11]. Chi-square [And88, Xie88, Xie11]. Choice [MV86, Bun88, Har82, Kin81, MDR87, MS80]. Choice-based [MV86]. chronic [Kad85]. Cincinnati [Ano81c]. citizens [Lan85c]. class [DS84, EK88, Far87, PP86, SS83]. Classical [Tur89, Das85, Kle88c, LV87, Sai89]. Classification [Ger80]. co [EY87]. co-integrated [EY87]. CO2 [Ano82r]. coarseness [Cam87]. Cobb [Sch88]. Cochrane [Mag87, Tay81, Tho87]. Cochrane-Orcutt [Mag87, Tay81]. coefficient [BM80, HS81a, Kie82, Kni80, LT80, MR81a, PZ85, Phi84a, RSU80, Rao82, Shi88, TS82, WE83]. coefficients [CGW87, De 83, FMKT82, Gre83b, HP85a, KS86, KP87, LY89, OT80, Oht87, Pag80, Pal80, Par82a, Par82b, SM80, ST80, TO86, VMD89]. cointegration [EGH89]. collateral [Tri80]. collinearity [Bel81, Bel82]. Combined [Ano80h, Ano81h]. Comment [WF83a, And80, Cho81b, CN80, Off80, Poi80a, Ran87, Sto80, SM81]. comments [Dij83]. commercial [Tis83]. common [Cap86, Con85, RT80a]. common-property [Cap86]. comparative [BH82, SI88]. Comparing [GMD83, CM89, GP81]. Comparison [Ros85, AP81, Ame83a, Bun88, CCSH84, Cho81a, MJ88, Paa84, PN86, Swa80, vdBFHM80]. comparisons [DF83, Mit84, Wes86]. Competition [Ano88e]. Competitions [Ano86e]. competitive [CH86, VZ85]. complete [BSB84, PS85b]. component [Bal81a]. components [Bal81b, Kan85, Mag82, ST85, Sic85, Sla89, WK89, Zie83]. composite [MD83]. comprehensive [Pes82]. Computation [Wal85, Bel80, Hal84]. Computational [CPW87, Jen80, BSDS88, CDG88]. Computations [GG80]. computed [VKGB85]. computer [Dun80, HS80]. computing [Ful80, Xie88, Xie11]. concept [Gra88, Mun81]. Conception [OF85]. Conditional [Phi88, WO81, Bol86, Cha87]. Conditioning [Bel88a]. conditions [Bek89, Whi82b]. Conference [Ano81s, Ano81-30]. Confidence [Hal83]. confluent [BJY84]. Congress [Ano83i, Ano81-32]. conjugate [RS88]. consensus [Dur88a, Fie88]. consequence [RS87]. conservation [Sat85]. considerations [Tay80, Zel81]. Consistency [CCH84a, DP85, Sha85, Web85]. Consistent [Bie82, HM86, Hsi89, MD83, Pal80, Ruu86, Hsi83, MW85]. constancy [TS85]. constrained [Cha85, GG80, HM83, Tho82]. constraints [GS88, GHM81, LP87, Mit85, PT84, Ran87, TWI86, WW83, Wol89]. constructed [PN84]. consumer [CL82, DP85, Har82, PW84, Ran87, SW80, WW83]. consumption [ZM85]. containing [Bai87]. contemporal [TG80]. contemporaneously [Tur89]. contents [Ano80h, Ano81h]. contents-parts [Ano80h, Ano81h]. continuous [Agb88, GK84, HS89a]. continuous-time [Agb88, GK84]. Continuously [Maa89]. contours [Hal83]. contract [Ken85]. control [Cho81b, Nor81a, Nor81b, RS87]. Convenient [DM84]. corporate [Lo86]. Correcting [BK85]. Correction [Ano86f, Hon88]. corrections [BN84, NS88a]. correlated [SL80, Tre84]. correlation [CL85, De 83, DD85, Han87a, HP81b, Kin83a, Lev83, She85]. Correspondence [DS83]. Corrigenda [Ano81i, Ano87c, Ano87d]. Corrigendum [Ano83c, Ano85d, Ano89b, Ano89c, Kra13, NS87, Xie11]. cost [BDFW85, Bre81, CG85, KD82, SL80, Zie83, CCSH84]. cost-of-living [BDFW85]. Costs [GK84, SJ86, SW88]. count [Mul86]. countries [AT87, Ano86e, Ano88e]. country [BYBM87, HHCS^{+}81]. covariance [And89, Are89b, Are89a, BP86, Cha85, Han85, Hsi83, Kie80, MW85, Sch83, Spe81, Str83, Vin82]. covariances [Che85]. Cox [ATG83, AP81, Das85, HW81, Kha89, Poi80a, Whi82b]. Cox-transformed [Kha89]. credit [BHL89]. criminal [SW89a]. criteria [Aka81, Atk81, Bou89, Cho81a, WF83b]. criterion [HS89b, Roe84]. critical [Roe84]. crop [AT87]. cross [BYBM87, Biø81, Dea85, Dea87, Kie80, LH81, PZ85, TWI86]. cross-country [BYBM87]. cross-equational [TWI86]. cross-price [Dea87]. cross-section [Biø81, PZ85]. cross-section/time-series [Biø81]. cross-sectional [LH81]. cross-sections [Dea85, Kie80]. currency [HM88]. Current [Pri88]. curvature [GG84]. Curves [Kra13, VA89]. customer [CCH84a, PA84]. cycle [TT83]. cyclical [Sla89].

Data [LP88, Pie80, AL84, AH82, Bak83, Bar83, BC88, Bel82, Biø81, Cam87, Cha82a, CCG^{+}85, Che85, DS89, Dea85, Dea87, DP85, Dun80, FM83, Gra81, GH86, Haj87, HH84, Har82, HT81, HR87, Kad85, KW83, Kin83a, Kro88, Lee81, Lee87, MaC82, Mul86, Mut83, Nic85, Old88, PN82, PM80, PZ85, Par80, PK81, Sch88, Tay80, VKD88, YN81, ZM85]. data-generating [VKD88]. day [HT84, HV84, PA84]. debt [Haj87]. decade [Ano85e]. December [Ano80y, Ano81-27, Ano82q, Ano83t, Ano84m, Ano85y, Ano86u]. Decision [MS80]. decisions [Mue81]. declaration [Çin87]. decomposability [SI88]. decomposable [Cow89]. decomposition [KD82, Zie83]. Dedication [Ano88f]. deductibles [Ano83c, VV81]. defects [Kiv85]. degeneracy [GKP80]. degree [App82]. degrees [BN84]. demand [ASL87, Ano83c, Att85, BHW86, BDFW85, BCH81, Ber82, BSB84, Cha83, CH86, EY85b, HIHI81, Iza80, Kie84, Lew87, MP89a, Mee80, NPM80, PN86, Ran87, SA89, Sta84, Tis83, VV81, VZ85, WW83]. demand-side [BHW86]. densities [Dur81, Gew89, Mis83, RW81, Tse87]. density [RT80b, RS88]. Departures [Bre81]. dependence [BKS85, Cha81]. dependencies [Mar85]. dependency [Lyn85]. dependent [And87, DW82, FG83, Gre83a, HM86, New86, New87a, PP86, Ruu86, Smi87]. deposits [PZ85, TGF81]. depreciation [HW81]. derivation [Pol84]. derived [BDFW85]. despite [Ruu86]. Detecting [Tal83, Bel81, GKP80, NN85b, PT84]. detection [Dia82]. determinants [BYBM87]. determination [ELW85, EY85a, Gal82, Kni80, Pal81]. deterministic [vdBFHM80]. developing [Ano86e, Ano88e]. development [Gra88, The89]. developments [Pri88]. deviations [BDFW85, KD82, PN84, Pow84]. Diagnostic [Tau85, And88]. diagnostics [Eng82, Kie85b, Kle88c, KVT85, ZM85]. dichotomous [GM81b, Kle88b, ZR84]. difference [Tse82]. different [CCSH84, VKD88]. differentials [GNW89]. differentiated [PW84]. dimensional [BLW85, GM85]. dimensionality [Nic85]. disaggregated [PN82, Spe81]. disaggregation [TSF80]. discovery [PS81]. Discrete [Har82, Agb88, BH88, CL85, Lee87, Man85]. discriminant [Lo86]. Discussion [Fie88, Hin85, Dor81, Dur88b]. disequilibrium [Ger80, GQ81b, sH80, KVT85, Lee82b, Lee86, Lub85b, May89, MO86]. disguised [TGF81]. distance [BSDS88]. distortions [SGJ86]. distributed [Lüt81, Maa89, MS86, PH81]. Distribution [DL87, ATG83, Bou89, BJ87, HS89b, Hor86, Lüt89, Mof85, PMWA83, Phi84a, Phi84b, Ruu86, SP84, UCS84, Xek84]. Distribution-free [DL87, ATG83, Hor86]. distributional [HV84, Lee83b, Lee84, New87b]. distributions [Don86, FMKT82, Mas85, MT84, TM84, WO81, ZH88]. disturbance [Duf88, GD88, HK85, MD83, Oht87, TO86]. disturbances [Cam86, Cra82, DS85, DG83, EK85, Gew89, GQ81a, Kin84, Kin85, Kin89, Kni85, Mae80, Mag85, Tur89, USC83]. dividend [LWD87]. does [BW87]. domain [GS81]. dominance [Bou89]. domination [Mit85]. Double [GD85, DS84]. Double- [GD85]. Douglas [Sch88]. Dublin [Ano81-31]. due [Lüt82]. duration [BKS85, Cha81, HS84, Ken85, Kie85b, Lan85a, Wal85]. Durbin [DD85, Fie88, Kin81, Kin83a, Krä85]. Dynamic [Mee80, NN85a, AH82, Bai87, Blo89, BH82, Cam86, CH86, CP87, DA89, ELW85, Gho89, Gra80, HS80, HR82, Kiv85, MP89a, MD81, MB81, Now83, Now85, PZ80, PZ81, Pet89, PN86, WE83, Wat89]. dynamics [FH82].

earnings [BH88, HS89b, MaC82, WO81]. eating [NPM80]. Econom [Kra13, Xie11]. Econometric [CC80, GQ81a, HS84, HNS88, NS88b, NPM80, AJR83, AMS83, And88, Bas88, BCC81, BHL89, Cha83, Duf88, FM83, GK84, Gra81, Gre80a, Haj87, HR87, HS80, Kie85b, Lee82b, Lee86, Pag89, SV88, TB88, Wol89, Ano81a]. Econometrics [Ano80g, Ano81g, Ano83c, Aig84, HR82, McC89, MO86, Phi86, Zel88a]. Economic [Bar80a, Bar80b, CN80, Off80, Ano81-32, Biø81, HS81a, HW81, KVT85, Nic85, Pie81, Poi80a, Sin85, Slo89, Web85]. economics [Kle88a, Zel88b]. Economies [Hum81]. economy [FJ81]. Edgeworth [Tse82]. Editor [AD89, WF83a, Aig84, Ame84a, Ano81m, Ano87h, Bar80c, BG85, Bel88b, Blu87, Den80a, Dun86a, Jud84, Kie85a, Mad81, Slo89, Tri87, Van85, Whi82a, Whi83]. Editorial [Ano80j, Ano82e, Ano82c, Ano85g, Ano86h, Ano86i, Ano87g, Ano88h, Ano89d, Ano89e, Ano80i, Ano80k, Ano81j, Ano81k, Ano81l, Ano82d, Ano82f, Ano83d, Ano83e, Ano84c, Ano84d, Ano85f, Ano85h, Ano85i, Ano86g, Ano87e, Ano87f, Ano88g, Ano88i, Ano89f]. Editors [AS80, AZ88, Ano83f, BF86a, HS82, HS89c, KH88]. education [Lan85c]. effect [Kie80, Sch83, Sch88]. Effects [Mar80, Sto80, Bre87, HT81, HE88, Kle88b, KP87, Mou86, Oha88, PW84, Sha85]. efficiencies [BC88]. Efficiency [JS85, Mag85, Nel84, CPW87, Cha86, Cha87, CPC89, DS85, Dor81, DG83, FLS80, Han89, Hul86, KD82, Pal81, RSU80, Sai89, Tay81, Tho87, Wal84, Zie83]. Efficient [New87a, Are89b, Are89a, Bel80, CCG^{+}85]. efficiently [Koh82]. eigenvalues [Str83]. elasticities [CG85, Dea87, MR81b]. electric [BCH81, SJ86]. electricity [Aig84, CC80, CCSH84, CCH84a, CCH84b, EGH89, GK84, HT84, HV84, KM84, PA84, PW84, Tis83]. elementary [Lan85c]. elicitation [Lub85b]. Elliptical [Kra13, VW89, VA89, CTJ88]. Empirical [CL82, BB84, CCG^{+}85, ED80, EY85a, HR82, LWD87, Slo89, SBN89, VZ85]. employed [HC86]. employment [Gar89, Mue81]. employment-output [Gar89]. endogeneity [Att85]. Endogenous [ED80, Nel84, New86, New87a, PR81, Sha88]. entropy [Vin82, ZH88]. Entry [VZ85, CH86]. envelopment [CCG^{+}85]. equality [TO86]. equation [Ben83, Cam86, Cha82b, CS87, FMKT82, HP81a, HP81b, Kie82, Kiv85, Kni80, MD81, Mut83, Now89, PZ80, PZ81, Pol84, Tsu82, ZBV88]. equational [TWI86]. equations [Ame83a, And80, Are89b, Att85, Bal81b, BH84, Con85, DS84, Jen80, Lan85b, Maa80, MJ88, MD83, MS80, MT84, RS88, SA89, Tse82, Tur89]. equilibrium [BF86b, Cap86, HK86, Kul85, SN86, Sla86]. equity [JS85]. equivalence [Kan85]. equivalent [Bas88]. Errata [Ano83g, Ano89g]. Erratum [Ano80l, Ano80m, Ano81n, Ano81o, Ano82g, Ano83h, Ano85j, Ano88j, Ano88k, Ano89h, HS85]. error [Bal81a, Bal81b, CGW87, CM89, Don86, Gho89, GD80, Ham86, HMP88, JYBB84, Kle88b, Kro88, Lee82a, Lev85, MaC82, Mag82, MP89b, MF82, Mor84, Now83, Now85, PM80, Pie81, ST85, Sic85, Ter82, Thu87, Var85, Vin82, WK89]. error-components [WK89]. Errors [GH86, Sta84, Ame85, BCC81, Blo89, CTJ88, Çin87, DA89, FG83, Gre83b, GB84, HS80, Hsi89, Kin83b, Kle88c, Now83, Now89, Oht82, PN84, Pal80, Pet89, Sar81, Sha85, Tre84, VR87]. Errors-in-variables [Sta84, Ame85, Blo89, DA89, Hsi89, Kle88c]. essentially [Gal81]. estimate [Pol85]. Estimated [SJ86, Gre83b, Ham86, Lüt89, NN85a]. Estimates [Bre81, RS87, Bel81, FP80, FMKT82, Hal84, LY89, MPV86, Mou86, PZ85, Vin82]. Estimating [BCC81, Biø81, CG85, CGW87, Con85, GM81a, GQ81b, HS89a, HV84, May89, PN84, PM80, SL80, Ter87b, NPM80, RSU80]. Estimation [AL84, Cra82, Dea87, Gre83a, Kie80, KG84, MT89, OT80, Sch88, VMD89, WW83, WK89, AH82, App82, Are89b, Bak83, Bal81a, BB84, Blo89, Bre87, Bun88, BW82, CPW87, Cha87, Cha82b, CS87, CHO83, Dav81, Den80b, DG83, DL87, Dun80, Dun87, Fer86, FL81, GD88, Gho89, Gre80b, Gre80a, Gre82, HH84, HM86, Hsi89, HW81, Jen80, JLMS82, JYYT87, Kra81, LE81, Lee83a, Lee84, Maa80, Maa86, MP89a, Mag87, MT86, MD81, MD83, New86, New87a, New88, Oht82, OS80, Pag80, PZ80, PZ81, PF85, Par80, Pow84, PS85b, PN86, Ruu86, Sal86, SP84, SW88, Ste80, ST80, SM83b, Tay80, Van81, WE83, Wes86, ZBV88, vdBFHM80]. estimator [AP81, Ame85, AS81, Are89a, Bel80, BB84, Cam86, Car81, Cha85, Dun86b, GM81b, GT84, GD80, Han87a, Han89, Hor86, Hsi83, JYBB84, Kha89, Man84, Man85, NS88a, Nel84, OH84, PH83, Phi84b, Rao82, RW81, Sha85, Tay81, Tho82, Tho87, Vin80, Vin81]. Estimators [Duf88, Zam81, Ame83b, And80, AN80, BY84, BH82, DG83, Dun87, Dur81, DS84, FG83, FGP84, GWK89, GB84, GP81, Han85, HK85, HZ84, MP82, MJ88, MW85, Mae80, Mag85, Mit84, Mit85, MT84, OSW80, Paa84, Pal80, Par82b, Phi84a, Pol84, PP86, RV88, Roe84, She85, SS83, Swa80, SM81, Tre84, USC83, UCS84, VKD88, Wal84, YJM84]. Euro [Ano83i]. European [Ano83i, Ano86a]. evaluating [HR85]. evaluation [Bek89, BY84, GS86, Pag89, RT80b, Tau85]. Evidence [Ste89, AS81, CL82, GM80, GMD83, GP83, Kie84, Mae80, MDR87, SW89b]. Ex [CU87]. Exact [De 80, DS84, Gew89, Agb88, Bal80, DR85, HMP88, MP89b, NS88a, Pet89, Phi84a, Phi84b]. examination [BSDS88, EW83, FG83]. exceed [Dun80]. exchange [HM88]. existence [Duf88, SM80, SM81]. existing [AG84]. exit [CH86]. exogeneity [BSB84, RV88]. exogenous [Phi84a, Tse82]. expansions [FMKT82, MT84]. expectations [Ant86, Bat81, BW82, CHO83, EY85a, FL81, HS81b, Mee80, Mor86, Mue81, Pes81a, PN86, Sal86, Shi81, Sta83, TB88, Wat89, WF83b, Wes86]. expected [BM80, Par82b, Thu87]. expenditure [Çin87, Jar87]. expenditures [Lan85c]. Experience [Poi80a, Ste89, vDK80]. experiment [PMWA83, Ste89]. experimental [AL84, Bal81a, GMD83]. experiments [CC80, CCH84a]. explanation [Ant86]. explanatory [BH88, Con85, New86, New87a]. Explicitly [GM85]. exploratory [Duf82]. exponential [Gei85, SP84]. exposé [TGF81]. extended [Çin87, JYYT87, RS88]. extension [DS83, Kel82]. extensions [NS88a]. external [Haj87].

factor [Bek89, CH86, Kan85, MP89a, Mee80, Pic89, PN86, Sic85, Sla86, WE83]. factors [HC86, SN86]. failure [Lee84]. Family [Bou89, BW87, GD85]. Farebrother [Bak83]. February [Ano84h, Ano87n, Ano80t, Ano81z, Ano82o, Ano83p, Ano85t, Ano86n, Ano88r, Ano89s]. Federal [Hum81]. fees [GLS87]. Fertility [RS87, OF85]. Fienberg [Dur88b]. filter [Sla89, TSF80]. filtering [BW82]. FIML [CPW87, HNS88]. financial [MR81b]. financing [GLS87]. Finding [RT80a]. findings [KM84]. Finite [AN80, Cam86, BH82, DS84, GM81a, LM89, MW85]. Finite-sample [Cam86]. firm [BC88]. firm-level [BC88]. first [Agb88, Bal80, DG83, FH80, HMP88, Kin89, LT80, MP89b, PH83, Str83, Tse82]. first-order [Agb88, Bal80, DG83, FH80, HMP88, Kin89, LT80, MP89b, PH83, Str83, Tse82]. Fiscal [MI80]. fitting [CDG88, Nic85, PH81]. fixed [Kie80, SN86, Sch83, Sch88]. fixed-effect [Sch88]. flexible [Bar85, BLW85, BLW87, Gal81, GG84, Gre80b, MS86]. food [Iza80]. force [FH82, NN85a]. forecast [HMP88, Koh82, MP89b]. Forecasting [Ano81s, EY87, TG80, ZH89, Ano80z, EGH89, GWK89, Par82a, Poi80a]. forecasts [EGH89]. Foreign [HM88]. form [Bar85, BJY84, DW88, Gal81, LE81, Maa86, MJ88, Par82a, Pol84, SM80, SM81]. forms [BLW85, BLW87, Bel82, BF86b, Gal81, GG84, McC81, Ros85]. Formulation [AH82, HR81, HR82]. formulations [MT81]. Foundations [CCG^{+}85, Far87]. Fourier [CG85, Gal81]. fourth [Kin84, Kin89]. fourth-order [Kin84, Kin89]. Fractional [Lew87]. framework [Kel82, Tsu82]. free [ATG83, DL87, Hor86]. freedom [BN84]. frequency [Cec86, GS81, Par80]. frontier [BC88, FLS80, Gre80b, Gre80a, Gre82, JLMS82, KD82, Lee83a, Lee83b, OSW80, SL84, Ste80, Wal82, Wal84, vdBFHM80]. frontiers [Kum87, SL80]. fuels [Har82]. Full [FP80, Wes86, Bel80]. Full-information [FP80, Bel80].

Full-versus [Wes86]. Fully [Lee81, Mon83]. function [BDFW85, BC88, BH88, CG85, JLMS82, KD82]. functional [Bar85, BLW85, BLW87, DW88, Gal81, GG84, HK86, LE81, Ros85]. functions [BJY84, CCG^{+}85, FLS80, Gre80a, Lee83b, Lüt81, Lüt89, MI80, Mun81, OSW80, RT80b, Ste80, VMD89, vdBFHM80]. Further [AS81, GM80, SV88, vDK80, EK88, HP81b, LWD87, MWD83, MPV86].

G2SLS [Ame83a]. gamma [Dia82, DS85]. gas [JS85]. General [Zel81, AG84, De 80, DF83, Eng82, Gre83b, HM86, HZ84]. Generalised [GMRT87a, Lan85a]. Generalized [Bol86, New85, Ame83b, Bal83, BLW85, BC88, Bre80, FGP84, Han87a, Han85, Ste80]. generated [De 80]. generating [SJ86, VKD88]. generation [Lyn85]. generator [BR85]. generic [Maa86]. geographic [BJ87]. German [BW82]. GFT [BDFW85]. GFT-based [BDFW85]. Gini [LY89, MR81a]. Global [Now85, BLW85, BLW87]. GLS [Ame83a, Are89a, Thu87]. goal [Oht87]. Gradual [TWI86, Tsu82]. Granger [Tjø81]. Granger-causality [Tjø81]. grants [CR82]. GRF [MJ88]. group [Ano80z, GNW89, Mou86]. grouped [Cam87, Che85, CI87]. grouped-data [Cam87]. grouping [Cam87, DS89]. groups [Sat85]. growth [Hul86, Sic85, ZH89]. GTE [PMWA83].

harmful [Bel81, Bel82]. Hatanaka [SW80]. Hatanaka-Wallace [SW80]. Haugh [EW83]. Hausman [NN85b]. having [NPM80]. health [Ano83c, BW87, GLS87, NPM80, VV81]. help [VMD89]. heterogeneity [Wal85]. heterogeneous [CH86, Lan85a]. heteroscedastic [Cra82, EK85, Tre84]. heteroscedasticity [AG84, EK88, Far87, GS86, HP81a, Hsi83, Koe81, Man84, TS85, YJM84]. heteroscedasticity-consistent [Hsi83]. heteroskedastic [SM83a]. heteroskedasticity [AS81, Bol86, Hon88, Kel82, LE81, MW85]. heteroskedasticity-consistent [MW85]. higher [HS89a]. higher-order [HS89a]. Homogeneity [Att85, Ber82]. homoscedasticity [OT80]. hours [Cha83, WO81]. household [BW87, Çin87, Dea87, Jar87, PMWA83]. housing [ELW85]. human [RS87]. hypergeometric [BJY84]. hyperinflation [BW82]. hyperparameter [Tri80]. hypotheses [Das83, Das85, DM83, FM81, GMT83, MWD83, Pes81b, Pes81c, Whi82b, Zel81]. hypothesis [EY85b, Gar89, GS81, HS81b].

identically [LC86]. Identifiability [WF83b]. Identification [BP86, Bek89, Blo89, HP85a, Now83, Now89, Pes81a, Sar81, TB88, BS82, Mar81, Now85, Pag80]. Identifying [Kie82]. IFC [Ano81s]. II [Ano80h, Ano81h, Ano81m]. Illinois [PMWA83]. illustrations [NS88b]. impact [BCC81, Cam87, HR85, HV84]. Impacts [BN84]. Imperfect [MR81b, LP87, Mas85, RS87]. Implications [GM80, GLS87, Mor84, Mun81]. implied [HS81b]. Imposing [GG84]. Improved [Car81, GD88, HJ87, Vin80, Vin81, MW85, YJM84]. Improving [LY89]. impulse [Lüt89]. include [DG83]. including [BH82]. Income [Bou89, NS89, Ano89l, Bla89, Çin87, DS89, GNW89, GS81, HS89a, Sat85, SW89b, ZM85]. income-wealth [Sat85]. incomplete [Biø81, BM89, WK89]. inconsistencies [BCC81]. inconsistent [MP82]. independence [HIHI81, Phi88]. independent [CM89, Dia82, SJ86]. Index [Ano80n, Ano80o, Ano80p, Ano81p, Ano81q, Ano81r, Ano82h, Ano82i, Ano82j, Ano83j, Ano83k, Ano83l, Ano84e, Ano84f, Ano84g, Ano85k, Ano85l, Ano85m, Ano85n, Ano85z, Ano86j, Ano86k, Ano86l, Ano87i, Ano87j, Ano87k, Ano88l, Ano88m, Ano88n, Ano89k, Ano89i, Ano89j, Ano82s, Ano89w, Bar80a]. indexes [BDFW85]. Indexing [FJ81]. Indicator [TSF80, Ful80]. individual [Har82, HT81, HE88]. Induced [Ste89]. industrial [Tis83]. industries [VZ85]. industry [Bre81, Cha83, CH86, SGJ86]. inefficiency [JLMS82, Kum87, SL80]. inequality [Bla89, BM89, Cow89, GHM81, JYBB84, Maa89, NS89, Slo89, SBN89, The89, Tho82, Wol89]. Inference [Bai87, Spe81, AP83, FMR81, Gew88, NS88b]. inferences [Hsu82]. infinite [GM85]. infinite-dimensional [GM85]. inflationary [FL81]. influence [Kra81]. information [Aka81, Atk81, Bel80, BH84, Cho81a, FP80, Kie82, Mas85, May89, NN85a, RV88, Tri80, TB88, Wes86]. informative [Don86]. Initial [Ano81b, DG83]. input [BB84]. Instrumental [Ame85, Cam86, KP87, MP82, New86]. instrumental-variables [Cam86, KP87]. instruments [TB88]. insurance [Ano83c, Mof85, VV81]. integers [Sha88]. integrated [EY87, HS89a, Mor86]. integration [BR85, Gew88, NS88b, VKGB85, vDK80]. intercept [Krä85, MP89b]. interdependence [HHCS^{+}81]. interest [Ano80z]. International [Ano81-31, YN81, Ano81-32, HHCS^{+}81, The89, ZM85, ZH89, Ano81s, Ano83u]. interpersonal [Sch83]. interpretation [HK85, PS88]. Interpreting [Bla89, SW89b]. interspatial [DF83]. intertemporal [DF83, Kie80, Sin85]. Interval [Gei85, GWK89, HH84]. intervals [OF85]. interventions [HR85]. Introduction [Aig84, And88, Bro82, AS80, AZ88, AD89, Ame84a, Ano81m, Ano83f, Ano87h, Bar80c, BG85, Bel88b, Ben83, BF86a, Blu87, Den80a, Dun86a, HS82, HS89c, Jud84, Kie85a, KH88, Mad81, Slo89, Tri87, Van85, Whi82a, Whi83]. invariance [Bak83, Bre80, Sat85]. inverse [BJY84, MS86, Xie88, Xie11]. investigation [LWD87, LM89]. investment [BBR88]. investments [RS87]. involving [Kle88b]. Ireland [Ano81-31]. ISI [Ano86m]. Item [Per81, Ano85o, Ano89m]. items [Ano83m]. iterative [Mag85]. ITSM [Ano81c, Ano81-31, Ano83u].

J [Ano81t, Fie88, Kra13, Xie11]. January [Ano80q, Ano81w, Ano82l, Ano83n, Ano84h, Ano85p, Ano87n, Ano88o, Ano89p]. Japan [TT83]. job [NN85c]. Joint [KS86, LE81, HM83]. Journal [Ano83c]. July [Ano83i, Ano84l, Ano85r, Ano86r, Ano87o, Ano88t, Ano89t]. June [Ano83o, Ano84i, Ano88q, Ano80u, Ano81y, Ano85w, Ano86o, Ano89r].

Kalman [BW82, Sla89]. kappa [HS89b]. key [AMS83]. Keynesian [TS82]. Klem [TWI86]. known [SBN89]. Koopmans [CCG^{+}85]. Kuhn [GHM81].

labor [FH82, Gar89, HH84, Mee80, MB81, NN85a]. lag [Lüt81, MS86, PH81, SW80]. lagged [FG83]. Lagrange [Eng82, HK85, Hon88, Pes81b, Pes81c, Rog86]. Laguerre [Kie85b]. Large [MD81, PZ80, PZ81, PA84, VKD88]. Large-sample [VKD88]. Latent [AD89, And83, Bar83, GS81, GS88, Mut83, And89, Ben83, BK85, CL85, Lev86]. latent-variable [Lev86]. Laurent [Bar85, BLW85, BLW87]. laws [Bat81, PS88, Sat85, Zel88b]. Layoffs [BKS85]. LDC’s [Haj87]. LDV [Jar87]. Least [Pow84, AP81, Ame83b, Bal83, BN84, BM80, CHO83, Dij83, Don86, Gre83a, Gre83b, GP81, Hor86, Kha89, MPV86, Mit84, Mit85, NN85b, NS88a, PF85, Tho82, Wam80]. least-squares [Kha89, MPV86, NS88a]. Lee [Ame83a]. length [BM80]. Leonard [Ano81t]. Leontief [BLW85]. level [BC88]. library [HS80]. Lie [Sat85]. life [TT83]. like [BY84, HZ84, Maa80]. Likelihood [Aka81, Atk81, CP87, KVT85, Ste80, AP81, Bel80, Bre87, Bun88, Cha85, Dij83, Dun80, Fer86, Gho89, GM81b, GHM81, Gre80a, Gre82, Han89, HM86, Lee83a, Lee84, Lev83, Mag82, Pol84, Tau85, Tsu82, Wal82]. Limited [RV88, And87, BH84, BY84, Gre83a, Kie82, May89, NN85a, New86, New87a, Ruu86, Smi87, Wes86]. limited-information [BH84, Wes86]. limited-translation [BY84]. limits [Dun80]. Linear [And89, DA89, Gol81, Lüt84, New86, PN82, ST80, Xek84, Agb88, ASL87, AG84, AP81, BP86, BH88, Cha85, Che85, CI87, Çin87, CPC89, Den80b, DS85, Duf82, FMR81, GG80, GD88, Gew89, GB84, HNS88, HP85a, Hol82, Hor86, Kin83b, KG84, KS86, Kiv85, Lee81, LV87, Mag82, Mar85, PKK89, PP86, Sal86, SA89, Ter82, Ter87a, Ter87b, Tho82, TO86, Tre84, TB88, Tur89, USC83, Wat89, Wol89]. link [DL87]. link-free [DL87]. living [BDFW85]. load [Aig84, CCH84b, GK84]. local [CDG88, PMWA83]. location [GD85, Tal83]. location-scale [GD85]. log [Lee81]. log-linear [Lee81]. Logit [Lo86, DM84, GM81b, Lee82a, McF87, Par80, Wil87]. Loglinear [FM83]. logs [Bel88a]. Long [Gra80, CU87, EGH89]. long-run [CU87, EGH89]. longitudinal [MaC82]. look [AJR83]. Lorenz [Kra13, VA89]. loss [Dor81, JYBB84, Par82b, Thu87]. lost [Lyn85]. Luxembourg [Ano89l].

M [Ano83c]. M.L.E. [DS85]. Maasoumi [Hen82]. macroeconometric [FP80]. macroeconomic [HS83, Pal81]. Maddala [Ame83a]. magazines [Cec86]. magnitude [Bla89]. major [AT87]. Malaysia [RS87]. manufacturing [Ken85]. March [Ano81-31, Ano84k, Ano85x, Ano86q, Ano87q, Ano88u, Ano89v]. marginal [Bre81]. marginal-cost [Bre81]. marginalposterior [ZH88]. market [GQ81b, Lee86]. markets [Lee82b, LP87]. married [NN85a]. Marsh [HP85b]. matrices [Han85, Vin82]. matrix [Cha85, Hsi83, MW85, Str83, Ter82]. maturity [HM88]. Maximum [Bre87, Gho89, Gre80a, Gre82, Lee84, Vin82, AP81, Bel80, Bun88, Cha85, Dij83, Dun80, Fer86, GM81b, Han87a, HM86, Lee83a, Lev83, Mag82, Man85, MT86, Pol84, Tau85, TV81, Tsu82, ZH88]. maximum-likelihood [Bel80, HM86, Pol84]. May [Ano81s, Ano80r, Ano81v, Ano82r, Ano83o, Ano84i, Ano85u, Ano87l, Ano88q, Ano89o]. MDS [RT80a]. Mean [Cra87, PF85, HMP88, MP89b, Ter82, Zam81]. mean-square [HMP88, MP89b]. measure [BM89]. measurement [BF86b, FLS80, Kle88b, Lev85, Mor86, Now89, Sla86, Var85, Vin82]. measures [Bla89, Cow89, GNW89, KD82, Maa89, NS89, SBN89]. Measuring [PW84]. medical [BW87, OS80]. Meeting [Ano81c, Ano81-31, Ano81a, Ano86a, YN81]. Meetings [Ano83u, Ano81b]. MELO [Par82b]. memory [Gra80]. Merging [EGH89]. method [AH89, Gre83a, Han85, LM88, Maa80, New85, Pes81b, Pes81c, Pes82, VKD88]. methodologies [CCSH84]. methodology [AL84, Duf82]. Methods [CDG88, CDJ85, De 83, Dij83, FH82, GD85, HR85, KW83, PN86, ST80, Wat89, ZBV88]. Mexico [Jar87]. micro [KVT85]. micro-economic [KVT85]. Microeconometric [Kie84, LP87]. mimic [WE83]. minflex [Bar85, BLW85, BLW87]. minimax [BB84]. minimum [Par82b]. missing [DD85, Kad85]. Misspecification [McC89, Mit84, RW81, Ruu86, Sai89]. misspecifications [TM84]. Misspecified [DW82, Ter82]. mixed [Agb88, OH84, VKGB85]. mixtures [Mas85]. MLE [Lev85]. MMDA [PZ85]. Mode [jL89]. Model [BJ82, Bie84, HR81, Kiv85, PK81, Aka81, AG84, Ame83a, Ame85, Ano81-30, Ano83c, Bal81a, Bie82, Bie87, BM87, BKS85, Cha83, CI87, CH86, Cho81a, DS85, DL87, Eng82, ELW85, ED80, FP80, FJ81, GD88, Ger80, Gho89, Gra81, Gre80b, GD80, GB84, Haj87, Hal84, Ham86, Han87a, Hol82, HMP88, HHCS^{+}81, sH80, JLMS82, JYYT87, Kan85, Kha89, Kie84, Kin83b, KG84, Kle88c, KVT85, Kra81, Lan85a, Lee82b, LV87, Lüt81, MaC82, MWD83, Mag85, MP89b, McF87, MR81b, Mit84, MB81, NS88a, New87b, Now83, Now85, Oht82, PM80, PKK89, PP86, Pow84, RSU80, Rao82, Sch83, SM83a, SM83b, TM84, Tho82, Tho87, Tre84, Tri80, TS82, TWI86, Tur89, UCS84, VV81, Wal84, WK89, Wes86, Wil87, Zam84]. modeling [Mut83]. Modelling [CCH84b, NN85c, Sla89, GQ81a, PT84, Pri88]. Models [Cha81, De 83, AD89, Ame84b, And83, AH82, AJR83, AMS83, And87, And89, And88, AN80, Are89a, Bai80, Bai87, Bar83, Bas88, BP86, Bek89, Bel88a, Ben83, BBR88, BCC81, Blo89, Bre80, Bre87, Bun88, BH82, Cam87, Cap86, CDJ85, Cha82a, Cha86, Cha82b, Che85, Cho81b, CM89, CP87, CPC89, CL85, Cra82, CHO83, DM83, DM84, DA89, Den80b, DW82, Duf88, Fer86, FM83, FH82, FMR81, FG83, GG80, GM81a, GS81, Gew89, GS88, GP83, GQ81b, GM81b, GHM81, Gra80, Gre82, Gre83a, Hal83, Har82, HP81a, HP81b, HR87, HK86, HS80, HR82, HS83, HE88, Hor86, Hsi89, Hsu82, Jar87, Kie80, Kie85b, Kul85, Lan85a, Lan85b, Lee81, Lee82a, Lee83a, Lee84, Lee86, LP87, Lee87, Lev86, LT80, LH81, Lub85a, Lub85b, Lüt89]. models [Maa86, MJ88, MP89a, MT81, Mae80, Mag82, MDR87, Mar81, May89, MF82, Mon83, MO86, Mul86, NN85a, Nel84, New86, New87a, New88, Nic85, Nor81a, Nor81b, Now89, Paa84, Pag80, Pag89, PH81, Par80, Pes81a, Pes82, Pet89, Pic89, Poi80b, PS85a, PN86, RV88, Ruu86, Sal82, Sal86, Sar81, SN86, SL84, SW89a, ST85, Shi81, Sin85, SM83a, Smi87, Spe81, SW88, SS89, ST80, SV88, Tau85, Ter82, Ter87a, Ter87b, TWI86, TB88, Van81, Wal85, WE83, Wat89, Web85, WF83b, Wei84, Wol89, ZR84, ZBV88]. moderate [Cra87]. Modified [Rog86, Mul86]. moment [Ben83, Cha87, New88, Pal80, VKD88]. moments [BJY84, BM89, De 80, Duf88, Gre83a, Han85, Kni85, Man84, McC81, NS88a, New85, Pet89, SS83, SM80, SM81, VKGB85, Xie88, Xie11, Zam81]. monetary [Bar80a, Bar80b, BHW86, CN80, MI80, Mar80, Off80, Sto80, TSF80]. money [HS89a, HIHI81, SW89b]. money-income [HS89a, SW89b]. Monte [BR85, CG85, Gew88, GP83, GS86, LM89, MDR87, Nic85, Oha88, OSW80, Paa84, Sla86, ZBV88, vDK80]. monthly [EGH89, Kin83a]. mother [BW87]. moving [AN80, Bal80, Dav81, Kin83b, MD83, PH83, Pie80]. MPS [FJ81]. multi [HMP88, HHCS^{+}81, Lee86, MP89b]. multi-country [HHCS^{+}81]. multi-market [Lee86]. multi-period [HMP88, MP89b]. multicollinearity [HZ84, HJ87]. multidimensional [HM83]. multinomial [Kad85, Lee82a, McF87, Par80, Wil87]. Multiple [SM83a, CR82, HM88, TG80, Tjø81]. multiplicative [SS89]. multiplier [Eng82, HK85, Hon88, Pes81b, Pes81c, Rog86]. multipliers [BCC81]. Multivariate [Cha82a, KW83, Mag82, PT84, ATG83, Cha85, DM83, Kan85, Lee81, Maa89, Mar81, Pol85, Sic85, Smi87, TWI86, VW89]. Muth [WF83b]. Muth-rational [WF83b].

natural [JS85, RS88]. nature [PS81]. near [SI88]. negative [Lüt81, Sha88, SS83]. negativity [LP87, Ran87, WW83]. neoclassical [BDFW85, Kie84]. Nerlove [MT81]. nested [BBR88, Das83, Das85, FM81, GP83, GMT83, Hal83, Pes81b, Pes81c, Pes82, SM83a, Whi82b]. Netherlands [Bie87]. News [Ano83m, Ano85o, Ano89m, Per81]. newsstand [Cec86]. nit [TSF80]. nit-picking [TSF80]. noise [Bel81, Bel82]. nominal [DS83]. Non [BH88, EY85b, Fer86, Han87a, Lee87, Lüt82, Var85, Agb88, ATG83, ASL87, AP81, BBR88, BJY84, CPC89, Das83, Das85, FM81, GP83, GQ81a, GMT83, Hal83, Han87b, HM86, HP85a, JYBB84, Kni85, Lee84, LP87, Lüt81, Mar85, MV86, Mor86, Mun81, Pes81b, Pes81c, Pes82, Pet89, Pol85, PP86, PN86, Ran87, Sha88, Sin82, SM83a, WW83, Whi82b, Xie88, Xie11]. Non-causality [Lüt82]. non-central [BJY84, Xie88, Xie11]. Non-linear [BH88, ASL87, AP81, CPC89, HP85a, Mar85, PP86]. non-negative [Lüt81, Sha88]. non-negativity [LP87, Ran87, WW83]. non-nested [BBR88, Das83, Das85, FM81, GP83, GMT83, Hal83, Pes81b, Pes81c, Pes82, SM83a, Whi82b]. non-normal [GQ81a, HM86, Kni85, Lee84, Pet89]. non-null [ATG83]. non-optimality [JYBB84]. Non-parametric [EY85b, Fer86, Han87a, Lee87, Var85, Han87b, MV86, Pol85]. non-positive [Lüt81]. non-significant [Mun81]. non-static [Mor86, PN86]. non-stationarity [Sin82]. non-stationary [Agb88]. nonlinear [ATG83, Ame85, Bel80, BH84, BCC81, FP80, GHM81, Hsi89, Kha89, Mag82, Sic85]. nonlinearity [HK85]. nonsymmetric [MP89a]. normal [CI87, GQ81a, HM86, Kni85, Lee84, Mas85, Pet89, USC83, Zam81]. normality [Lee82b, Smi87]. normalized [DW88]. North [Ano81b]. note [Bal80, Bal83, Ber82, Cha85, DS85, GD80, Har82, HP85b, Kan85, Kha89, Koe81, Lüt89, MP89a, Mag87, Mar81, McC81, MO86, PH83, Rao82, Sch83, ST85, SS83, Tho87, Tse87, Wil87, Zie83]. Notice [Ano81b]. novel [NS88b]. November [Ano85s, Ano86p, Ano81x, Ano82m, Ano87p, Ano88s, Ano89u]. null [ATG83]. number [Bar80a]. numbers [Con85]. numerical [AMS83, BY84, NS88b, Wes86]. nutrition [BW87].

Object [Old88]. Object-oriented [Old88]. observability [Poi80b]. observables [Gei85]. observation [Dor81, PS88]. observationally [Bas88]. observations [Con85, DW82, DG83, DD85, HM86, PP86]. occupations [HS89b]. occurrence [PK81]. October [Ano84j, Ano87m, Ano88p, Ano80v, Ano82k, Ano83q, Ano85s, Ano86p, Ano89q]. odds [Atk81, Zel81]. official [Ano86m]. Ohio [Ano81c]. oligopoly [App82]. OLS [Kni85, Mit84, Pet89, Sha85, Thu87]. Omission [Dor81]. omitted [Lee82a, Lüt82]. On-site [Sha88]. one [KS86, Mis83, PP86, Rog86]. one-sided [KS86, Rog86]. one-step [PP86]. open [Agb88]. Operational [Ano83i, MT86]. operations [Hum81]. opportunities [Ano85e]. Optimal [DS89, Roe84, EK85, Kin85, Pal81]. optimality [JYBB84]. optimization [CU87]. optimizing [Var85]. optimum [Sin85]. Orcutt [Mag87, Tay81, Tho87]. order [Agb88, Bal80, De 80, DG83, FH80, GM81a, HS89a, HMP88, Kin84, Kin89, LT80, MP89b, PH83, She85, Str83, Tse82]. ordered [Bar83, HR87]. ordering [SI88]. ordinal [Ter87b]. ordinary [Gre83a, NN85b]. oriented [Old88]. orthogonal [BSDS88, Lüt89]. other [Bel82, BF86b, CP87, MJ88, ZH89]. outlook [Ano85e]. output [CH86, Cla89, Gar89, HC86, OS80]. overall [PT84]. overidentifying [GS88, Sta83]. overview [HR85]. own [Dea87]. own- [Dea87].

P [Ano83c]. Pages [Ano80w, Ano80x, Ano80y, Ano80t, Ano80q, Ano80u, Ano80r, Ano80v, Ano80s, Ano81-29, Ano81-28, Ano81-27, Ano81z, Ano81w, Ano81y, Ano81v, Ano81x, Ano81u, Ano82p, Ano82n, Ano82r, Ano82q, Ano82o, Ano82l, Ano82m, Ano82k, Ano83s, Ano83r, Ano83t, Ano83p, Ano83n, Ano83o, Ano83q, Ano84m, Ano84l, Ano84h, Ano84k, Ano84i, Ano84j, Ano85q, Ano85y, Ano85t, Ano85r, Ano85p, Ano85w, Ano85x, Ano85u, Ano85s, Ano85v, Ano86s, Ano86t, Ano86u, Ano86n, Ano86o, Ano86r, Ano86q, Ano86p, Ano87o, Ano87n, Ano87q, Ano87l, Ano87p, Ano87m, Ano88r, Ano88o, Ano88t, Ano88u, Ano88q, Ano88s, Ano88p, Ano89s, Ano89p, Ano89r, Ano89t, Ano89v, Ano89o, Ano89u, Ano89q, Ano89n]. Pakistan [ASL87]. Panel [Dea85, HT81, AH82, BC88, Cha82a, GH86, Haj87, Lee87, PK81, Sch88, Tay80]. panels [WK89]. paper [Hin85]. papers [Ano81a]. parallel [SP84]. parameter [Dia82, JYYT87, NS85, SP84]. parameters [AMS83, GHM81, HH84, SJ86]. parametric [Cha86, Dun86b, EY85b, Far87, Fer86, Han87b, Han87a, Lee87, MV86, Pol85, Var85]. Parametrization [Pic89]. Pareto [AP83, CCG^{+}85, Gei85]. part [Ano81m, MDR87, Bar80c]. Partial [Poi80b, Dij83, Han89]. Partially [Ame83b, Bal83, McC81, Par82a, PP86, SM80, SM81]. participation [Ste89]. parts [Ano80h, Ano81h]. past [HS89b, Koh82, NN85a]. patterns [RT80a]. PC [HNS88]. PC-FIML [HNS88]. peak [Aig84, CCH84b, GK84]. peak-load [Aig84, CCH84b, GK84]. perennial [AT87]. performance [NN85b, RW81, Tre84]. period [HMP88, MP89b]. permanent [GS81, HT84]. Peru [GLS87]. phenomenon [HIHI81]. Phillips [Hen82]. philosophical [Dur88a, Fie88]. picking [TSF80]. Pierce [EW83]. Pitfalls [Pes81b, Pes81c]. plants [CH86]. point [EK85, Kin85, Wal82]. points [BS82]. policy [MI80, Mar80, Pal81, Sto80, TV81]. poly [BR85, RT80b]. poly- [BR85, RT80b]. polynomial [MS86, PH81, Pal81]. pool [Iza80]. Pooling [Bal81a, Oht87]. population [SW89a]. populations [AP83]. positive [Lüt81]. possible [Man84, YJM84]. post [CU87]. Posterior [VKGB85, Zel81, Atk81, Cho81a]. potential [BCH81]. power [App82, GS86, HW81, Krä85, LM89]. pp [Ano83c]. Praag [Ano83c]. pre [GT84, KG84, Man84, Mit84, Roe84]. pre-test [GT84, Man84, Mit84, Roe84]. pre-testing [KG84]. precision [Mou86]. Predicting [SW89a]. Prediction [And87, BC88, Lüt88, CTJ88, CP87, Gei85, HJ87, KG84, Maa86, PN84, ST80]. Predictions [Bai80, Ful80, PN84]. Predictive [CPC89, Gew89]. predictor [TM84]. Predictors [FH80, MT89]. Preferences [Lan85c]. preliminary [CGW87, GB84, Mor84, OT80, TO86]. presence [Bel82, DM83, HZ84, HJ87, MWD83, Pal80, Sin85, Thu87]. present [Kin89]. Press [MT81]. price [BSB84, Cec86, Dea87, ELW85, JS85, May89, MO86]. price-adjustment [May89]. prices [Cec86, CL82, HW81, PW84, SW80, Web85]. pricing [ASL87, AL84, Aig84, Bre81, CC80, CCSH84, CCH84a, GK84, Gar89, HV84, Tis83]. principle [Sat85]. prior [Lub85b, RS88]. priori [May89]. private [Ano83c, VV81]. probabilities [Ger80]. Probability [Sky88, Cho81a, GD85, Lee81, TV81]. probit [Ano83c, DM84, Poi80b, RV88, VV81]. problem [BHL89, Iza80, NN85b]. Problems [Dav81, Sha88, Car81, Haj87, Lub85b, NPM80, Rog86, Sin85, Vin80, Vin81, Wam80]. procedure [Bak83, EY85a, JYYT87, SW80]. procedures [Bal81a, EY85b, FM81, Kiv85, Mar80, MD81, PZ80, PZ81, Pie80, Sto80, Wam80, ZH89]. process [Bal80, FH80, LWD87, NN85c, PH83, Str83]. processes [Dav81, De 80, Lüt84, MaC82, PF85, VKD88]. processing [Hum81]. producing [AT87]. product [TT83]. production [AT87, BC88, Cap86, CCG^{+}85, ED80, FLS80, Gal82, GM85, Gre80b, Gre82, JLMS82, Kum87, Lee83a, OSW80, Ros85, SL80, TWI86, vdBFHM80]. Productivity [BF86b, Hul86, Mor86, DF83, MB81, Sic85, Sla86]. Professors [Hen82]. profit [Kum87]. Program [Ano81-30, HS80]. Properties [USC83, Wal84, AN80, BLW85, BLW87, BH82, Cam86, CDG88, DS84, GNW89, GM81b, Gra81, GB84, MW85, Mae80, Man85, Nic85, OH84, Par82b, She85, Tho82, VKD88]. property [Bak83, Cap86, Par82a, Sha85]. proposals [GLS87]. proxies [Kro88]. proxy [KP87, Ter87a]. psychometric [FM83, Ano86a]. public [Lan85c]. purchase [HT84].

Quadratic [Mit85, DW88, HP85a]. qualitative [FL81, KW83, Lee81, Ter87b]. quality [Bel81]. quantal [ZR84]. quantile [GWK89]. quantiles [Pow86]. quantitative [KW83]. quasi [SN86]. quasi-fixed [SN86].

Random [LT80, Mou86, BR85, BJY84, Bre87, Dia82, HS81a, RSU80, Rao82, TS82, Xie88, Xie11]. randomness [DR85]. rank [Bek89, GKP80, Han87a, HR87, VR87]. rank-ordered [HR87]. ranks [McC89]. rate [CCH84b, HT84]. rates [BJ87, HM88, HC86, KM84, PA84, SN86, ZH89]. ratio [GHM81, LM89]. rational [BW82, CHO83, HS81b, Mee80, Mue81, Pes81a, Sal86, Shi81, Sta83, TB88, Wat89, WF83b, Wes86]. rational-expectations [Wes86]. Rationality [HS83]. rationing [LP87]. ratios [Atk81, UCS84, Zel81]. re [EW83]. re-examination [EW83]. reactor [DR89]. real [Cla89]. recidivism [SW89a]. reciprocal [Zam81]. rectangular [HM83]. Recursive [Cha82b, Duf82, Wat89, Che85, Lee81, RS88]. Reduced [Maa86, VR87, MJ88, McC81, Par82a, Pol84, SM80, SM81]. reduced-form [MJ88, Par82a, Pol84]. reexamination [Iza80]. reform [Law87]. regions [Roe84]. Regression [CDG88, McF87, ATG83, Bel81, BH88, BSDS88, Bre80, BH82, Cam87, Car81, Cha82a, Cha85, CGW87, CM89, Con85, Cra82, Dor81, Duf82, Dun86b, Fer86, FGP84, GD88, GM81a, GP83, Gol81, Hal83, Han87a, Hor86, Hsu82, Kin83b, KS86, Kin89, Law87, jL89, Lev86, LV87, LH81, Lub85a, Mag85, Mag82, MF82, Mou86, Mun81, New88, OT80, Oht82, OH84, Oht87, Paa84, Pag80, Pal80, PN82, PZ85, Pes82, Pet89, PKK89, Pol85, PP86, Pow84, Pow86, Rao82, RS88, Sha88, She85, Shi88, SM83a, ST80, SM83b, Tho87, Tre84, Tsu82, TWI86, USC83, UCS84, VMD89, VR87, Vin80, Vin81, WE83, Xek84, Zel81, ZM85]. Regression-based [McF87]. regressions [Bie84, CTJ88, DG83, DD85, Fis83, Hsi83, Kin83a, Kle88b, Krä85, Phi86, TO86, TS85]. regressive [GLS87]. Regressor [Kle88c]. regressors [Ter87a, Ter87b, Tho82]. Regularity [Whi82b]. regulation [JS85]. regulatory [BHW86, SGJ86]. related [Sat85]. relation [AH89]. relations [Biø81]. relationship [AJR83, FLS80, Gar89, SW80, SBN89]. relationships [Duf82, Gra80, Poi88]. relative [DS85, DG83, Par80, Sai89]. reliability [SP84]. remark [Lev83]. Remembrance [HS89b]. renewable [Cap86]. renewable-resource [Cap86]. repayments [Haj87]. replicability [Poi88]. Reply [Dur88b, HP85b, Bar80b, Hen82, Nor81b]. representation [AH89]. representations [Bas88, Old88]. Research [Ano83i, Ano85e]. reserve [Gar89, Hum81]. reserve-labor [Gar89]. residential [CC80, CCSH84, CCH84a, CCH84b, Har82, KM84]. Residual [CI87]. residuals [GMRT87a, GMRT87b, Lan85a, Lüt89, She85]. resource [Cap86]. Response [KM84, AL84, CCH84a, Lüt89, Man85, Man86, MT89, PA84, ZR84]. response-based [Man86]. responses [FL81]. Restricted [Mit84, Bek89, Den80b, GP81, JYBB84, MPV86, McC81, Mit85, Par82a, SM80, SM81]. Restrictions [Don82, Don83, Are89b, Are89a, BP86, Cha87, CGW87, GG84, HS81b, Kie82, New88, Oht87, PH81, Sta83]. result [MPV86]. results [Bre80, CG85, DR85, FJ81, FGP84, GMD83, HP81b, MWD83, Pag80, Tho82, Zel81]. retail [KVT85]. retrospective [BS82]. return [SN86]. returns [Mar85, Ste89]. Reverse [Lev86]. reversion [Cla89]. Review [Ano80g, Ano81g]. revisions [Pie80]. Ridge [SM83b, FGP84, Maa80]. ridge-like [Maa80]. Risk [DR89, Mit84, HZ84, Mit85, Roe84]. Robust [Hsu82, Van81, DR85, Tal83]. robustness [AS81]. role [Kra81, Pag89]. roots [Oha88]. Rossi [Hin85]. Rotterdam [SM83b]. rule [Car81, Mit84, Phi84b, Vin80, Vin81]. ruled [Mit85]. Rules [GKP80, CR82, MS80]. run [Cha83, CU87, EGH89, ED80, Gar89, MB81, TV81].

S [Ano83c]. safety [DR89]. sales [EGH89]. same [BDFW85]. Sample [CR82, Agb88, Ano83c, AN80, BH82, Cam86, De 80, DL87, DR85, Duf88, DS84, GP83, GB84, MW85, MDR87, Mas85, MD81, Nel84, Nic85, NS89, OH84, PZ80, PZ81, Sha88, Tay80, VKD88, VV81, YN81]. samples [Cra87, Kiv85, LM89, Man86, MV86, Swa80, Tri80, Vin82]. Sampling [Cow89, Par82b, SP84, TM84, UCS84, Wei84]. sanitation [BW87]. Sargan [Mis83, Tse87]. Savage [Ano81t]. scale [Dia82, GD85, Hum81]. scales [Çin87]. scaling [HM83]. Schedule [Ano83u, HT84]. schedules [Mee80]. schooling [BW87, Kie82]. sciences [Ano85e]. Score [Che85, LC86, Man85, MT86]. scoring [BHL89]. search [Lan85b, LP88, NN85c]. seasonal [EGH89, HS81a, Mar80, Pie80, RT80a, Sto80]. Seasonality [BH82]. Second [She85]. Second-order [She85]. secondary [Lan85c]. section [Ano86m, PZ85]. section/time [Biø81]. sectional [LH81]. sections [Dea85, Kie80]. sector [Cap86]. seemingly [DG83, RS88]. Selecting [LV87]. selection [Ano81-30, Ano83c, CR82, Cho81a, DL87, Gol81, HR81, JYYT87, Kiv85, Kra81, MDR87, Nel84, PK81, VV81, Zam84]. semi [Cha86, Dun86b]. semi-parametric [Cha86, Dun86b]. semiflexible [DW88]. Semiparametric [Man85, Man86]. sensitivity [Çin87, Lev85]. separable [HK86]. separate [Fis83, MF82]. separation [Mas85]. September [Ano80s, Ano81u, Ano84j, Ano85v, Ano87m, Ano88p, Ano89n]. sequences [Dia82]. sequential [YN81]. Serial [CL85, DD85, HP81b, Kin83a, Lev83, Mar85, She85]. Series [Ano81c, Ano81-31, Ano83u, AJR83, Ano80z, AH89, BBR88, Bie84, Biø81, BJ87, CDJ85, Cra82, De 80, Dea85, DS83, GS81, Gra81, HS81a, HP85a, Hsi83, Kie80, Koh82, MaC82, Mar81, Mon83, Now89, Pie81, Poi80a, Pri88, RT80a, Sal82, TG80, Tjø81, Wei84, Ano81b]. services [KVT85, NPM80]. sets [Dun80, PK81, TO86]. Seventh [Ano81-32, Ano81-31]. several [AG84, SBN89]. share [CS87]. shift [Hsu82, Tal83]. shifts [PZ85]. shock [Gho89, Now83, Now85]. shock-error [Gho89, Now83, Now85]. Short [MB81, Cha83, CU87, EGH89, ED80, Gar89, TV81]. short-and [CU87, EGH89]. Short-run [MB81, Cha83, ED80, Gar89, TV81]. shrinkage [USC83, UCS84, Zam84, ZH89]. side [BHW86]. sided [KS86, Rog86]. sign [MPV86]. signal [Bel81, Bel82]. signal-to-noise [Bel81, Bel82]. significance [FM81]. significant [Mun81]. Simple [SL84, BJY84, CPC89, MS86, TS82, Xie88, Xie11]. simplified [Dun87, EY85a]. simplify [HR81]. Simulated [GMRT87b]. Simulation [FJ81, Pag89]. Simultaneous [Bal81b, Ben83, Jen80, Lan85b, Ame83a, And80, Are89b, BJ82, BH84, Cam86, Cha82b, DS84, FMKT82, HP81a, HP81b, JYYT87, Kiv85, Kni80, Maa80, MJ88, MT84, Now89, Pol84, RV88, Smi87, Tsu82, Tur89, ZBV88]. simultaneous-equation [Cam86, Pol84]. simultaneous-equations [Ame83a]. Single [GQ81b, GD85, Kiv85]. single-bisection [GD85]. Single-market [GQ81b]. singular [Vin82]. site [Sha88]. Sixth [Ano83i]. size [Bou89, Hon88, LM89]. Slutsky [SA89]. Small [Duf88, GP83, Nic85, OH84, Tay80, Tri80, Cra87, GB84, HNS88, Kiv85, Mae80]. Small-sample [Duf88, Nic85, GB84]. social [Ano85e, Bou89, Ste89]. Society [Ano81a, Ano86a]. Software [Hal84, GKP80, Old88]. Solution [Sal86, Wat89]. Some [CG85, Das83, DA89, Dij83, DR85, FGP84, Gra81, Gra88, HP81b, Lub85b, MWD83, MW85, Pag80, Par82b, Tal83, Tho82, TS85, Wes86, YJM84, AMS83, GB84, GS86, Hsi89, Mul86, PS85a, Spe81, Zel81]. Sources [Pie81, Hul86, PZ85]. space [Ham86]. Spain [Ano81s]. spatial [BJ87]. spatial-time [BJ87]. special [Ano80z, Spe81]. specific [Sic85, Zel81]. Specification [FMR81, GT84, Kie85b, Lee82a, LC86, Mul86, New87b, BJ82, BBR88, Bie82, Bie84, Bie87, CS87, DM83, DM84, Don86, GM80, Gra81, HR81, HS83, Kan85, Kro88, Kum87, Lee84, Lee86, Lo86, MWD83, MF82, McF87, Mor84, New85, Thu87, Wil87, ZBV88]. specifications [SL84, Sin85]. Specifying [HR87]. spectrum [BH82]. speculative [Web85]. spells [Mof85]. split [SW89a]. spurious [Oha88, Phi86]. square [And88, HMP88, MP89b, Ter82, Xie88, Xie11]. squared [JYBB84]. squares [AP81, Ame83b, Bal83, BN84, BM80, CHO83, Dij83, Gre83a, Gre83b, GP81, Hor86, Kha89, MPV86, Mit84, Mit85, NN85b, NS88a, PF85, Tho82, Wam80]. stability [Duf82, Lüt88, PKK89, Shi88]. stable [Bat81]. stage [AP81, Ame83b, BN84, CHO83, Dun87, HH84, Kha89]. standard [BCC81, Gre83b, Ham86, Kel82, PN84]. State [Lyn85, Ham86, Slo89]. state-space [Ham86]. static [Kul85, Mor86, PN86, SN86]. statics [SI88]. stationarity [Sin82]. stationary [Agb88, LM88, ST80, Wal82]. statistic [Kin81, NS85]. Statistical [GNW89, Far87, Kle88a, Mor84, Old88, Pag89, Phi88, Tho82]. statisticians [Ano86e, Ano88e]. statistics [Ano86m, Dur88a, Fie88, Hal83, LC86, Nic85, Sai89]. stein [Vin80, BB84, BY84, Car81, HZ84, JYYT87, Mit84, Mit85, Phi84b, Vin81]. Stein-like [BY84, HZ84]. stein-rule [Vin80, Car81, Phi84b, Vin81]. Stein-ruled [Mit85]. step [CHO83, FG83, Nel84, PP86]. Stephen [Dur88b]. Stochastic [CS87, PZ85, BP86, Gre82, JLMS82, Kum87, Lee83a, Lee83b, LM88, OSW80, SL80, SL84, Sla89, Ste80, ST80, Ter87a, Tse82, TWI86, Wal82, Wal84, vdBFHM80]. stochastically [Pag80]. stock [Mar85]. strategies [NS88b]. stratification [Sha88]. strikes [Ken85]. Structural [Sal82, AJR83, Cho81b, FH82, HH84, Hsu82, Lüt88, Maa86, MS80, MT84, Mut83, Nor81a, Nor81b, Par82b, PKK89]. structure [And80, Ben83, DA89, HK86, MaC82, Now89, PS81, Shi81, Spe81]. structures [And89, CCH84b, PS85b]. student [CR82]. Studentizing [Koe81]. study [AG84, Ano89l, Bal81a, BH82, CCSH84, Cec86, Nic85, Oha88, OSW80, RW81]. Subject [Ano82s, Ano85z, Ano89w, MF82, PH81]. subjective [GD85]. subset [PT84]. substitution [CG85, OF85]. subsystems [Cam86]. sufficient [Dur81]. summer [Ano89l, Ano81a]. supply [AT87, BHW86, BYBM87, CH86, HH84, VZ85]. supply-side [BHW86]. surprise [Bai87]. survey [Ame84b, Dea87, FLS80, YN81]. surveys [NS89]. survival [SW89a]. Swamy [Rao82]. switching [Hol82, Lub85a, Oht82, PR81, PS85a, Tsu82, TWI86]. symmetry [SA89]. system [Agb88, BDFW85, Çin87, FMKT82, Kiv85, MT84, SN86, Sch88, TS82]. Systematic [Wei84]. systems [Att85, Ben83, BH84, BSB84, CS87, DR89, EY87, GMD83, HNS88, Kel82, Kni80, Lew87, MD81, PZ80, PZ81, Pol84, Ran87, RS88, SA89, Sta84, WW83].

t [NS85]. tables [HM83]. tax [Law87]. tea [AT87]. technical [BC88, JLMS82, KD82, Kum87, SL80, Sla89, Wal84, Zie83]. technique [MS86]. techniques [Law87, Zam84]. technological [Ant86]. technologies [Gal82, GM85, GM80]. technology [EY85a, OS80, Sic85]. telephone [PMWA83]. temporal [GMD83, Wei84]. temporally [PN82]. Temporary [Cap86, BF86b, HK86, Sla86]. term [Shi81]. terms [MD83]. Test [Lee82b, Wam80, ATG83, Bel81, Bel82, CGW87, CM89, Das85, EK85, GT84, GB84, Hal83, Hon88, sH80, Kel82, Kin81, Kin83a, Kin84, Kin85, Kiv85, Koe81, Krä85, Lee83b, Lee84, LC86, Lo86, LM89, Man84, Mit84, OT80, PKK89, Roe84, Sai89, SN86, TO86, TT83, Whi82b]. Testing [DM83, GMT83, HP81a, HP81b, HK86, HS81b, HE88, Kin83b, Kin89, MF82, NS85, SA89, Sin85, Smi87, Sta83, TO86, Wol89, Bal81a, Ber82, Bie84, Bie87, Cra82, Das83, EY87, EY85b, GQ81b, HK85, HR87, KG84, LE81, LC86, Lee87, MD81, Mor84, Mul86, New85, Oht87, PZ80, PZ81, Pes81b, Pes81c, Pes82, SM83a, SV88, Tau85]. Tests [DP85, Fis83, GP83, Kul85, MWD83, AG84, And87, And88, Bas88, BJ82, BBR88, Bie82, Che85, CU87, DM84, DD85, DR85, EK88, Far87, FM81, GMD83, GHM81, GS86, HS83, Kan85, KS86, Lüt88, McC89, McF87, NN85b, New87b, RV88, Rog86, SL84, Shi81, Shi88, Tal83, TS85, Tur89, Wil87]. their [FLS80, Gra81, Kiv85, PN84, UCS84]. theoretic [BHW86]. theoretical [FGP84]. theories [Gar89]. Theory [ED80, Bar80a, DA89, Lan85b]. things [HS89b]. thoughts [SV88]. three [BLW85, BN84]. three-dimensional [BLW85]. threshold [PS85a]. Time [Ano80z, Ano81c, Ano81b, Ano81-31, Ano83u, Agb88, AL84, AJR83, AH89, BBR88, Bie84, BJ87, CDJ85, CC80, CCSH84, CCH84a, Cra82, Dea85, DS83, GK84, GS81, Gra81, HS89a, HT84, HS81a, HP85a, HV84, Hsi83, Kie80, KM84, Koh82, Lee84, LH81, MaC82, Mar81, Mon83, Now89, OF85, PA84, PW84, Pie81, Poi80a, Pri88, RT80a, Sal82, SW89a, TG80, Tis83, Tjø81, Wei84]. time-differentiated [PW84]. time-of-day [HT84, HV84, PA84]. time-of-use [AL84, CC80, CCSH84, CCH84a, KM84, Tis83]. time-series [AJR83, BBR88, Biø81, Cra82, Mar81]. time-varying [LH81]. Tobin [Iza80]. Tobit [Ame83a, Ame84b, AS81, BM87, Hal84, New87b]. too [NPM80]. Toronto [Ano81b]. Total [Sla86]. Total-factor-productivity [Sla86]. trade [TT83]. trait [And83]. transaction [SW88]. transfer [MI80]. transferability [AL84]. transferable [KM84]. transformation [Bal80, HW81, Poi80a]. transformations [Han87b, Lüt84, Sat85]. transformed [Kha89]. transition [SGJ86]. translation [BY84]. translog [Bar85, BLW85, TWI86]. Trend [Cla89]. trended [PM80]. triangular [Are89a]. Trost [Ame83a]. true [BDFW85]. truncation [BK85, Sha88]. TSA&F [Ano80z]. Tucker [GHM81]. turnover [BKS85]. Two [CHO83, HH84, Pol84, AP81, Ame83b, AH89, Bal81a, BLW87, BN84, Dun87, Fis83, FG83, Hal83, Kan85, Kha89, MDR87, Nel84, Tho82, TO86]. two-factor [Kan85]. two-part [MDR87]. Two-stage [HH84, AP81, Ame83b, CHO83, Dun87, Kha89]. Two-step [CHO83, FG83, Nel84]. two-way [Bal81a]. type [HP85a].

U.S. [BJ87, Cha83, FJ81, Ken85, Sic85, SGJ86, TT83]. Unbiased [Gal82, Gal81]. unbiasedness [Rao82]. uncertain [HR81, Maa86]. unconditional [Phi88]. uncorrelated [Tur89]. Underestimation [Ter82]. undersized [Swa80, Vin82]. Understanding [Phi86]. Unemployment [Mof85, Bie87, BJ87, Cla89, Lyn85]. unequal [CM89, Con85]. unit [Oha88]. units [SJ86]. unknown [GM81a]. unobservable [HT81, OS80]. unobserved [BW82]. unrelated [DG83, RS88]. unrestricted [Par82a]. upon [GWK89]. usage [BW87, HT84]. use [AL84, CC80, CCSH84, CCH84a, Ful80, Gra81, Har82, KM84, MaC82, Sta83, Tis83, Zam84]. Useful [Bre80]. Usefulness [Ter87a]. user [GLS87]. Using [BM89, AH82, HH84, HK86, HNS88, HW81, Kin83a, NN85a, PN84, PN82, PZ85, Poi80a, Roe84, Sch88, SW89a, ZBV88, ZH89, vDK80]. utility [Bou89]. utilization [BF86b, ED80, HC86, Hul86, Mor86].

Valencia [Ano81s]. validity [Kul85]. value [Mas85]. valued [AH89]. values [AMS83]. var [Lüt89, Nic85]. variable [Ame85, And87, And89, Bar83, BR85, Ben83, BK85, CL85, GS81, Gre83a, HC86, Lee82a, Lev86, MP82, Mut83, New86, New87a, OS80, Phi84a, Ruu86, Smi87, Xie88, Xie11]. variables [AD89, Ame85, Bai87, BH88, Blo89, BJY84, Cam86, Con85, DA89, Dia82, Don82, Don83, FG83, Ful80, GS88, GH86, Hsi89, Kle88b, Kle88c, KP87, Lüt82, New86, New87a, PN84, Pal80, SJ86, Sta84, Ter87a, Tse82]. variance [CGW87, Cow89, Cra87, Duf88, GD88, Kan85, LM89, MT84, Sha85]. variances [CM89, Oht87, TO86]. variations [BF86b]. various [BH82]. varying [LH81, Mor86, Pag80, WE83]. vector [AH89, BM80, Ham86, Lüt84, Oha88]. vector-valued [AH89]. Ven [Ano83c]. versus [JS85, Lo86, MI80, Wes86]. VI [Ano83i]. via [Mit85, New88]. victimization [Kad85]. Vienna [Ano83i]. Vintage [AT87, HW81]. Volumes [Ano82s, Ano85z, Ano89w].

wage [Kie82]. wages [WO81]. Wald [GHM81]. Wallace [SW80]. Watson [DD85, Kin81, Kin83a, Krä85]. way [Bal81a, Xie88, Xie11]. weak [Bel82]. wealth [DS89, Sat85]. wedge [BHW86]. Weilbull [Lan85a]. welfare [Aig84, GLS87, PW84, Ste89]. well [SBN89]. well-known [SBN89]. wheat [ASL87]. where [OS80]. which [DG83, Mis83, NN85a]. whites [WO81]. wholesale [CL82, SW80]. without [Krä85, Zam81]. women [NN85a, Ste89]. work [Slo89, Ste89]. workers [Cha83]. workshop [Ano89l]. world [Ano81-32]. Wu [NN85b]. Wynand [Ano83c].

young [Ano86e, Ano88e]. youth [Lyn85]. Yule [Xek84].

zero [Che85, LC86, PT84].

References

Aigner:1989:LVM

[AD89] Dennis J. Aigner and Manfred Deistler. Latent variables models: Editors’ introduction. Journal of Econometrics, 41(1):1–3, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

            (electronic). URL http://www.sciencedirect.com/science/
            article/pii/0304407689900390.

Ali:1984:SSN

[AG84] Mukhtar M. Ali and Carmelo Giaccotto. A study of several new and existing tests for heteroscedasticity in the general linear model *. Journal of Econometrics, 26(3):355–373, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900265.

Agbeyegbe:1988:EDA

[Agb88] Terence D. Agbeyegbe. An exact discrete analog of an open linear non-stationary first-order continuous-time system with mixed sample. Journal of Econometrics, 39(3):237–250, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900577.

Anderson:1982:FED

[AH82] T. W. Anderson and Cheng Hsiao. Formulation and estimation of dynamic models using panel data. Journal of Econometrics, 18(1): 47–82, January 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900951.

Aoki:1989:MAR

[AH89] Masanao Aoki and Arthur Havenner. A method for approximate representation of vector-valued time series and its relation to two alternatives. Journal of Econometrics, 42(2):181–199, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768990002X.

Aigner:1984:WEP

[Aig84] Dennis J. Aigner. The welfare econometrics of peak-load pricing for electricity: Editor’s introduction. Journal of Econometrics, 26(1–2):1–15, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900101.

Anderson:1983:NLR

[AJR83] Richard G. Anderson, James M. Johannes, and Robert H. Rasche. A new look at the relationship between time-series and structural econometric models. Journal of Econometrics, 23(2):235– 251, October 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390079K.

Akaike:1981:LMI

[Aka81] Hirotugu Akaike. Likelihood of a model and information criteria. Journal of Econometrics, 16(1):3–14, May 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900713.

Aigner:1984:ETU

[AL84] Dennis J. Aigner and Edward E. Leamer. Estimation of timeof-use pricing response in the absence of experimental data: An application of the methodology of data transferability. Journal of Econometrics, 26(1–2):205–227, September/October 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900186.

Amemiya:1983:CAG

[Ame83a] Takeshi Amemiya. A comparison of the Amemiya GLS and the Lee–Maddala–Trost G2SLS in a simultaneous-equations Tobit model. Journal of Econometrics, 23(3):295–300, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900581.

Amemiya:1983:PGL

[Ame83b] Takeshi Amemiya. Partially generalized least squares and two-stage least squares estimators. Journal of Econometrics, 23(2):275–283, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390082G.

Amemiya:1984:EI

[Ame84a] Takeshi Amemiya. Editor’s introduction. Journal of Econometrics, 24(1–2):1–2, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0304407684900733.

Amemiya:1984:TMS

[Ame84b] Takeshi Amemiya. Tobit models: A survey. Journal of Econometrics, 24(1–2):3–61, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900745.

Amemiya:1985:IVE

[Ame85] Yasuo Amemiya. Instrumental variable estimator for the nonlinear errors-in-variables model. Journal of Econometrics, 28(3): 273–289, June 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900016.

Anderson:1983:NVS

[AMS83] T. W. Anderson, Kimio Morimune, and Takamitsu Sawa. The numerical values of some key parameters in econometric models. Journal of Econometrics, 21(2):229–243, February 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900155.

Ansley:1980:FSP

[AN80] Craig F. Ansley and Paul Newbold. Finite sample properties of estimators for autoregressive moving average models. Journal of Econometrics, 13(2):159–183, June 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900135.

Anderson:1980:SSE

[And80] G. J. Anderson. The structure of simultaneous equations estimators: A comment. Journal of Econometrics, 14(2):271–276, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900974.

Andersen:1983:LTM

[And83] Erling B. Andersen. Latent trait models. Journal of Econometrics, 22(1–2):215–227, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901008.

Anderson:1987:PTL

[And87] G. J. Anderson. Prediction tests in limited dependent variable models. Journal of Econometrics, 34(1–2):253–261, January/ February 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900741.

Andrews:1988:CSD

[And88] Donald W. K. Andrews. Chi-square diagnostic tests for econometric models: Introduction and applications. Journal of Econometrics, 37(1):135–156, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900796.

Anderson:1989:LLV

[And89] T. W. Anderson. Linear latent variable models and covariance structures. Journal of Econometrics, 41(1):91–119, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900444.

Anonymous:1980:Aa

[Ano80a] Anonymous. Acknowledgement. Journal of Econometrics, 12(2):247–248, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900111.

Anonymous:1980:Af

[Ano80b] Anonymous. Acknowledgement. Journal of Econometrics, 14(3):397–398, December 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900366.

Anonymous:1980:Ab

[Ano80c] Anonymous. Announcement. Journal of Econometrics, 12(3): 249, April 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768090055X.

Anonymous:1980:Ac

[Ano80d] Anonymous. Announcement. Journal of Econometrics, 13(2): 267–268, June 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900184.

Anonymous:1980:Ae

[Ano80e] Anonymous. Announcement. Journal of Econometrics, 14(3): 285, December 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900287.

Anonymous:1980:Ad

[Ano80f] Anonymous. Announcements. Journal of Econometrics, 14(2): 283, October 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680901001.

Anonymous:1980:BRA

[Ano80g] Anonymous. Book review: Annals of Applied Econometrics 1980-3. Journal of Econometrics, 14(1):159, September 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900792.

Anonymous:1980:CCP

[Ano80h] Anonymous. Combined contents-parts i and II. Journal of Econometrics, 14(1):1–2, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900688.

Anonymous:1980:EBa

[Ano80i] Anonymous. Editorial Board. Journal of Econometrics, 12(1): ifc, January 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900470.

Anonymous:1980:EBb

[Ano80j] Anonymous. Editorial board. Journal of Econometrics, 13(1):ii, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768090038X.

Anonymous:1980:EBc

[Ano80k] Anonymous. Editorial Board. Journal of Econometrics, 14(1): ifc, September 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900676.

Anonymous:1980:Ea

[Ano80l] Anonymous. Erratum. Journal of Econometrics, 12(3):393, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900652.

Anonymous:1980:Ec

[Ano80m] Anonymous. Erratum. Journal of Econometrics, 14(2):281, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900998.

Anonymous:1980:Ia

[Ano80n] Anonymous. Index. Journal of Econometrics, 12(3):395–396, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900664.

Anonymous:1980:Ib

[Ano80o] Anonymous. Index. Journal of Econometrics, 13(3):403–404, August 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900871.

Anonymous:1980:Ic

[Ano80p] Anonymous. Index. Journal of Econometrics, 14(3):399–400, December 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900378.

Anonymous:1980:PJa

[Ano80q] Anonymous. Pages 1–102 (January 1980). Journal of Econometrics, 12(1):??, January 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PM

[Ano80r] Anonymous. Pages 1–138 (May 1980). Journal of Econometrics, 13(1):??, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PS

[Ano80s] Anonymous. Pages 1–159 (September 1980). Journal of Econometrics, 14(1):??, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PF

[Ano80t] Anonymous. Pages 103–248 (February 1980). Journal of Econometrics, 12(2):??, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PJb

[Ano80u] Anonymous. Pages 139–268 (June 1980). Journal of Econometrics, 13(2):??, June 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PO

[Ano80v] Anonymous. Pages 161–283 (October 1980). Journal of Econometrics, 14(2):??, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PAa

[Ano80w] Anonymous. Pages 249–396 (April 1980). Journal of Econometrics, 12(3):??, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PAb

[Ano80x] Anonymous. Pages 269–404 (August 1980). Journal of Econometrics, 13(3):??, August 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:PD

[Ano80y] Anonymous. Pages 285–400 (December 1980). Journal of Econometrics, 14(3):??, December 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1980:TSA

[Ano80z] Anonymous. Time series analysis and forecasting (TSA&F) special interest group. Journal of Econometrics, 14(3):395, December 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900354.

Anonymous:1981:SME

[Ano81a] Anonymous. 1982 Summer meeting of the Econometric Society: A call for papers. Journal of Econometrics, 17(2):259–260, November 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900300.

Anonymous:1981:NAT

[Ano81b] Anonymous. 1983 North American Time Series meetings in Toronto: Initial notice. Journal of Econometrics, 17(3):396, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900130.

Anonymous:1981:ATS

[Ano81c] Anonymous. 3rd American Time Series Meeting (8th ITSM): Cincinnati (Ohio), 19–21 August 1982. Journal of Econometrics, 17(3):395–396, December 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681900129.

Anonymous:1981:Ac

[Ano81d] Anonymous. Acknowledgement. Journal of Econometrics, 17(3):397–398, December 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681900142.

Anonymous:1981:Aa

[Ano81e] Anonymous. Announcement. Journal of Econometrics, 16(2): 275, June 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901135.

Anonymous:1981:Ab

[Ano81f] Anonymous. Announcement. Journal of Econometrics, 17(3): 261, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900014.

Anonymous:1981:BRA

[Ano81g] Anonymous. Book review: Annals of Applied Econometrics 1981-1. Journal of Econometrics, 15(1):175, January 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901354.

Anonymous:1981:CCP

[Ano81h] Anonymous. Combined contents-parts i and II. Journal of Econometrics, 15(1):1–2, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901238.

Anonymous:1981:C

[Ano81i] Anonymous. Corrigenda. Journal of Econometrics, 15(3):405, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901032.

Anonymous:1981:EBa

[Ano81j] Anonymous. Editorial Board. Journal of Econometrics, 15(1): iv, January 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901226.

Anonymous:1981:EBb

[Ano81k] Anonymous. Editorial Board. Journal of Econometrics, 16(1):ifc, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900695.

Anonymous:1981:EBc

[Ano81l] Anonymous. Editorial Board. Journal of Econometrics, 17(1): ifc, September 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900555.

Anonymous:1981:EIP

[Ano81m] Anonymous. Editor’s introduction to part II. Journal of Econometrics, 15(1):3–9, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768190124X.

Anonymous:1981:Ea

[Ano81n] Anonymous. Erratum. Journal of Econometrics, 15(2):309, February 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901202.

Anonymous:1981:Ec

[Ano81o] Anonymous. Erratum. Journal of Econometrics, 17(3):393–394, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900105.

Anonymous:1981:Ia

[Ano81p] Anonymous. Index. Journal of Econometrics, 15(3):407–408, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901056.

Anonymous:1981:Ib

[Ano81q] Anonymous. Index. Journal of Econometrics, 16(3):401, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900385.

Anonymous:1981:Ic

[Ano81r] Anonymous. Index. Journal of Econometrics, 17(3):399–400, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900154.

Anonymous:1981:IFC

[Ano81s] Anonymous. International Forecasting Conference (IFC): Valencia (Spain), 24–28 May 1982. Journal of Econometrics, 17(1): 139, September 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900683.

Anonymous:1981:LJS

[Ano81t] Anonymous. Leonard J. Savage award. Journal of Econometrics, 17(3):395, December 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681900117.

Anonymous:1981:PS

[Ano81u] Anonymous. Pages 1–139 (September 1981). Journal of Econometrics, 17(1):??, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PM

[Ano81v] Anonymous. Pages 1–170 (May 1981). Journal of Econometrics, 16(1):??, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PJa

[Ano81w] Anonymous. Pages 1–175 (January 1981). Journal of Econometrics, 15(1):??, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PN

[Ano81x] Anonymous. Pages 141–260 (November 1981). Journal of Econometrics, 17(2):??, November 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PJb

[Ano81y] Anonymous. Pages 171–275 (June 1981). Journal of Econometrics, 16(2):??, June 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PF

[Ano81z] Anonymous. Pages 177–309 (February 1981). Journal of Econometrics, 15(2):??, February 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PD

[Ano81-27] Anonymous. Pages 261–400 (December 1981). Journal of Econometrics, 17(3):??, December 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PAb

[Ano81-28] Anonymous. Pages 277–402 (August 1981). Journal of Econometrics, 16(3):??, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PAa

[Ano81-29] Anonymous. Pages 311–408 (April 1981). Journal of Econometrics, 15(3):??, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1981:PCM

[Ano81-30] Anonymous. Program of the conference on model selection: April 18–21, 1980. Journal of Econometrics, 16(1):169–170, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/030440768190097X.

Anonymous:1981:SIT

[Ano81-31] Anonymous. Seventh International Time Series meeting (ITSM): Dublin (Ireland), 15–19 March 1982. Journal of Econometrics, 17(1):139, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681900671.

Anonymous:1981:SWC

[Ano81-32] Anonymous. Seventh world congress of the international economic association in 1983. Journal of Econometrics, 15(3):406, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901044.

Anonymous:1982:Ab

[Ano82a] Anonymous. Acknowledgement. Journal of Econometrics, 20(2):337–338, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900252.

Anonymous:1982:Aa

[Ano82b] Anonymous. Announcement. Journal of Econometrics, 20(2): 159, November 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768290015X.

Anonymous:1982:Ea

[Ano82c] Anonymous. Editorial. Journal of Econometrics, 20(2):161, November 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900161.

Anonymous:1982:EBa

[Ano82d] Anonymous. Editorial Board. Journal of Econometrics, 18(1): ii, January 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900926.

Anonymous:1982:EBb

[Ano82e] Anonymous. Editorial board. Journal of Econometrics, 19(1): CO2, May 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768290046X.

Anonymous:1982:EBc

[Ano82f] Anonymous. Editorial Board. Journal of Econometrics, 20(1): ii, October 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900999.

Anonymous:1982:Ec

[Ano82g] Anonymous. Erratum. Journal of Econometrics, 20(2):335, November 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900240.

Anonymous:1982:Ia

[Ano82h] Anonymous. Index. Journal of Econometrics, 18(3):403, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900914.

Anonymous:1982:Ib

[Ano82i] Anonymous. Index. Journal of Econometrics, 19(2–3):391, August 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900148.

Anonymous:1982:Ic

[Ano82j] Anonymous. Index. Journal of Econometrics, 20(2):339, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900264.

Anonymous:1982:PO

[Ano82k] Anonymous. Pages 1–157 (October 1982). Journal of Econometrics, 20(1):??, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PJ

[Ano82l] Anonymous. Pages 1–190 (January 1982). Journal of Econometrics, 18(1):??, January 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PN

[Ano82m] Anonymous. Pages 159–339 (November 1982). Journal of Econometrics, 20(2):??, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PAb

[Ano82n] Anonymous. Pages 183–391 (August 1982). Journal of Econometrics, 19(2–3):??, August 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PF

[Ano82o] Anonymous. Pages 191–294 (February 1982). Journal of Econometrics, 18(2):??, February 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PAa

[Ano82p] Anonymous. Pages 295–403 (April 1982). Journal of Econometrics, 18(3):??, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PD

[Ano82q] Anonymous. Pages 341–413 (December 1982). Journal of Econometrics, 20(3):??, December 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:PCM

[Ano82r] Anonymous. Pages CO2, 1–182 (May 1982). Journal of Econometrics, 19(1):??, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1982:SIV

[Ano82s] Anonymous. Subject index: Volumes 11–20, 1979–1982. Journal of Econometrics, 20(3):341–413, December 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900458.

Anonymous:1983:Ab

[Ano83a] Anonymous. Acknowledgement. Journal of Econometrics, 23(3):409–411, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683900696.

Anonymous:1983:Aa

[Ano83b] Anonymous. Announcement. Journal of Econometrics, 23(3): 293, December 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768390057X.

Anonymous:1983:CDD

[Ano83c] Anonymous. Corrigendum: “The demand for deductibles in private health insurance: A probit model with sample selection,” Journal of Econometrics: Wynand P. M. M. van de Ven and Bernard M. S. van Praag 17 (1981) pp. 229– 252. Journal of Econometrics, 22(3):395, August 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901124. See [VV81].

Anonymous:1983:EBa

[Ano83d] Anonymous. Editorial Board. Journal of Econometrics, 21(1): ii, January 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768390115X.

Anonymous:1983:EBb

[Ano83e] Anonymous. Editorial Board. Journal of Econometrics, 22(1–2): ifc, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900908.

Anonymous:1983:EI

[Ano83f] Anonymous. Editors’ introduction. Journal of Econometrics, 22(1–2):1–12, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768390091X.

Anonymous:1983:Ec

[Ano83g] Anonymous. Errata. Journal of Econometrics, 23(3):407, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900684.

Anonymous:1983:Ea

[Ano83h] Anonymous. Erratum. Journal of Econometrics, 21(3):403, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900544.

Anonymous:1983:EVS

[Ano83i]  Anonymous.  Euro VI — Sixth European Congress on

Operational Research: Vienna, Austria, 19–22 July 1983. Journal of Econometrics, 21(3):404, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900556.

Anonymous:1983:Ia

[Ano83j] Anonymous. Index. Journal of Econometrics, 21(3):405–406, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900568.

Anonymous:1983:Ib

[Ano83k] Anonymous. Index. Journal of Econometrics, 22(3):399, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683901148.

Anonymous:1983:Ic

[Ano83l] Anonymous. Index. Journal of Econometrics, 23(3):413–414, December 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900702.

Anonymous:1983:NI

[Ano83m] Anonymous. News items. Journal of Econometrics, 22(3):397– 398, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901136.

Anonymous:1983:PJ

[Ano83n] Anonymous. Pages 1–160 (January 1983). Journal of Econometrics, 21(1):??, January 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PMJ

[Ano83o] Anonymous. Pages 1–243 (May–June 1983). Journal of Econometrics, 22(1–2):??, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PF

[Ano83p] Anonymous. Pages 161–261 (February 1983). Journal of Econometrics, 21(2):??, February 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PO

[Ano83q] Anonymous. Pages 165–292 (October 1983). Journal of Econometrics, 23(2):??, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PAb

[Ano83r] Anonymous. Pages 245–399 (August 1983). Journal of Econometrics, 22(3):??, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PAa

[Ano83s] Anonymous. Pages 263–406 (April 1983). Journal of Econometrics, 21(3):??, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:PD

[Ano83t] Anonymous. Pages 293–414 (December 1983). Journal of Econometrics, 23(3):??, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1983:SIT

[Ano83u] Anonymous. Schedule of International Time Series Meetings (ITSM), 1983–1985. Journal of Econometrics, 21(2):261, February 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900192.

Anonymous:1984:Aa

[Ano84a] Anonymous. Announcement. Journal of Econometrics, 25(3): 395, July 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900083.

Anonymous:1984:Ab

[Ano84b] Anonymous. Announcement. Journal of Econometrics, 26(3): 253, December 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900204.

Anonymous:1984:EBa

[Ano84c] Anonymous. Editorial Board. Journal of Econometrics, 24 (1–2):iv, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407684900721.

Anonymous:1984:EBb

[Ano84d] Anonymous. Editorial Board. Journal of Econometrics, 25(1–2):CO2, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407684900307.

Anonymous:1984:Ia

[Ano84e] Anonymous. Index. Journal of Econometrics, 24(3):405, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900629.

Anonymous:1984:Ib

[Ano84f] Anonymous. Index. Journal of Econometrics, 25(3):397–398, July 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900095.

Anonymous:1984:Ic

[Ano84g] Anonymous. Index. Journal of Econometrics, 26(3):399, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900290.

Anonymous:1984:PJF

[Ano84h] Anonymous. Pages 1–222 (January–February 1984). Journal of Econometrics, 24(1–2):??, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1984:PMJ

[Ano84i] Anonymous. Pages 1–239 (May–June 1984). Journal of Econometrics, 25(1–2):??, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1984:PSO

[Ano84j] Anonymous. Pages 1–252 (September–October 1984). Journal of Econometrics, 26(1–2):??, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1984:PM

[Ano84k] Anonymous. Pages 223–406 (March 1984). Journal of Econometrics, 24(3):??, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1984:PJ

[Ano84l] Anonymous. Pages 241–398 (July 1984). Journal of Econometrics, 25(3):??, July 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1984:PD

[Ano84m] Anonymous. Pages 253–399 (December 1984). Journal of Econometrics, 26(3):??, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:Aa

[Ano85a] Anonymous. Acknowledgement. Journal of Econometrics, 27(1):139–140, January 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900508.

Anonymous:1985:Ab

[Ano85b] Anonymous. Acknowledgement. Journal of Econometrics, 27(3): 387, March 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900144.

Anonymous:1985:Ac

[Ano85c] Anonymous. Acknowledgement. Journal of Econometrics, 29(3):351–352, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901617.

Anonymous:1985:C

[Ano85d] Anonymous. Corrigendum. Journal of Econometrics, 28(2):269, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901253.

Anonymous:1985:DOR

[Ano85e] Anonymous. A decade outlook on research opportunities in the behavioral and social sciences. Journal of Econometrics, 27(1): 141, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768590051X.

Anonymous:1985:EBa

[Ano85f] Anonymous. Editorial Board. Journal of Econometrics, 27(1): ifc, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900405.

Anonymous:1985:EBb

[Ano85g] Anonymous. Editorial board. Journal of Econometrics, 28(1): ifc, April 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900624.

Anonymous:1985:EBc

[Ano85h] Anonymous. Editorial Board. Journal of Econometrics, 29(1–2): ii, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900284.

Anonymous:1985:EBd

[Ano85i] Anonymous. Editorial Board. Journal of Econometrics, 30 (1–2):ii, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901277.

Anonymous:1985:E

[Ano85j] Anonymous. Erratum. Journal of Econometrics, 27(2):271, February 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900922.

Anonymous:1985:Ia

[Ano85k] Anonymous. Index. Journal of Econometrics, 27(3):389, March 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900156.

Anonymous:1985:Ib

[Ano85l] Anonymous. Index. Journal of Econometrics, 28(3):371, June 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900065.

Anonymous:1985:Ic

[Ano85m] Anonymous. Index. Journal of Econometrics, 29(3):353, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901629.

Anonymous:1985:Id

[Ano85n] Anonymous. Index. Journal of Econometrics, 30(1–2):473– 474, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901526.

Anonymous:1985:NI

[Ano85o] Anonymous. News item. Journal of Econometrics, 28(2):271, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901265.

Anonymous:1985:PJa

[Ano85p] Anonymous. Pages 1–141 (January 1985). Journal of Econometrics, 27(1):??, January 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PA

[Ano85q] Anonymous. Pages 1–169 (April 1985). Journal of Econometrics, 28(1):??, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PJAa

[Ano85r] Anonymous. Pages 1–211 (July–August 1985). Journal of Econometrics, 29(1–2):??, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PON

[Ano85s] Anonymous. Pages 1–474 (October–November 1985). Journal of Econometrics, 30(1–2):??, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PF

[Ano85t] Anonymous. Pages 143–271 (February 1985). Journal of Econometrics, 27(2):??, February 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PMb

[Ano85u] Anonymous. Pages 171–271 (May 1985). Journal of Econometrics, 28(2):??, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PS

[Ano85v] Anonymous. Pages 213–353 (September 1985). Journal of Econometrics, 29(3):??, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PJb

[Ano85w] Anonymous. Pages 273–371 (June 1985). Journal of Econometrics, 28(3):??, June 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PMa

[Ano85x] Anonymous. Pages 273–389 (March 1985). Journal of Econometrics, 27(3):??, March 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:PD

[Ano85y] Anonymous. Pages 475–543 (December 1985). Journal of Econometrics, 30(3):??, December 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1985:SIV

[Ano85z] Anonymous. Subject index: Volumes 21–30, 1983–1985. Journal of Econometrics, 30(3):475–543, December 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900272.

Anonymous:1986:EMP

[Ano86a] Anonymous. 1987 European meeting of the Psychometric Society. Journal of Econometrics, 33(3):399, December 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900059.

Anonymous:1986:Ac

[Ano86b] Anonymous. Acknowledgement. Journal of Econometrics, 33(3):400–402, December 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407686900060.

Anonymous:1986:Aa

[Ano86c] Anonymous. Announcement. Journal of Econometrics, 31(1): 1, February 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900527.

Anonymous:1986:Ab

[Ano86d] Anonymous. Announcement. Journal of Econometrics, 32(2): 295, July 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900448.

Anonymous:1986:CYS

[Ano86e] Anonymous. Competitions for young statisticians from developing countries: 1987 and 1989. Journal of Econometrics, 31(2): 233, March 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900503.

Anonymous:1986:C

[Ano86f] Anonymous. Correction. Journal of Econometrics, 32(2):293, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900436.

Anonymous:1986:EBa

[Ano86g] Anonymous. Editorial Board. Journal of Econometrics, 31(1): ii–iii, February 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900515.

Anonymous:1986:EBb

[Ano86h] Anonymous. Editorial board. Journal of Econometrics, 32(1):ii, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900084.

Anonymous:1986:EBc

[Ano86i] Anonymous. Editorial board. Journal of Econometrics, 33 (1–2):ii, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407686900230.

Anonymous:1986:Ia

[Ano86j] Anonymous. Index. Journal of Econometrics, 31(3):363, April 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900667.

Anonymous:1986:Ib

[Ano86k] Anonymous. Index. Journal of Econometrics, 32(3):399, August 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900229.

Anonymous:1986:Ic

[Ano86l] Anonymous. Index. Journal of Econometrics, 33(3):403, December 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900072.

Anonymous:1986:ISO

[Ano86m] Anonymous. ISI section on official statistics. Journal of Econometrics, 31(1):119, February 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768690059X.

Anonymous:1986:PF

[Ano86n] Anonymous. Pages 1–119 (February 1986). Journal of Econometrics, 31(1):??, February 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PJa

[Ano86o] Anonymous. Pages 1–187 (June 1986). Journal of Econometrics, 32(1):??, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PON

[Ano86p] Anonymous. Pages 1–310 (October–November 1986). Journal of Econometrics, 33(1–2):??, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PM

[Ano86q] Anonymous. Pages 121–233 (March 1986). Journal of Econometrics, 31(2):??, March 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PJb

[Ano86r] Anonymous. Pages 189–295 (July 1986). Journal of Econometrics, 32(2):??, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PAa

[Ano86s] Anonymous. Pages 235–363 (April 1986). Journal of Econometrics, 31(3):??, April 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PAb

[Ano86t] Anonymous. Pages 297–399 (August 1986). Journal of Econometrics, 32(3):??, August 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1986:PD

[Ano86u] Anonymous. Pages 311–403 (December 1986). Journal of Econometrics, 33(3):??, December 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:Ab

[Ano87a] Anonymous. Acknowledgement. Journal of Econometrics, 36(3):391–393, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768790011X.

Anonymous:1987:Aa

[Ano87b] Anonymous. Announcement. Journal of Econometrics, 35(2– 3):397, July 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900376.

Anonymous:1987:Ca

[Ano87c] Anonymous. Corrigenda. Journal of Econometrics, 35(2–3):393, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900352.

Anonymous:1987:Cb

[Ano87d] Anonymous. Corrigenda. Journal of Econometrics, 35(2–3):395, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900364.

Anonymous:1987:EBa

[Ano87e] Anonymous. Editorial Board. Journal of Econometrics, 34 (1–2):ifc, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900637.

Anonymous:1987:EBb

[Ano87f] Anonymous. Editorial Board. Journal of Econometrics, 35(1):ii, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900765.

Anonymous:1987:EBc

[Ano87g] Anonymous. Editorial board. Journal of Econometrics, 36 (1–2):ii, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768790039X.

Anonymous:1987:EI

[Ano87h] Anonymous. Editor’s introduction. Journal of Econometrics, 35 (1):1, May 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900777.

Anonymous:1987:Ia

[Ano87i] Anonymous. Index. Journal of Econometrics, 34(3):393, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900224.

Anonymous:1987:Ib

[Ano87j] Anonymous. Index. Journal of Econometrics, 35(2–3):399, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900388.

Anonymous:1987:Ic

[Ano87k] Anonymous. Index. Journal of Econometrics, 36(3):395–396, November 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900121.

Anonymous:1987:PMb

[Ano87l] Anonymous. Pages 1–190 (May 1987). Journal of Econometrics, 35(1):??, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:PSO

[Ano87m] Anonymous. Pages 1–230 (September–October 1987). Journal of Econometrics, 36(1–2):??, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:PJF

[Ano87n] Anonymous. Pages 1–274 (January–February 1987). Journal of Econometrics, 34(1–2):??, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:PJ

[Ano87o] Anonymous. Pages 191–399 (July 1987). Journal of Econometrics, 35(2–3):??, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:PN

[Ano87p] Anonymous. Pages 231–396 (November 1987). Journal of Econometrics, 36(3):??, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1987:PMa

[Ano87q] Anonymous. Pages 275–393 (March 1987). Journal of Econometrics, 34(3):??, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:Ad

[Ano88a] Anonymous. Acknowledgement. Journal of Econometrics, 39(3):399–400, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768890067X.

Anonymous:1988:Aa

[Ano88b] Anonymous. Announcement. Journal of Econometrics, 37(2): 193, February 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900012.

Anonymous:1988:Ab

[Ano88c] Anonymous. Announcement. Journal of Econometrics, 38(3): 401, July 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900541.

Anonymous:1988:Ac

[Ano88d] Anonymous. Announcement. Journal of Econometrics, 39(3): 235, November 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900565.

Anonymous:1988:CYS

[Ano88e] Anonymous. Competition for young statisticians from developing countries. Journal of Econometrics, 37(3):396, March 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900152.

Anonymous:1988:D

[Ano88f] Anonymous. Dedication. Journal of Econometrics, 38(1–2):1, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900231.

Anonymous:1988:EBa

[Ano88g] Anonymous. Editorial Board. Journal of Econometrics, 37(1): ii, January 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900693.

Anonymous:1988:EBb

[Ano88h] Anonymous. Editorial board. Journal of Econometrics, 38(1–2): ifc, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768890022X.

Anonymous:1988:EBc

[Ano88i] Anonymous. Editorial Board. Journal of Econometrics, 39 (1–2):ifc, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768890036X.

Anonymous:1988:Ea

[Ano88j] Anonymous. Erratum. Journal of Econometrics, 37(3):395, March 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900140.

Anonymous:1988:Ec

[Ano88k] Anonymous. Erratum. Journal of Econometrics, 39(3):397, November 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900668.

Anonymous:1988:Ia

[Ano88l] Anonymous. Index. Journal of Econometrics, 37(3):397, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900164.

Anonymous:1988:Ib

[Ano88m] Anonymous. Index. Journal of Econometrics, 38(3):403, July 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900553.

Anonymous:1988:Ic

[Ano88n] Anonymous. Index. Journal of Econometrics, 39(3):401, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900681.

Anonymous:1988:PJa

[Ano88o] Anonymous. Pages 1–192 (January 1988). Journal of Econometrics, 37(1):??, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PSO

[Ano88p] Anonymous. Pages 1–234 (September–October 1988). Journal of Econometrics, 39(1–2):??, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PMJ

[Ano88q] Anonymous. Pages 1–268 (May–June 1988). Journal of Econometrics, 38(1–2):??, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PF

[Ano88r] Anonymous. Pages 193–292 (February 1988). Journal of Econometrics, 37(2):??, February 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PN

[Ano88s] Anonymous. Pages 235–401 (November 1988). Journal of Econometrics, 39(3):??, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PJb

[Ano88t] Anonymous. Pages 269–403 (July 1988). Journal of Econometrics, 38(3):??, July 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1988:PM

[Ano88u] Anonymous. Pages 293–397 (March 1988). Journal of Econometrics, 37(3):??, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:A

[Ano89a] Anonymous. Announcement. Journal of Econometrics, 41(2): 187, June 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900924.

Anonymous:1989:Ca

[Ano89b] Anonymous. Corrigendum. Journal of Econometrics, 40(2):339, February 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900900.

Anonymous:1989:Cb

[Ano89c] Anonymous. Corrigendum. Journal of Econometrics, 41(2):279– 281, June 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900985.

Anonymous:1989:EBa

[Ano89d] Anonymous. Editorial board. Journal of Econometrics, 40(1): ifc, January 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900249.

Anonymous:1989:EBb

[Ano89e] Anonymous. Editorial board. Journal of Econometrics, 41(1):ii, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900389.

Anonymous:1989:EBc

[Ano89f] Anonymous. Editorial Board. Journal of Econometrics, 42(1): ii, September 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900699.

Anonymous:1989:Eb

[Ano89g] Anonymous. Errata. Journal of Econometrics, 42(3):381, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407689900614.

Anonymous:1989:Ea

[Ano89h] Anonymous. Erratum. Journal of Econometrics, 42(2):285, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900080.

Anonymous:1989:Ia

[Ano89i] Anonymous. Index. Journal of Econometrics, 40(2):340, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900912.

Anonymous:1989:Ib

[Ano89j] Anonymous. Index. Journal of Econometrics, 41(3):385, July 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900687.

Anonymous:1989:I

[Ano89k] Anonymous. Index. Journal of Econometrics, 42(3):383–384, November 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0304407689900626.

Anonymous:1989:LIS

[Ano89l] Anonymous. Luxembourg income study summer workshop. Journal of Econometrics, 41(2):283, June 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900997.

Anonymous:1989:NI

[Ano89m] Anonymous. News item. Journal of Econometrics, 42(2):286, October 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900092.

Anonymous:1989:PS

[Ano89n] Anonymous. Pages 1–155 (September 1989). Journal of Econometrics, 42(1):??, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PMb

[Ano89o] Anonymous. Pages 1–185 (May 1989). Journal of Econometrics, 41(1):??, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PJa

[Ano89p] Anonymous. Pages 1–202 (January 1989). Journal of Econometrics, 40(1):??, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PO

[Ano89q] Anonymous. Pages 157–286 (October 1989). Journal of Econometrics, 42(2):??, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PJb

[Ano89r] Anonymous. Pages 187–283 (June 1989). Journal of Econometrics, 41(2):??, June 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PF

[Ano89s] Anonymous. Pages 203–340 (February 1989). Journal of Econometrics, 40(2):??, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PJc

[Ano89t] Anonymous. Pages 285–385 (July 1989). Journal of Econometrics, 41(3):??, July 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PN

[Ano89u] Anonymous. Pages 287–384 (November 1989). Journal of Econometrics, 42(3):??, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:PMa

[Ano89v] Anonymous. Pages 341–460 (March 1989). Journal of Econometrics, 40(3):??, March 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1989:SAI

[Ano89w] Anonymous. Subject and author index: Volumes 31–40, 1986– 1989. Journal of Econometrics, 40(3):341–460, March 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900377.

Antle:1986:AEE

[Ant86] John M. Antle. Aggregation, expectations, and the explanation of technological change. Journal of Econometrics, 33(1–2):213–236, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900345.

Amemiya:1981:CBC

[AP81] Takeshi Amemiya and James L. Powell. A comparison of the Box– Cox maximum likelihood estimator and the non-linear two-stage least squares estimator. Journal of Econometrics, 17(3):351–381, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900075.

Arnold:1983:BIP

[AP83] Barry C. Arnold and S. James Press. Bayesian inference for Pareto populations. Journal of Econometrics, 21(3):287–306, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900477.

Appelbaum:1982:EDO

[App82] Elie Appelbaum. The estimation of the degree of oligopoly power. Journal of Econometrics, 19(2–3):287–299, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900069.

Arellano:1989:EGE

[Are89a] Manuel Arellano. An efficient GLS estimator of triangular models with covariance restrictions. Journal of Econometrics, 42(2):267– 273, October 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900067.

Arellano:1989:EES

[Are89b] Manuel Arellano. On the efficient estimation of simultaneous equations with covariance restrictions. Journal of Econometrics, 42(2):247–265, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900055.

Aigner:1980:EI

[AS80] Dennis J. Aigner and Peter Schmidt. Editors’ introduction. Journal of Econometrics, 13(1):1–3, May 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900391.

Arabmazar:1981:FER

[AS81] Abbas Arabmazar and Peter Schmidt. Further evidence on the robustness of the Tobit estimator to heteroskedasticity. Journal of Econometrics, 17(2):253–258, November 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900294.

Ahmad:1987:DWU

[ASL87] Ehtisham Ahmad, Nicholas Stern, and H.-M. Leung. The demand for wheat under non-linear pricing in Pakistan. Journal of Econometrics, 36(1–2):55–65, September/October 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900431.

Akiyama:1987:VPA

[AT87] T. Akiyama and P. K. Trivedi. Vintage production approach to perennial crop supply: An application to tea in major producing countries. Journal of Econometrics, 36(1–2):133–161, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900479.

Aguirre-Torres:1983:NNN

[ATG83] Victor Aguirre-Torres and A. Ronald Gallant. The null and nonnull asymptotic distribution of the Cox test for multivariate nonlinear regression: Alternatives and a new distribution-free Cox test. Journal of Econometrics, 21(1):5–33, January 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901173.

Atkinson:1981:LRP

[Atk81] A. C. Atkinson. Likelihood ratios, posterior odds and information criteria. Journal of Econometrics, 16(1):15–20, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900725.

Attfield:1985:HES

[Att85] Clifford L. F. Attfield. Homogeneity and endogeneity in systems of demand equations. Journal of Econometrics, 27(2):197– 209, February 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900879.

Aigner:1988:EI

[AZ88] Dennis J. Aigner and Arnold Zellner. Editors’ introduction. Journal of Econometrics, 39(1–2):1–5, September/October 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900371.

Baillie:1980:PAM

[Bai80] Richard T. Baillie. Predictions from ARMAX models. Journal of Econometrics, 12(3):365–374, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900627.

Baillie:1987:IDM

[Bai87] Richard T. Baillie. Inference in dynamic models containing ‘surprise’ variables. Journal of Econometrics, 35(1):101–117, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900832.

Baksalary:1983:IPF

[Bak83] Jerzy K. Baksalary. An invariance property of Farebrother’s procedure for estimation with aggregated data. Journal of Econometrics, 22(3):317–322, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901070.

Balestra:1980:NET

[Bal80] Pietro Balestra. A note on the exact transformation associated with the first-order moving average process. Journal of Econometrics, 14(3):381–394, December 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900342. See note [PH83].

Baltagi:1981:PES

[Bal81a] Badi H. Baltagi. Pooling: An experimental study of alternative testing and estimation procedures in a two-way error component model. Journal of Econometrics, 17(1):21–49, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900579.

Baltagi:1981:SEE

[Bal81b] Badi H. Baltagi. Simultaneous equations with error components. Journal of Econometrics, 17(2):189–200, November 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900269.

Balestra:1983:NAP

[Bal83] Pietro Balestra. A note on Amemiya’s partially generalized least squares. Journal of Econometrics, 23(2):285–290, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390083H.

Barnett:1980:EMAa

[Bar80a] William A. Barnett. Economic monetary aggregates an application of index number and aggregation theory. Journal of Econometrics, 14(1):11–48, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900706.

Barnett:1980:EMAb

[Bar80b] William A. Barnett. Economic monetary aggregates-reply. Journal of Econometrics, 14(1):57–59, September 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900731.

Barnett:1980:EIP

[Bar80c] William A. Barnett. Editor’s introduction to Part I. Journal of Econometrics, 14(1):3–8, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768090069X.

Bartholomew:1983:LVM

[Bar83] D. J. Bartholomew. Latent variable models for ordered categorical data. Journal of Econometrics, 22(1–2):229–243, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768390101X.

Barnett:1985:MLT

[Bar85] William A. Barnett. The minflex–Laurent translog flexible functional form. Journal of Econometrics, 30(1–2):33–44, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901307.

Basmann:1988:CTO

[Bas88] R. L. Basmann. Causality tests and observationally equivalent representations of econometric models. Journal of Econometrics, 39(1–2):69–104, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900413.

Batchelor:1981:AEU

[Bat81] R. A. Batchelor. Aggregate expectations under the stable laws. Journal of Econometrics, 16(2):199–210, June 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768190107X.

Berger:1984:BIS

[BB84] J. Berger and L. M. Berliner. Bayesian input in Stein estimation and a new minimax empirical Bayes estimator. Journal of Econometrics, 25(1–2):87–108, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900393.

Bernanke:1988:ANN

[BBR88] Ben Bernanke, Henning Bohn, and Peter C. Reiss. Alternative non-nested specification tests of time-series investment models. Journal of Econometrics, 37(3):293–326, March 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900085.

Battese:1988:PFL

[BC88] George E. Battese and Tim J. Coelli. Prediction of firm-level technical efficiencies with a generalized frontier production function and panel data. Journal of Econometrics, 38(3):387–399, July 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768890053X.

Bianchi:1981:EAS

[BCC81] Carlo Bianchi, Giorgio Calzolari, and Paolo Corsi. Estimating asymptotic standard errors and inconsistencies of impact multipliers in nonlinear econometric models. Journal of Econometrics, 16(3):277–294, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900312.

Beggs:1981:APD

[BCH81] S. Beggs, S. Cardell, and J. Hausman. Assessing the potential demand for electric cars. Journal of Econometrics, 17(1):1–19, September 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900567.

Basmann:1985:DBN

[BDFW85] R. L. Basmann, C. A. Diamond, J. C. Frentrup, and S. N. White. On deviations between neoclassical and GFT-based true costof-living indexes derived from the same demand function system. Journal of Econometrics, 30(1–2):45–66, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901319.

Bekker:1989:IRF

[Bek89] Paul A. Bekker. Identification in restricted factor models and the evaluation of rank conditions. Journal of Econometrics, 41 (1):5–16, May 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900407.

Belsley:1980:ECN

[Bel80] David A. Belsley. On the efficient computation of the nonlinear full-information maximum-likelihood estimator. Journal of Econometrics, 14(2):203–225, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900913.

Belsley:1981:AQR

[Bel81] David A. Belsley. Assessing the quality of regression estimates through a test for signal-to-noise and its application to detecting harmful collinearity. Journal of Econometrics, 16(1):167, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900968.

Belsley:1982:APH

[Bel82] David A. Belsley. Assessing the presence of harmful collinearity and other forms of weak data through a test for signal-to-noise. Journal of Econometrics, 20(2):211–253, November 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900203.

Belsley:1988:CML

[Bel88a] David A. Belsley. Conditioning in models with logs. Journal of Econometrics, 38(1–2):127–143, May/June 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900309.

Belsley:1988:EI

[Bel88b] David A. Belsley. Editor’s introduction. Journal of Econometrics, 38(1–2):3–5, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407688900243.

Bentler:1983:SES

[Ben83] P. M. Bentler. Simultaneous equation systems as moment structure models: With an introduction to latent variable models. Journal of Econometrics, 22(1–2):13–42, May/June 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900921.

Bera:1982:NTD

[Ber82] Anil K. Bera. A note on testing demand homogeneity. Journal of Econometrics, 18(2):291–294, February 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900446.

Berndt:1986:EI

[BF86a] Ernst R. Berndt and Melvyn A. Fuss. Editors’ introduction. Journal of Econometrics, 33(1–2):1–5, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900242.

Berndt:1986:PMA

[BF86b] Ernst R. Berndt and Melvyn A. Fuss. Productivity measurement with adjustments for variations in capacity utilization and other forms of temporary equilibrium. Journal of Econometrics, 33(1–2):7–29, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900254.

Barnett:1985:EI

[BG85] William A. Barnett and A. Ronald Gallant. Editor’s introduction. Journal of Econometrics, 30(1–2):1, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901289.

Bunzel:1982:SDR

[BH82] Henning Bunzel and Svend Hylleberg. Seasonality in dynamic regression models: A comparative study of finite sample properties of various regression estimators including band spectrum regression. Journal of Econometrics, 19(2–3):345–366, August 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900100.

Berg:1984:BLI

[BH84] Peter Ter Berg and Rins Harkema. Bayesian limited-information analysis of nonlinear simultaneous equations systems. Journal of Econometrics, 24(3):379–395, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900605.

Bierens:1988:NLR

[BH88] Herman J. Bierens and Joop Hartog. Non-linear regression with discrete explanatory variables, with an application to the earnings function. Journal of Econometrics, 38(3):269–299, July 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900474.

Boyes:1989:EAB

[BHL89] William J. Boyes, Dennis L. Hoffman, and Stuart A. Low. An econometric analysis of the bank credit scoring problem. Journal of Econometrics, 40(1):3–14, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900262.

Barnett:1986:RWB

[BHW86] William A. Barnett, Melvin J. Hinich, and Warren E. Weber. The regulatory wedge between the demand-side and supplyside aggregation-theoretic monetary aggregates. Journal of Econometrics, 33(1–2):165–185, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900321.

Bierens:1982:CMS

[Bie82] Herman J. Bierens. Consistent model specification tests. Journal of Econometrics, 20(1):105–134, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682901051.

Bierens:1984:MST

[Bie84] Herman J. Bierens. Model specification testing of time series regressions. Journal of Econometrics, 26(3):323–353, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900253.

Bierens:1987:AMS

[Bie87] Herman J. Bierens. Armax model specification testing, with an application to unemployment in the Netherlands. Journal of Econometrics, 35(1):161–190, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900868.

Biorn:1981:EER

[Biø81] Erik Biørn. Estimating economic relations from incomplete crosssection/time-series data. Journal of Econometrics, 16(2):221–236, June 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901093.

Bera:1982:MST

[BJ82] Anil K. Bera and Carlos M. Jarque. Model specification tests: A simultaneous approach. Journal of Econometrics, 20(1):59– 82, October 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682901038.

Bronars:1987:GDU

[BJ87] Stephen G. Bronars and Dennis W. Jansen. The geographic distribution of unemployment rates in the U.S.: A spatial-time series analysis. Journal of Econometrics, 36(3):251–279, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900029.

Bock:1984:SFI

[BJY84] M. E. Bock, G. G. Judge, and T. A. Yancey. A simple form for the inverse moments of non-central χ^{2} and F random variables and certain confluent hypergeometric functions. Journal of Econometrics, 25(1–2):217–234, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900484.

Bloom:1985:CTB

[BK85] David E. Bloom and Mark R. Killingsworth. Correcting for truncation bias caused by a latent truncation variable. Journal of Econometrics, 27(1):131–135, January 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768590048X.

Burdett:1985:LDD

[BKS85] Kenneth Burdett, Nicholas M. Kiefer, and Sunil Sharma. Layoffs and duration dependence in a model of turnover. Journal of Econometrics, 28(1):51–69, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900661.

Blackburn:1989:IMC

[Bla89] McKinley L. Blackburn. Interpreting the magnitude of changes in measures of income inequality. Journal of Econometrics, 42(1):21–25, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900729.

Bloch:1989:IED

[Blo89] Anthony M. Bloch. Identification and estimation of dynamic errors-in-variables models. Journal of Econometrics, 41(1):145– 158, May 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900468.

Blundell:1987:EI

[Blu87] Richard Blundell. Editor’s introduction. Journal of Econometrics, 34(1–2):1–4, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900649.

Barnett:1985:TDG

[BLW85] William A. Barnett, Yul W. Lee, and Michael D. Wolfe. The three-dimensional global properties of the minflex Laurent, generalized Leontief, and translog flexible functional forms. Journal of Econometrics, 30(1–2):3–31, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901290.

Barnett:1987:GPT

[BLW87] William A. Barnett, Yul W. Lee, and Michael Wolfe. The global properties of the two minflex Laurent flexible functional forms. Journal of Econometrics, 36(3):281–298, November 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900030.

Brook:1980:ELL

[BM80] Richard J. Brook and Terry Moore. On the expected length of the least squares coefficient vector. Journal of Econometrics, 12(2):245–246, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768090010X.

Blundell:1987:BAT

[BM87] Richard Blundell and Costas Meghir. Bivariate alternatives to the Tobit model. Journal of Econometrics, 34(1–2):179–200, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900728.

Butler:1989:UIM

[BM89] Richard J. Butler and James B. McDonald. Using incomplete moments to measure inequality. Journal of Econometrics, 42(1):109– 119, September 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900791.

Binkley:1984:IAD

[BN84] James K. Binkley and Glenn Nelson. Impacts of alternative degrees of freedom corrections in two and three stage least squares. Journal of Econometrics, 24(3):223–233, March 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900502.

Bollerslev:1986:GAC

[Bol86] Tim Bollerslev. Generalized autoregressive conditional heteroskedasticity. Journal of Econometrics, 31(3):307–327, April 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900631.

Bourguignon:1989:FSS

[Bou89] François Bourguignon. Family size and social utility: Income distribution dominance criteria. Journal of Econometrics, 42(1):67– 80, September 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900766.

Bekker:1986:ILS

[BP86] Paul A. Bekker and D. S. G. Pollock. Identification of linear stochastic models with covariance restrictions. Journal of Econometrics, 31(2):179–208, March 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900473.

Bauwens:1985:PRV

[BR85] Luc Bauwens and Jean-François Richard. A 1–1 poly-t random variable generator with application to Monte Carlo integration. Journal of Econometrics, 29(1–2):19–46, July/August 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900314.

Breusch:1980:UIR

[Bre80] Trevor S. Breusch. Useful invariance results for generalized regression models. Journal of Econometrics, 13(3):327–340, August 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900834.

Bresnahan:1981:DMC

[Bre81] Timothy F. Bresnahan. Departures from marginal-cost pricing in the American automobile industry: Estimates for 1977–1978. Journal of Econometrics, 17(2):201–227, November 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900270.

Breusch:1987:MLE

[Bre87] Trevor S. Breusch. Maximum likelihood estimation of random effects models. Journal of Econometrics, 36(3):383–389, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900108.

Broemeling:1982:I

[Bro82] Lyle Broemeling. Introduction. Journal of Econometrics, 19(1): 1–5, May 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900471.

Booth:1982:BAR

[BS82] N. B. Booth and A. F. M. Smith. A Bayesian approach to retrospective identification of change-points. Journal of Econometrics, 19(1):7–22, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900483.

Bronsard:1984:PEC

[BSB84] Camille Bronsard and Lise Salvas-Bronsard. On price exogeneity in complete demand systems. Journal of Econometrics, 24(3): 235–247, March 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900514.

Boggs:1988:CEO

[BSDS88] Paul T. Boggs, Clifford H. Spiegelman, Janet R. Donaldson, and Robert B. Schnabel. A computational examination of orthogonal distance regression. Journal of Econometrics, 38(1–2):169–201, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900322.

Bunch:1988:CAM

[Bun88] David S. Bunch. A comparison of algorithms for maximum likelihood estimation of choice models. Journal of Econometrics, 38(1–2):145–167, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900310.

Burmeister:1982:KFE

[BW82] Edwin Burmeister and Kent D. Wall. Kalman filtering estimation of unobserved rational expectations with an application to the German hyperinflation. Journal of Econometrics, 20(2):255– 284, November 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900215.

Behrman:1987:HDM

[BW87] Jere R. Behrman and Barbara L. Wolfe. How does mother’s schooling affect family health, nutrition, medical care usage, and household sanitation? Journal of Econometrics, 36(1–2):185–204, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900492.

Bohrer:1984:ANE

[BY84] R. Bohrer and T. A. Yancey. Algorithms for numerical evaluation of Stein-like and limited-translation estimators. Journal of Econometrics, 25(1–2):235–239, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900496.

Binswanger:1987:DCC

[BYBM87] Hans Binswanger, Maw-Cheng Yang, Alan Bowers, and Yair Mundlak. On the determinants of cross-country aggregate agricultural supply. Journal of Econometrics, 36(1–2):111–131, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900467.

Campos:1986:FSP

[Cam86] Julia Campos. Finite-sample properties of the instrumentalvariables estimator for dynamic simultaneous-equation subsystems with ARMA disturbances. Journal of Econometrics, 32(3):333–366, August 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407686900199.

Cameron:1987:IGC

[Cam87] Trudy Ann Cameron. The impact of grouping coarseness in alternative grouped-data regression models. Journal of Econometrics, 35(1):37–57, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900807.

Capalbo:1986:TEP

[Cap86] Susan M. Capalbo. Temporary equilibrium production models for a common-property renewable-resource sector. Journal of Econometrics, 33(1–2):263–284, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900369.

Carter:1981:ISR

[Car81] R. A. L. Carter. Improved Stein-rule estimator for regression problems. Journal of Econometrics, 17(1):113–123, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900634.

Caves:1980:EAR

[CC80] Douglas W. Caves and Laurits R. Christensen. Econometric analysis of residential time-of-use electricity pricing experiments. Journal of Econometrics, 14(3):287–306, December 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900299.

Charnes:1985:FDE

[CCG^{+}85] A. Charnes, W. W. Cooper, B. Golany, L. Seiford, and J. Stutz. Foundations of data envelopment analysis for Pareto– Koopmans efficient empirical production functions. Journal of Econometrics, 30(1–2):91–107, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901332.

Caves:1984:CRC

[CCH84a] Douglas W. Caves, Laurits R. Christensen, and Joseph A. Herriges. Consistency of residential customer response in time-of-use electricity pricing experiments. Journal of Econometrics, 26(1–2):179–203, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900174.

Caves:1984:MAR

[CCH84b] Douglas W. Caves, Laurits R. Christensen, and Joseph A. Herriges. Modelling alternative residential peak-load electricity rate structures. Journal of Econometrics, 24(3):249–268, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900526.

Caves:1984:CDM

[CCSH84] Douglas W. Caves, Laurits R. Christensen, Philip E. Schoech, and Wallace Hendricks. A comparison of different methodologies in a case study of residential time-of-use electricity pricing: Cost–Benefit Analysis. Journal of Econometrics, 26(1– 2):17–34, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407684900113.

Cleveland:1988:RLF

[CDG88] William S. Cleveland, Susan J. Devlin, and Eric Grosse. Regression by local fitting: Methods, properties, and computational algorithms. Journal of Econometrics, 37(1):87–114, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900772.

Carlin:1985:MMB

[CDJ85] J. B. Carlin, A. P. Dempster, and A. B. Jonas. On models and methods for Bayesian time series analysis. Journal of Econometrics, 30(1–2):67–90, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901320.

Cecchetti:1986:FPA

[Cec86] Stephen G. Cecchetti. The frequency of price adjustment: A study of the newsstand prices of magazines. Journal of Econometrics, 31(3):255–274, April 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900618.

Chalfant:1985:ESE

[CG85] James A. Chalfant and A. Ronald Gallant. Estimating substitution elasticities with the Fourier cost function: Some Monte Carlo results. Journal of Econometrics, 28(2):205–222, May 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901204.

Clarke:1987:EEV

[CGW87] Judith A. Clarke, David E. A. Giles, and T. Dudley Wallace. Estimating the error variance in regression after a preliminary test of restrictions on the coefficients. Journal of Econometrics, 34(3): 293–304, March 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900145.

Chetty:1986:DMA

[CH86] V. K. Chetty and J. J. Heckman. A dynamic model of aggregate output supply, factor demand and entry and exit for a competitive industry with heterogeneous plants. Journal of Econometrics, 33(1–2):237–262, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900357.

Chamberlain:1981:MDD

[Cha81] Gary Chamberlain. Models of duration dependence. Journal of Econometrics, 16(1):164, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900932.

Chamberlain:1982:MRM

[Cha82a] Gary Chamberlain. Multivariate regression models for panel data. Journal of Econometrics, 18(1):5–46, January 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768290094X.

Chavas:1982:RES

[Cha82b] Jean-Paul Chavas. Recursive estimation of simultaneous equation models. Journal of Econometrics, 18(2):207–217, February 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900367.

Chang:1983:EMS

[Cha83] Julius C. Chang. An econometric model of the short-run demand for workers and hours in the U.S. auto industry. Journal of Econometrics, 22(3):301–316, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901069.

Chang:1985:NCM

[Cha85] Y. C. Chang. A note on the covariance matrix of the maximum likelihood estimator in constrained multivariate linear regression. Journal of Econometrics, 28(2):247–252, May 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768590123X.

Chamberlain:1986:AES

[Cha86] Gary Chamberlain. Asymptotic efficiency in semi-parametric models with censoring. Journal of Econometrics, 32(2):189–218, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900382.

Chamberlain:1987:AEE

[Cha87] Gary Chamberlain. Asymptotic efficiency in estimation with conditional moment restrictions. Journal of Econometrics, 34(3): 305–334, March 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900157.

Chesher:1985:STZ

[Che85] Andrew Chesher. Score tests for zero covariances in recursive linear models for grouped or censored data. Journal of Econometrics, 28(3):291–305, June 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900028.

Chow:1981:CIP

[Cho81a] Gregory C. Chow. A comparison of the information and posterior probability criteria for model selection. Journal of Econometrics, 16(1):21–33, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900737.

Chow:1981:CSM

[Cho81b] Gregory C. Chow. On the control of structural models — comment. Journal of Econometrics, 15(1):25–28, January 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407681901263. See [Nor81a] and reply [Nor81b].

Cumby:1983:TST

[CHO83] Robert E. Cumby, John Huizinga, and Maurice Obstfeld. Twostep two-stage least squares estimation in models with rational expectations. Journal of Econometrics, 21(3):333–355, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900490.

Chesher:1987:RAG

[CI87] Andrew Chesher and Margaret Irish. Residual analysis in the grouped and censored normal linear model. Journal of Econometrics, 34(1–2):33–61, January/February 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900662.

Cinar:1987:SEL

[Çin87] E. Miné Çinar. The sensitivity of extended linear expenditure system household scales to income declaration errors. Journal of Econometrics, 34(3):361–372, March 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900194.

Colclough:1982:EEC

[CL82] William G. Colclough and Mark D. Lange. Empirical evidence of causality from consumer to wholesale prices. Journal of Econometrics, 19(2–3):379–384, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900124.

Cosslett:1985:SCL

[CL85] Stephen R. Cosslett and Lung-Fei Lee. Serial correlation in latent discrete variable models. Journal of Econometrics, 27(1):79– 97, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900454.

Clark:1989:TRR

[Cla89] Peter K. Clark. Trend reversion in real output and unemployment. Journal of Econometrics, 40(1):15–32, January 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900274.

Conerly:1989:ATC

[CM89] Michael D. Conerly and Edward R. Mansfield. An approximate test for comparing independent regression models with unequal error variances. Journal of Econometrics, 40(2):239–259, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900845.

Clements:1980:EMA

[CN80] Kenneth W. Clements and Phuong Nguyen. Economic monetary aggregates-comment. Journal of Econometrics, 14(1):49–53, September 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900718.

Conniffe:1985:ERE

[Con85] Denis Conniffe. Estimating regression equations with common explanatory variables but unequal numbers of observations. Journal of Econometrics, 27(2):179–196, February 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900867.

Cowell:1989:SVD

[Cow89] Frank A. Cowell. Sampling variance and decomposable inequality measures. Journal of Econometrics, 42(1):27–41, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900730.

Cooley:1987:LOA

[CP87] Thomas F. Cooley and William R. Parke. Likelihood and other approaches to prediction in dynamic models. Journal of Econometrics, 35(1):119–142, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900844.

Cooley:1989:PES

[CPC89] Thomas F. Cooley, William R. Parke, and Siddhartha Chib. Predictive efficiency for simple non-linear models. Journal of Econometrics, 40(1):33–44, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900286.

Calzolari:1987:CEF

[CPW87] Giorgio Calzolari, Lorenzo Panattoni, and Claus Weihs. Computational efficiency of FIML estimation. Journal of Econometrics, 36(3):299–310, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900042.

Catsiapis:1982:SSB

[CR82] George Catsiapis and Chris Robinson. Sample selection bias with multiple selection rules: An application to student aid grants. Journal of Econometrics, 18(3):351–368, April 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900884.

Cragg:1982:ETT

[Cra82] J. G. Cragg. Estimation and testing in time-series regression models with heteroscedastic disturbances. Journal of Econometrics, 20(1):135–157, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682901063.

Cramer:1987:MVR

[Cra87] J. S. Cramer. Mean and variance of R^{2} in small and moderate samples. Journal of Econometrics, 35(2–3):253–266, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900273.

Chavas:1987:SSE

[CS87] Jean-Paul Chavas and Kathleen Segerson. Stochastic specification and estimation of share equation systems. Journal of Econometrics, 35(2–3):337–358, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900327.

Chib:1988:BPR

[CTJ88] Siddhartha Chib, Ram C. Tiwari, and S. Rao Jammalamadaka. Bayes prediction in regressions with elliptical errors. Journal of Econometrics, 38(3):349–360, July 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900504.

Conrad:1987:EPT

[CU87] Klaus Conrad and Ralph Unger. Ex post tests for short-and long-run optimization. Journal of Econometrics, 36(3):339–358, November 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900066.

Deistler:1989:LDE

[DA89] M. Deistler and B. D. O. Anderson. Linear dynamic errorsin-variables models: Some structure theory. Journal of Econometrics, 41(1):39–63, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900420.

Dastoor:1983:SAT

[Das83] Naorayex K. Dastoor. Some aspects of testing non-nested hypotheses. Journal of Econometrics, 21(2):213–228, February 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900143.

Dastoor:1985:CAC

[Das85] Naorayex K. Dastoor. A classical approach to Cox’s test for non-nested hypotheses. Journal of Econometrics, 27(3):363– 370, March 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900119.

Davidson:1981:PEM

[Dav81] James E. H. Davidson. Problems with the estimation of moving average processes. Journal of Econometrics, 16(3):295–310, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900324.

Dufour:1985:DWT

[DD85] Jean-Marie Dufour and Marcel G. Dagenais. Durbin–Watson tests for serial correlation in regressions with missing observations. Journal of Econometrics, 27(3):371–381, March 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900120.

DeGooijer:1980:EMS

[De 80] Jan G. De Gooijer. Exact moments of the sample autocorrelations from series generated by general ARIMA processes of order (p, d, q), d = 0 or 1. Journal of Econometrics, 14(3):365– 379, December 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900330.

DeLeeuw:1983:MMA

[De 83] Jan De Leeuw. Models and methods for the analysis of correlation coefficients. Journal of Econometrics, 22(1–2):113–137, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900969.

Deaton:1985:PDT

[Dea85] Angus Deaton. Panel data from time series of cross-sections. Journal of Econometrics, 30(1–2):109–126, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901344.

Deaton:1987:EOC

[Dea87] Angus Deaton. Estimation of own- and cross-price elasticities from household survey data. Journal of Econometrics, 36(1–2):7–30, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900418.

Dent:1980:EI

[Den80a] Warren T. Dent. Editor’s introduction. Journal of Econometrics, 12(1):1–2, January 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900482.

Dent:1980:REL

[Den80b] Warren T. Dent. On restricted estimation in linear models. Journal of Econometrics, 12(1):49–58, January 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900524.

Denny:1983:GAI

[DF83] Michael Denny and Melvyn Fuss. A general approach to intertemporal and interspatial productivity comparisons. Journal of Econometrics, 23(3):315–330, December 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768390060X.

Doran:1983:REE

[DG83] H. E. Doran and W. E. Griffiths. On the relative efficiency of estimators which include the initial observations in the estimation of seemingly unrelated regressions with first-order autoregressive disturbances. Journal of Econometrics, 23(2):165–191, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390075G.

Diaz:1982:BDC

[Dia82] Joaquin Diaz. Bayesian detection of a change of scale parameter in sequences of independent gamma random variables. Journal of Econometrics, 19(1):23–29, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900495.

Dijkstra:1983:SCM

[Dij83] Theo Dijkstra. Some comments on maximum likelihood and partial least squares methods. Journal of Econometrics, 22(1–2):67– 90, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900945.

Duan:1987:DFL

[DL87] Naihua Duan and Ker-Chau Li. Distribution-free and link-free estimation for the sample selection model. Journal of Econometrics, 35(1):25–35, May 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900790.

Davidson:1983:TSM

[DM83] Russell Davidson and James G. MacKinnon. Testing the specification of multivariate models in the presence of alternative hypotheses. Journal of Econometrics, 23(3):301–313, December 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900593.

Davidson:1984:CST

[DM84] Russell Davidson and James G. MacKinnon. Convenient specification tests for logit and probit models. Journal of Econometrics, 25(3):241–262, July 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407684900010.

Don:1982:RV

[Don82] F. J. Henk Don. Restrictions on variables. Journal of Econometrics, 18(3):369–393, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900896.

Don:1983:RV

[Don83] F. J. H. Don. Restrictions on variables. Journal of Econometrics, 23(2):291–292, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390084I.

Don:1986:SLI

[Don86] F. J. Henk Don. The specification of least informative error distributions. Journal of Econometrics, 31(1):81–91, February 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900576.

Doran:1981:OOR

[Dor81] Howard E. Doran. Omission of an observation from a regression analysis: A discussion on efficiency loss, with applications. Journal of Econometrics, 16(3):367–374, August 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768190035X.

Diewert:1985:TCC

[DP85] W. E. Diewert and C. Parkan. Tests for the consistency of consumer data. Journal of Econometrics, 30(1–2):127–147, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901356.

Dufour:1985:SRE

[DR85] Jean-Marie Dufour and Roch Roy. Some robust exact results on sample autocorrelations and tests of randomness. Journal of Econometrics, 29(3):257–273, September 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901551.

Dubin:1989:RRS

[DR89] Jeffrey A. Dubin and Geoffrey S. Rothwell. Risk and reactor safety systems adoption. Journal of Econometrics, 42(2):201– 218, October 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900031.

Deville:1983:CAE

[DS83] J.-C. Deville and G. Saporta. Correspondence analysis, with an extension towards nominal time series. Journal of Econometrics, 22(1–2):169–189, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900982.

Dwivedi:1984:EFS

[DS84] T. D. Dwivedi and V. K. Srivastava. Exact finite sample properties of double k-class estimators in simultaneous equations. Journal of Econometrics, 25(3):263–283, July 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900022.

Deprins:1985:NAR

[DS85] D. Deprins and L. Simar. A note on the asymptotic relative efficiency of M.L.E. in a linear model with Gamma disturbances. Journal of Econometrics, 27(3):383–386, March 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900132.

Davies:1989:OGI

[DS89] J. B. Davies and A. F. Shorrocks. Optimal grouping of income and wealth data. Journal of Econometrics, 42(1):97–108, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768990078X.

Dufour:1982:RSA

[Duf82] Jean-Marie Dufour. Recursive stability analysis of linear regression relationships: An exploratory methodology. Journal of Econometrics, 19(1):31–76, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900501.

Dufour:1988:EDV

[Duf88] Jean-Marie Dufour. Estimators of the disturbance variance in econometric models: Small-sample bias and the existence of moments. Journal of Econometrics, 37(2):277–292, February 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900073.

Duncan:1980:AML

[Dun80] Gregory M. Duncan. Approximate maximum likelihood estimation with data sets that exceed computer limits. Journal of Econometrics, 14(2):257–264, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900950.

Duncan:1986:EI

[Dun86a] Gregory M. Duncan. Editor’s introduction. Journal of Econometrics, 32(1):1–3, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407686900096.

Duncan:1986:SPC

[Dun86b] Gregory M. Duncan. A semi-parametric censored regression estimator. Journal of Econometrics, 32(1):5–34, June 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900102.

Duncan:1987:SAE

[Dun87] Gregory M. Duncan. A simplified approach to M-estimation with application to two-stage estimators. Journal of Econometrics, 34(3):373–389, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900200.

Durbin:1981:ADS

[Dur81] J. Durbin. Approximations for densities of sufficient estimators. Journal of Econometrics, 16(1):165, May 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900944.

Durbin:1988:PCS

[Dur88a] J. Durbin. Is a philosophical consensus for statistics attainable? Journal of Econometrics, 37(1):51–61, January 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407688900735. See [Fie88] and reply [Dur88b].

Durbin:1988:RSF

[Dur88b]  J. Durbin.  Reply to Stephen E. Fienberg’s discussion.

Journal of Econometrics, 37(1):65, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407688900759. See [Dur88a, Fie88].

Domowitz:1982:MMD

[DW82] Ian Domowitz and Halbert White. Misspecified models with dependent observations. Journal of Econometrics, 20(1):35–58, October 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682901026.

Diewert:1988:NQS

[DW88] W. E. Diewert and T. J. Wales. A normalized quadratic semiflexible functional form. Journal of Econometrics, 37(3):327– 342, March 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900097.

Epstein:1980:ECU

[ED80] L. Epstein and M. Denny. Endogenous capital utilization in a short-run production model: Theory and an empirical application. Journal of Econometrics, 12(2):189–207, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900068.

Engle:1989:MSL

[EGH89] R. F. Engle, C. W. J. Granger, and J. J. Hallman. Merging short-and long-run forecasts: An application of seasonal cointegration to monthly electricity sales forecasting. Journal of Econometrics, 40(1):45–62, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900298.

Evans:1985:POT

[EK85] Merran A. Evans and Maxwell L. King. A point optimal test for heteroscedastic disturbances. Journal of Econometrics, 27(2):163–178, February 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900855.

Evans:1988:FCT

[EK88] Merran A. Evans and Maxwell L. King. A further class of tests for heteroscedasticity. Journal of Econometrics, 37(2):265–276, February 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900061.

Engle:1985:DMH

[ELW85] Robert F. Engle, David M. Lilien, and Mark Watson. A dynamic model of housing price determination. Journal of Econometrics, 28(3):307–326, June 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768590003X.

Engle:1982:GAL

[Eng82] Robert F. Engle. A general approach to Lagrange multiplier model diagnostics. Journal of Econometrics, 20(1):83–104, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768290104X.

Evans:1983:PHC

[EW83] Lewis Evans and Graeme Wells. Pierce and haugh on characterizations of causality: A re-examination. Journal of Econometrics, 23(3):331–335, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900611.

Epstein:1985:EDT

[EY85a] Larry G. Epstein and Adonis J. Yatchew. The empirical determination of technology and expectations: A simplified procedure. Journal of Econometrics, 27(2):235–258, February 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900909.

Epstein:1985:NPH

[EY85b] Larry G. Epstein and Adonis J. Yatchew. Non-parametric hypothesis testing procedures and applications to demand analysis. Journal of Econometrics, 30(1–2):149–169, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901368.

Engle:1987:FTC

[EY87] Robert F. Engle and Byung Sam Yoo. Forecasting and testing in co-integrated systems. Journal of Econometrics, 35(1): 143–159, May 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900856.

Farebrother:1987:SFC

[Far87] R. W. Farebrother. The statistical foundations of a class of parametric tests for heteroscedasticity. Journal of Econometrics, 36(3):359–368, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900078.

Fernandez:1986:NPM

[Fer86] Luis Fernandez. Non-parametric maximum likelihood estimation of censored regression models. Journal of Econometrics, 32(1): 35–57, June 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900114.

Fomby:1983:ETS

[FG83] Thomas B. Fomby and David K. Guilkey. An examination of two-step estimators for models with lagged dependent variables and autocorrelated errors. Journal of Econometrics, 22(3):291– 300, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901057.

Fourgeaud:1984:STR

[FGP84] C. Fourgeaud, C. Gourieroux, and J. Pradel. Some theoretical results for generalized ridge regression estimators. Journal of Econometrics, 25(1–2):191–203, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900460.

Fuller:1980:PFO

[FH80] Wayne A. Fuller and David P. Hasza. Predictors for the firstorder autoregressive process. Journal of Econometrics, 13(2): 139–157, June 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900123.

Flinn:1982:NMA

[FH82] C. Flinn and J. Heckman. New methods for analyzing structural models of labor force dynamics. Journal of Econometrics, 18(1):115–168, January 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900975.

Fienberg:1988:DPC

[Fie88] Stephen E. Fienberg. Discussion of: Is a philosophical consensus for statistics attainable? by J. Durbin. Journal of Econometrics, 37(1):63–64, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407688900747. See [Dur88a, Dur88b].

Fisher:1983:TTS

[Fis83] Gordon R. Fisher. Tests for two separate regressions. Journal of Econometrics, 21(1):117–132, January 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901227.

Flannery:1981:IUE

[FJ81] Mark J. Flannery and Lewis Johnson. Indexing the U.S. economy: Simulation results with the MPS model. Journal of Econometrics, 15(1):93–114, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901317.

Fishe:1981:EIE

[FL81] Raymond P. H. Fishe and Kajal Lahiri. On the estimation of inflationary expectations from qualitative responses. Journal of Econometrics, 16(1):89–102, May 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900774.

Forsund:1980:SFP

[FLS80] Finn R. Førsund, C. A. Knox Lovell, and Peter Schmidt. A survey of frontier production functions and of their relationship to efficiency measurement. Journal of Econometrics, 13(1):5–25, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900408.

Fisher:1981:APA

[FM81] Gordon R. Fisher and Michael McAleer. Alternative procedures and associated tests of significance for non-nested hypotheses. Journal of Econometrics, 16(1):103–119, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900786.

Fienberg:1983:LMC

[FM83] Stephen E. Fienberg and Michael M. Meyer. Loglinear models and categorical data analysis with psychometric and econometric applications. Journal of Econometrics, 22(1–2):191–214, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900994.

Fujikoshi:1982:AED

[FMKT82] Yasunori Fujikoshi, Kimio Morimune, Naoto Kunitomo, and Masanobu Taniguchi. Asymptotic expansions of the distributions of the estimates of coefficients in a simultaneous equation system. Journal of Econometrics, 18(2):191–205, February 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900355.

Florens:1981:SIL

[FMR81] J.-P. Florens, M. Mouchart, and J.-F. Richard. Specification and inference in linear models. Journal of Econometrics, 16(1):153, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/030440768190083X.

Fair:1980:FIE

[FP80] Ray C. Fair and William R. Parke. Full-information estimates of a nonlinear macroeconometric model. Journal of Econometrics, 13(3):269–291, August 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900809.

Fuller:1980:UIV

[Ful80] Wayne A. Fuller. The use of indicator variables in computing predictions. Journal of Econometrics, 12(2):231–243, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900093.

Gallant:1981:BFF

[Gal81] A. Ronald Gallant. On the bias in flexible functional forms and an essentially unbiased form: The Fourier flexible form. Journal of Econometrics, 15(2):211–245, February 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901159.

Gallant:1982:UDP

[Gal82] A. Ronald Gallant. Unbiased determination of production technologies. Journal of Econometrics, 20(2):285–323, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900227.

Garber:1989:RLH

[Gar89] Steven Garber. The reserve-labor hypothesis, short-run pricing theories, and the employment-output relationship. Journal of Econometrics, 42(2):219–245, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900043.

Griffiths:1984:SSP

[GB84] W. E. Griffiths and P. A. A. Beesley. The small-sample properties of some preliminary test estimators in a linear model with autocorrelated errors. Journal of Econometrics, 25(1–2):49–61, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900368.

Griffiths:1980:NBE

[GD80] William Griffiths and Dan Dao. A note on a Bayesian estimator in an autocorrelated error model. Journal of Econometrics, 12(3): 389–392, April 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900640.

Garthwaite:1985:DSB

[GD85] Paul H. Garthwaite and James M. Dickey. Double- and singlebisection methods for subjective probability assessment in a location-scale family. Journal of Econometrics, 29(1–2):149–163, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900363.

Gelfand:1988:IED

[GD88] Alan E. Gelfand and Dipak K. Dey. Improved estimation of the disturbance variance in a linear regression model. Journal of Econometrics, 39(3):387–395, November 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900656.

Geisser:1985:IPP

[Gei85] Seymour Geisser. Interval prediction for Pareto and exponential observables. Journal of Econometrics, 29(1–2):173–185, July/ August 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900387.

Gersovitz:1980:CPD

[Ger80] Mark Gersovitz. Classification probabilities for the disequilibrium model. Journal of Econometrics, 14(2):239–246, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900937.

Geweke:1988:AAM

[Gew88] John Geweke. Antithetic acceleration of Monte Carlo integration in Bayesian inference. Journal of Econometrics, 38(1–2):73–89, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900279.

Geweke:1989:EPD

[Gew89] John Geweke. Exact predictive densities for linear models with arch disturbances. Journal of Econometrics, 40(1):63–86, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900304.

Gallant:1980:CCL

[GG80] A. Ronald Gallant and Thomas M. Gerig. Computations for constrained linear models. Journal of Econometrics, 12(1):59– 84, January 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900536.

Gallant:1984:ICR

[GG84] A. Ronald Gallant and Gene H. Golub. Imposing curvature restrictions on flexible functional forms. Journal of Econometrics, 26(3):295–321, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900241.

Griliches:1986:EVP

[GH86] Zvi Griliches and Jerry A. Hausman. Errors in variables in panel data. Journal of Econometrics, 31(1):93–118, February 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900588.

Gourieroux:1981:KTL

[GHM81] Christian Gouriéroux, Alberto Holly, and Alain Monfort. Kuhn– Tucker, likelihood ratio and Wald tests for nonlinear models with inequality constraints on the parameters. Journal of Econometrics, 16(1):166, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900956.

Ghosh:1989:MLE

[Gho89] Damayanti Ghosh. Maximum likelihood estimation of the dynamic shock-error model. Journal of Econometrics, 41(1):121– 143, May 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900456.

Gallant:1984:CBP

[GK84] A. Ronald Gallant and Roger W. Koenker. Costs and benefits of peak-load pricing of electricity: A continuous-time econometric approach. Journal of Econometrics, 26(1–2):83–113, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900149.

Golub:1980:RSD

[GKP80] Gene Golub, Virginia Klema, and Stephen C. Peters. Rules and software for detecting rank degeneracy. Journal of Econometrics, 12(1):41–48, January 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900512.

Gertler:1987:UFR

[GLS87] Paul Gertler, Luis Locay, and Warren Sanderson. Are user fees regressive?: The welfare implications of health care financing proposals in Peru. Journal of Econometrics, 36(1–2):67–88, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900443.

Geary:1980:IST

[GM80] Patrick T. Geary and Edward J. McDonnell. Implications of the specification of technologies: Further evidence. Journal of Econometrics, 14(2):247–255, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900949.

Geweke:1981:ERM

[GM81a] John Geweke and Richard Meese. Estimating regression models of finite but unknown order. Journal of Econometrics, 16(1):162, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900919.

Gourieroux:1981:APM

[GM81b] Christian Gourieroux and Alain Monfort. Asymptotic properties of the maximum likelihood estimator in dichotomous logit models. Journal of Econometrics, 17(1):83–97, September 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900609.

Gallant:1985:EID

[GM85] A. Ronald Gallant and John F. Monahan. Explicitly infinitedimensional Bayesian analysis of production technologies. Journal of Econometrics, 30(1–2):171–201, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768590137X.

Geweke:1983:CAT

[GMD83] John Geweke, Richard Meese, and Warren Dent. Comparing alternative tests of causality in temporal systems: Analytic results and experimental evidence. Journal of Econometrics, 21(2):161– 194, February 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768390012X.

Gourieroux:1987:GR

[GMRT87a] Christian Gourieroux, Alain Monfort, Eric Renault, and Alain Trognon. Generalised residuals. Journal of Econometrics, 34(1–2):5–32, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900650.

Gourieroux:1987:SR

[GMRT87b] Christian Gourieroux, Alain Monfort, Eric Renault, and Alain Trognon. Simulated residuals. Journal of Econometrics, 34(1–2):201–252, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768790073X.

Gourieroux:1983:TNN

[GMT83] Christian Gourieroux, Alain Monfort, and Alain Trognon. Testing nested or non-nested hypotheses. Journal of Econometrics, 21(1):83–115, January 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683901215.

Gastwirth:1989:SPM

[GNW89] Joseph L. Gastwirth, Tapan K. Nayak, and Jane-Ling Wang. Statistical properties of measures of between-group income differentials. Journal of Econometrics, 42(1):5–19, September 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900717.

Goldberger:1981:LRA

[Gol81] Arthur S. Goldberger. Linear regression after selection. Journal of Econometrics, 15(3):357–366, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901007.

Guilkey:1981:CRL

[GP81] David K. Guilkey and J. Michael Price. On comparing restricted least squares estimators. Journal of Econometrics, 15(3):397– 404, April 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901020.

Godfrey:1983:TNN

[GP83] L. G. Godfrey and M. H. Pesaran. Tests of non-nested regression models: Small sample adjustments and Monte Carlo evidence. Journal of Econometrics, 21(1):133–154, January 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901239.

Goldfeld:1981:EMN

[GQ81a]  Stephen M. Goldfeld and Richard E. Quandt.  Econo-

metric modelling with non-normal disturbances. Journal of Econometrics, 17(2):141–155, November 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900233.

Goldfeld:1981:SMD

[GQ81b] Stephen M. Goldfeld and Richard E. Quandt. Single-market disequilibrium models: Estimating and testing. Journal of Econometrics, 16(1):157, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900877.

Granger:1980:LMR

[Gra80] C. W. J. Granger. Long memory relationships and the aggregation of dynamic models. Journal of Econometrics, 14(2):227– 238, October 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900925.

Granger:1981:SPT

[Gra81] C. W. J. Granger. Some properties of time series data and their use in econometric model specification. Journal of Econometrics, 16(1):121–130, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900798.

Granger:1988:SRD

[Gra88] C. W. J. Granger. Some recent development in a concept of causality. Journal of Econometrics, 39(1–2):199–211, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900450.

Greene:1980:MLE

[Gre80a] William H. Greene. Maximum likelihood estimation of econometric frontier functions. Journal of Econometrics, 13(1):27–56, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768090041X.

Greene:1980:EFF

[Gre80b] William H. Greene. On the estimation of a flexible frontier production model. Journal of Econometrics, 13(1):101–115, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900457.

Greene:1982:MLE

[Gre82] William H. Greene. Maximum likelihood estimation of stochastic frontier production models. Journal of Econometrics, 18(2):285– 289, February 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900434.

Greene:1983:ELD

[Gre83a] William H. Greene. Estimation of limited dependent variable models by ordinary least squares and the method of moments. Journal of Econometrics, 21(2):195–212, February 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900131.

Greenwald:1983:GAB

[Gre83b] Bruce C. Greenwald. A general analysis of bias in the estimated standard errors of least squares coefficients. Journal of Econometrics, 22(3):323–338, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901082.

Geweke:1981:LVM

[GS81] John F. Geweke and Kenneth J. Singleton. Latent variable models for time series: A frequency domain approach with an application to the permanent income hypothesis. Journal of Econometrics, 17(3):287–304, December 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900038.

Griffiths:1986:MCE

[GS86] W. E. Griffiths and K. Surekha. A Monte Carlo evaluation of the power of some tests for heteroscedasticity. Journal of Econometrics, 31(2):219–231, March 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900497.

Glymour:1988:LVC

[GS88] Clark Glymour and Peter Spirtes. Latent variables, causal models and overidentifying constraints. Journal of Econometrics, 39(1–2):175–198, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900449.

Gourieroux:1984:SPT

[GT84] C. Gourieroux and A. Trognon. Specification pre-test estimator. Journal of Econometrics, 25(1–2):15–27, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900332.

Granger:1989:IFA

[GWK89] C. W. J. Granger, Halbert White, and Mark Kamstra. Interval forecasting: An analysis based upon ARCH-quantile estimators. Journal of Econometrics, 40(1):87–96, January 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900316.

Hajivassiliou:1987:EDR

[Haj87] Vassilis A. Hajivassiliou. The external debt repayments problems of LDC’s: An econometric model based on panel data. Journal of Econometrics, 36(1–2):205–230, September/October 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900509.

Hall:1983:CCT

[Hal83] A. D. Hall. Confidence contours for two test statistics for nonnested regression models. Journal of Econometrics, 21(1):155– 160, January 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901240.

Hall:1984:SCT

[Hal84] Bronwyn H. Hall. Software for the computation of Tobit model estimates. Journal of Econometrics, 24(1–2):215–222, January/ February 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900800.

Hamilton:1986:SEE

[Ham86] James D. Hamilton. A standard error for the estimated state vector of a state-space model. Journal of Econometrics, 33(3):387– 397, December 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900047.

Hansen:1985:MCB

[Han85] Lars Peter Hansen. A method for calculating bounds on the asymptotic covariance matrices of generalized method of moments estimators. Journal of Econometrics, 30(1–2):203–238, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901381.

Han:1987:NPAb

[Han87a] Aaron K. Han. Non-parametric analysis of a generalized regression model: The maximum rank correlation estimator. Journal of Econometrics, 35(2–3):303–316, July 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900303.

Han:1987:NPAa

[Han87b] Aaron K. Han. A non-parametric analysis of transformations. Journal of Econometrics, 35(2–3):191–209, July 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900236.

Han:1989:AEC

[Han89] Aaron K. Han. Asymptotic efficiency calculations of the partial likelihood estimator. Journal of Econometrics, 41(2):237– 250, June 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768990095X.

Hartman:1982:NUA

[Har82] Raymond S. Hartman. A note on the use of aggregate data in individual choice models: Discrete consumer choice among alternative fuels for residential appliances. Journal of Econometrics, 18(3): 313–335, April 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900860.

Helliwell:1986:AOV

[HC86] John F. Helliwell and Alan Chung. Aggregate output with variable rates of utilization of employed factors. Journal of Econometrics, 33(1–2):285–310, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900370.

Holtz-Eakin:1988:TIE

[HE88] Douglas Holtz-Eakin. Testing for individual effects in autoregressive models. Journal of Econometrics, 39(3):297–307, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900607.

Hendry:1982:RPM

[Hen82] David F. Hendry. A reply to professors maasoumi and Phillips. Journal of Econometrics, 19(2–3):203–213, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900021.

Ham:1984:TSE

[HH84] John Ham and Cheng Hsiao. Two-stage estimation of structural labor supply parameters using interval data from the 1971 Canadian census. Journal of Econometrics, 24(1–2):133–158, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900769.

Howe:1981:AII

[HHCS^{+}81] Howard Howe, Ernesto Hernandez-Cata, Guy Stevens, Richard Berner, Peter Clark, and Sung Y. Kwack. Assessing international interdependence with a multi-country model. Journal of Econometrics, 15(1):65–92, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901305.

Hernandez-Iglesias:1981:CIP

[HIHI81] C. Hernández-Iglesias and F. Hernández-Iglesias. Causality and the independence phenomenon: The case of the demand for money. Journal of Econometrics, 15(2):247–263, February 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901160.

Hinkley:1985:DRP

[Hin85] David Hinkley. Discussion of Rossi’s paper. Journal of Econometrics, 30(1–2):363–364, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901460.

Hill:1987:IPP

[HJ87] R. Carter Hill and George Judge. Improved prediction in the presence of multicollinearity. Journal of Econometrics, 35(1): 83–100, May 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900820.

Harrison:1985:LMI

[HK85] M. J. Harrison and Gary Keogh. A Lagrange multiplier interpretation of disturbance estimators with an application to testing for nonlinearity. Journal of Econometrics, 27(2):259–269, February 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900910.

Hazilla:1986:TSF

[HK86] Michael Hazilla and Raymond J. Kopp. Testing for separable functional structure using temporary equilibrium models. Journal of Econometrics, 33(1–2):119–141, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900308.

Heiser:1983:ART

[HM83] Willem J. Heiser and Jacqueline Meulman. Analyzing rectangular tables by joint and constrained multidimensional scaling. Journal of Econometrics, 22(1–2):139–167, May/June 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900970.

Heijmans:1986:CML

[HM86] Risto D. H. Heijmans and Jan R. Magnus. Consistent maximum-likelihood estimation with dependent observations: The general (non-normal) case and the normal case. Journal of Econometrics, 32(2):253–285, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900400.

Havenner:1988:FER

[HM88] Arthur Havenner and Bagher Modjtahedi. Foreign exchange rates: A multiple currency and maturity analysis. Journal of Econometrics, 37(2):251–264, February 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768890005X.

Hoque:1988:EMP

[HMP88] Asraul Hoque, Jan R. Magnus, and Bahram Pesaran. The exact multi-period mean-square forecast error for the first-order autoregressive model. Journal of Econometrics, 39(3):327–346, November 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900620.

Hendry:1988:EAS

[HNS88] David F. Hendry, Adrian J. Neale, and Frank Srba. Econometric analysis of small linear systems using PC-FIML. Journal of Econometrics, 38(1–2):203–226, May/June 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900334.

Holbert:1982:BAS

[Hol82] Donald Holbert. A Bayesian analysis of a switching linear model. Journal of Econometrics, 19(1):77–87, May 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900513.

Honda:1988:SCL

[Hon88] Yuzo Honda. A size correction to the Lagrange multiplier test for heteroskedasticity. Journal of Econometrics, 38(3):375–386, July 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900528.

Horowitz:1986:DFL

[Hor86] Joel L. Horowitz. A distribution-free least squares estimator for censored linear regression models. Journal of Econometrics, 32(1): 59–84, June 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900126.

Harvey:1981:THS

[HP81a] A. C. Harvey and G. D. A. Phillips. Testing for heteroscedasticity in simultaneous equation models. Journal of Econometrics, 15(3): 311–340, April 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900981.

Harvey:1981:TSC

[HP81b] A. C. Harvey and G. D. A. Phillips. Testing for serial correlation in simultaneous equation models: Some further results. Journal of Econometrics, 17(1):99–105, September 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900610.

Hinich:1985:ICN

[HP85a] Melvin J. Hinich and Douglas M. Patterson. Identification of the coefficients in a non-linear: time series of the quadratic type. Journal of Econometrics, 30(1–2):269–288, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768590140X.

Hinich:1985:RMN

[HP85b] Melvin J. Hinich and Douglas M. Patterson. Reply to Marsh’s note. Journal of Econometrics, 30(1–2):297–299, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901423.

Hendry:1981:MFS

[HR81] David F. Hendry and Jean-François Richard. Model formulation to simplify selection when specification is uncertain. Journal of Econometrics, 16(1):159, May 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900890.

Hendry:1982:FEM

[HR82] David F. Hendry and Jean-François Richard. On the formulation of empirical models in dynamic econometrics. Journal of Econometrics, 20(1):3–33, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682901014.

Heckman:1985:AME

[HR85] James J. Heckman and Richard Robb. Alternative methods for evaluating the impact of interventions: An overview. Journal of Econometrics, 30(1–2):239–267, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901393.

Hausman:1987:STE

[HR87] Jerry A. Hausman and Paul A. Ruud. Specifying and testing econometric models for rank-ordered data. Journal of Econometrics, 34(1–2):83–104, January/February 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900686.

Hendry:1980:ACP

[HS80] David F. Hendry and Frank Srba. Autoreg: a computer program library for dynamic econometric models with autoregressive errors. Journal of Econometrics, 12(1):85–102, January 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900548.

Havenner:1981:RCA

[HS81a] A. Havenner and P. A. V. B. Swamy. A random coefficient approach to seasonal adjustment of economic time series. Journal of Econometrics, 15(2):177–209, February 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901147.

Hoffman:1981:TRI

[HS81b] Dennis L. Hoffman and Peter Schmidt. Testing the restrictions implied by the rational expectations hypothesis. Journal of Econometrics, 15(2):265–287, February 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901172.

Heckman:1982:EI

[HS82] J. Heckman and B. Singer. Editors’ introduction. Journal of Econometrics, 18(1):1–3, January 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900938.

Hoffman:1983:RST

[HS83] Dennis L. Hoffman and Don E. Schlagenhauf. Rationality, specification tests, and macroeconomic models. Journal of Econometrics, 21(3):367–386, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900519.

Heckman:1984:EDA

[HS84] James J. Heckman and Burton Singer. Econometric duration analysis. Journal of Econometrics, 24(1–2):63–132, January/ February 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900757.

Heckman:1985:E

[HS85] James J. Heckman and Burton Singer. Erratum. Journal of Econometrics, 27(1):137–138, January 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900491.

Harvey:1989:EIH

[HS89a] A. C. Harvey and James H. Stock. Estimating integrated higherorder continuous time autoregressions with an application to money-income causality. Journal of Econometrics, 42(3):319– 336, November 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0304407689900560.

Hirschberg:1989:RTP

[HS89b] Joseph G. Hirschberg and D. J. Slottje. Remembrance of things past the distribution of earnings across occupations and the kappa criterion. Journal of Econometrics, 42(1):121–130, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900808.

Hoffman:1989:EI

[HS89c] Dennis Hoffman and Peter Schmidt. Editors’ introduction. Journal of Econometrics, 40(1):1, January 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900250.

Hsieh:1983:HCC

[Hsi83] David A. Hsieh. A heteroscedasticity-consistent covariance matrix estimator for time series regressions. Journal of Econometrics, 22(3):281–290, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683901045.

Hsiao:1989:CES

[Hsi89] Cheng Hsiao. Consistent estimation for some nonlinear errors-in-variables models. Journal of Econometrics, 41(1):159–185, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768990047X.

Hsu:1982:RIS

[Hsu82] D. A. Hsu. Robust inferences for structural shift in regression models. Journal of Econometrics, 19(1):89–107, May 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900525.

Hausman:1981:PDU

[HT81] Jerry A. Hausman and William E. Taylor. Panel data and unobservable individual effects. Journal of Econometrics, 16(1):155, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900853.

Hausman:1984:APU

[HT84] Jerry A. Hausman and John Trimble. Appliance purchase and usage adaptation to a permanent time-of-day electricity rate schedule. Journal of Econometrics, 26(1–2):115–139, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900150.

Hulten:1986:PCC

[Hul86] Charles R. Hulten. Productivity change, capacity utilization, and the sources of efficiency growth. Journal of Econometrics, 33(1–2):31–50, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900266.

Humphrey:1981:ESF

[Hum81] David Burras Humphrey. Economies to scale in Federal Reserve check processing operations. Journal of Econometrics, 15(1):155– 173, January 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901342.

Howrey:1984:EDI

[HV84] E. Philip Howrey and Hal R. Varian. Estimating the distributional impact of time-of-day pricing of electricity. Journal of Econometrics, 26(1–2):65–82, September/October 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900137.

Hulten:1981:EED

[HW81] Charles R. Hulten and Frank C. Wykoff. The estimation of economic depreciation using vintage asset prices: An application of the Box–Cox power transformation. Journal of Econometrics, 15(3):367–396, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901019.

Hill:1984:RGS

[HZ84] R. C. Hill and R. F. Ziemer. The risk of general Steinlike estimators in the presence of multicollinearity. Journal of Econometrics, 25(1–2):205–216, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900472.

Izan:1980:PPR

[Iza80] Haji Y. Izan. To pool or not to pool?: A reexamination of Tobin’s food demand problem. Journal of Econometrics, 13(3):391– 402, August 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768090086X.

Jarque:1987:ALM

[Jar87] Carlos M. Jarque. An application of LDV models to household expenditure analysis in Mexico. Journal of Econometrics, 36(1–2):31–53, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768790042X.

Jennings:1980:SEE

[Jen80] L. S. Jennings. Simultaneous equations estimation: Computational aspects. Journal of Econometrics, 12(1):23–39, January 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900500.

Lee:1989:MR

[jL89] Myoung jae Lee. Mode regression. Journal of Econometrics, 42(3):337–349, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407689900572.

Jondrow:1982:ETI

[JLMS82] James Jondrow, C. A. Knox Lovell, Ivan S. Materov, and Peter Schmidt. On the estimation of technical inefficiency in the stochastic frontier production function model. Journal of Econometrics, 19(2–3):233–238, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900045.

Jorgenson:1985:EVE

[JS85] Dale W. Jorgenson and Daniel T. Slesnick. Efficiency versus equity in natural gas price regulation. Journal of Econometrics, 30(1–2):301–316, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901435.

Judge:1984:EI

[Jud84] George Judge. Editor’s introduction. Journal of Econometrics, 25(1–2):1–2, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407684900319.

Judge:1984:NOI

[JYBB84] G. G. Judge, T. A. Yancey, M. E. Bock, and R. Bohrer. The nonoptimality of the inequality restricted estimator under squared error loss. Journal of Econometrics, 25(1–2):165–177, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900447.

Judge:1987:ESP

[JYYT87] George Judge, Gang Yi, Thomas Yancey, and Timo Teräsvirta. The extended Stein procedure for simultaneous model selection and parameter estimation. Journal of Econometrics, 35(2–3): 375–391, July 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900340.

Kadane:1985:VCB

[Kad85] Joseph B. Kadane. Is victimization chronic? a Bayesian analysis of multinomial missing data. Journal of Econometrics, 29(1–2):47–67, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900326.

Kang:1985:NES

[Kan85] Suk Kang. A note on the equivalence of specification tests in the two-factor multivariate variance components model. Journal of Econometrics, 28(2):193–203, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901198.

Kopp:1982:DFC

[KD82] Raymond J. Kopp and W. Erwin Diewert. The decomposition of frontier cost function deviations into measures of technical and allocative efficiency. Journal of Econometrics, 19(2–3):319– 331, August 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900082.

Kelejian:1982:EST

[Kel82] Harry H. Kelejian. An extension of a standard test for heteroskedasticity to a systems framework. Journal of Econometrics, 20(2):325–333, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900239.

Kennan:1985:DCS

[Ken85] John Kennan. The duration of contract strikes in U.S. manufacturing. Journal of Econometrics, 28(1):5–28, April 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900648.

King:1984:APT

[KG84] M. L. King and D. E. A. Giles. Autocorrelation pre-testing in the linear model: Estimation, testing and prediction. Journal of Econometrics, 25(1–2):35–48, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900356.

Kloek:1988:EI

[KH88] Teun Kloek and Yoel Haitovsky. Editors’ introduction. Journal of Econometrics, 37(1):1–6, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768890070X.

Khazzoom:1989:NAN

[Kha89] J. Daniel Khazzoom. A note on the application of the nonlinear two-stage least-squares estimator to a Box–Cox-transformed model. Journal of Econometrics, 42(3):377–379, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407689900602.

Kiefer:1980:EFE

[Kie80] Nicholas M. Kiefer. Estimation of fixed effect models for time series of cross-sections with arbitrary intertemporal covariance. Journal of Econometrics, 14(2):195–202, October 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900901.

Kiefer:1982:IRL

[Kie82] Nicholas M. Kiefer. Identifying restrictions in limited information analysis of the schooling coefficient in a wage equation. Journal of Econometrics, 18(2):219–237, February 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900379.

Kiefer:1984:MEN

[Kie84] Nicholas M. Kiefer. Microeconometric evidence on the neoclassical model of demand. Journal of Econometrics, 25(3):285–302, July 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900034.

Kiefer:1985:EI

[Kie85a] Nicholas M. Kiefer. Editor’s introduction. Journal of Econometrics, 28(1):1–3, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900636.

Kiefer:1985:SDB

[Kie85b] Nicholas M. Kiefer. Specification diagnostics based on Laguerre alternatives for econometric models of duration. Journal of Econometrics, 28(1):135–154, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900727.

King:1981:ADW

[Kin81] M. L. King. The alternative Durbin–Watson test: An assessment of Durbin and Watson’s choice of test statistic. Journal of Econometrics, 17(1):51–66, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900580.

King:1983:DWT

[Kin83a] Maxwell L. King. The Durbin–Watson test for serial correlation: Bounds for regressions using monthly data. Journal of Econometrics, 21(3):357–366, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900507.

King:1983:TAA

[Kin83b] Maxwell L. King. Testing for autoregressive against moving average errors in the linear regression model. Journal of Econometrics, 21(1):35–51, January 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901185.

King:1984:NTF

[Kin84] Maxwell L. King. A new test for fourth-order autoregressive disturbances. Journal of Econometrics, 24(3):269–277, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900538.

King:1985:POT

[Kin85] Maxwell L. King. A point optimal test for autoregressive disturbances. Journal of Econometrics, 27(1):21–37, January 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900429.

King:1989:TFO

[Kin89] Maxwell L. King. Testing for fourth-order autocorrelation in regression disturbances when first-order autocorrrelation is present. Journal of Econometrics, 41(3):285–301, July 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900638.

Kiviet:1985:MST

[Kiv85] Jan F. Kiviet. Model selection test procedures in a single linear equation of a dynamic simultaneous system and their defects in small samples. Journal of Econometrics, 28(3):327–362, June 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900041.

Klein:1988:SAE

[Kle88a] Lawrence R. Klein. The statistical approach to economics. Journal of Econometrics, 37(1):7–26, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900711.

Klepper:1988:BEM

[Kle88b] Steven Klepper. Bounding the effects of measurement error in regressions involving dichotomous variables. Journal of Econometrics, 37(3):343–359, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900103.

Klepper:1988:RDC

[Kle88c] Steven Klepper. Regressor diagnostics for the classical errors-in-variables model. Journal of Econometrics, 37(2):225–250, February 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900048.

Kohler:1984:RRT

[KM84] Daniel F. Kohler and Bridger M. Mitchell. Response to residential time-of-use electricity rates: How transferable are the findings? Journal of Econometrics, 26(1–2):141–177, September/October 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900162.

Knight:1980:CDS

[Kni80] John L. Knight. The coefficient of determination and simultaneous equation systems. Journal of Econometrics, 14(2):265– 270, October 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900962.

Knight:1985:MOW

[Kni85] John L. Knight. The moments of OLS and 2SLS when the disturbances are non-normal. Journal of Econometrics, 27(1):39– 60, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900430.

Koenker:1981:NST

[Koe81] Roger Koenker. A note on Studentizing a test for heteroscedasticity. Journal of Econometrics, 17(1):107–112, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900622.

Kohn:1982:WAT

[Koh82] Robert Kohn. When is an aggregate of a time series efficiently forecast by its past? Journal of Econometrics, 18(3):337–349, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900872.

Krasker:1987:BEP

[KP87] William S. Krasker and John W. Pratt. Bounding the effects of proxy variables on instrumental-variables coefficients. Journal of Econometrics, 35(2–3):233–252, July 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900261.

Krasker:1981:RBI

[Kra81] William S. Krasker. The role of bounded-influence estimation in model selection. Journal of Econometrics, 16(1):131–138, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900804.

Kramer:1985:PDW

[Krä85] W. Krämer. The power of the Durbin–Watson test for regressions without an intercept. Journal of Econometrics, 28(3):363– 370, June 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900053.

Krause:2013:CEL

[Kra13] Melanie Krause. Corrigendum to “Elliptical Lorenz Curves” [J. Econom. 40 (1989) 327–338]. Journal of Econometrics, 174(1):44, May 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000158. See [VA89].

Kroch:1988:BSE

[Kro88] Eugene Kroch. Bounds on specification error arising from data proxies. Journal of Econometrics, 37(1):171–192, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900814.

King:1986:JOS

[KS86] Maxwell L. King and Murray D. Smith. Joint one-sided tests of linear regression coefficients. Journal of Econometrics, 32(3):367– 383, August 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900205.

Kulatilaka:1985:TVS

[Kul85] Nalin Kulatilaka. Tests on the validity of static equilibrium models. Journal of Econometrics, 28(2):253–268, May 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901241.

Kumbhakar:1987:STA

[Kum87] Subal C. Kumbhakar. The specification of technical and allocative inefficiency in stochastic production and profit frontiers. Journal of Econometrics, 34(3):335–348, March 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900169.

Kooiman:1985:LDB

[KVT85] Peter Kooiman, Herman K. Van Dijk, and A. Roy Thurik. Likelihood diagnostics and Bayesian analysis of a microeconomic disequilibrium model for retail services. Journal of Econometrics, 29(1–2):121–148, July/August 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900351.

Keller:1983:MMQ

[KW83] Wouter J. Keller and Tom Wansbeek. Multivariate methods for quantitative and qualitative data. Journal of Econometrics, 22(1–2):91–111, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683900957.

Lancaster:1985:GRH

[Lan85a] Tony Lancaster. Generalised residuals and heterogeneous duration models: With applications to the weilbull model. Journal of Econometrics, 28(1):155–169, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900739.

Lancaster:1985:SEM

[Lan85b] Tony Lancaster. Simultaneous equations models in applied search theory. Journal of Econometrics, 28(1):113–126, April 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900703.

Lankford:1985:PCP

[Lan85c] R. Hamilton Lankford. Preferences of citizens for public expenditures on elementary and secondary education. Journal of Econometrics, 27(1):1–20, January 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900417.

Lawrence:1987:ACR

[Law87] David B. Lawrence. The application of censored regression techniques to the analysis of tax reform. Journal of Econometrics, 35(2–3):287–302, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900297.

Lee:1986:STW

[LC86] Lung-Fei Lee and Andrew Chesher. Specification testing when score test statistics are identically zero. Journal of Econometrics, 31(2):121–149, March 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768690045X.

Lahiri:1981:JET

[LE81] Kajal Lahiri and Daniel Egy. Joint estimation and testing for functional form and heteroskedasticity. Journal of Econometrics, 15(2):299–307, February 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901196.

Lee:1981:FRP

[Lee81] Lung-Fei Lee. Fully recursive probability models and multivariate log-linear probability models for the analysis of qualitative data. Journal of Econometrics, 16(1):51–69, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900750.

Lee:1982:SEM

[Lee82a] Lung-Fei Lee. Specification error in multinomial logit models: Analysis of the omitted variable bias. Journal of Econometrics, 20(2):197–209, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900197.

Lee:1982:TNE

[Lee82b] Lung-Fei Lee. Test for normality in the econometric disequilibrium markets model. Journal of Econometrics, 19(1):109–123, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900537.

Lee:1983:MLE

[Lee83a] Lung-Fei Lee. On maximum likelihood estimation of stochastic frontier production models. Journal of Econometrics, 23(2):269– 274, October 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390081F.

Lee:1983:TDA

[Lee83b] Lung-Fei Lee. A test for distributional assumptions for the stochastic frontier functions. Journal of Econometrics, 22(3):245– 267, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901021.

Lee:1984:MLE

[Lee84] Lung-Fei Lee. Maximum likelihood estimation and a specification test for non-normal distributional assumption for the accelerated failure time models. Journal of Econometrics, 24(1–2):159–179, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900770.

Lee:1986:SMM

[Lee86] Lung-Fei Lee. The specification of multi-market disequilibrium econometric models. Journal of Econometrics, 32(3):297–332, August 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900187.

Lee:1987:NPT

[Lee87] Lung-Fei Lee. Non-parametric testing of discrete panel data models. Journal of Econometrics, 34(1–2):147–177, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900716.

Levine:1983:RSC

[Lev83] David Levine. A remark on serial correlation in maximum likelihood. Journal of Econometrics, 23(3):337–342, December 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407683900623.

Levine:1985:SMM

[Lev85] David Levine. The sensitivity of MLE to measurement error. Journal of Econometrics, 28(2):223–230, May 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901216.

Levine:1986:RRL

[Lev86] David K. Levine. Reverse regression for latent-variable models. Journal of Econometrics, 32(2):291–292, July 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900424.

Lewbel:1987:FDS

[Lew87] Arthur Lewbel. Fractional demand systems. Journal of Econometrics, 36(3):311–337, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900054.

Liu:1981:BAT

[LH81] Lon-Mu Liu and Dominique M. Hanssens. A Bayesian approach to time-varying cross-sectional regression models. Journal of Econometrics, 15(3):341–356, April 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900993.

Liddle:1988:SSA

[LM88] Roger F. Liddle and John F. Monahan. A stationary stochastic approximation method. Journal of Econometrics, 38(1–2):91–102, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900280.

Lo:1989:SPV

[LM89] Andrew W. Lo and A. Craig MacKinlay. The size and power of the variance ratio test in finite samples: A Monte Carlo investigation. Journal of Econometrics, 40(2):203–238, February 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900833.

Lo:1986:LVD

[Lo86] Andrew W. Lo. Logit versus discriminant analysis: A specification test and application to corporate bankruptcies. Journal of Econometrics, 31(2):151–178, March 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900461.

Lee:1987:MMR

[LP87] Lung-Fei Lee and Mark M. Pitt. Microeconometric models of rationing, imperfect markets, and non-negativity constraints. Journal of Econometrics, 36(1–2):89–110, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900455.

Lubinsky:1988:DAS

[LP88] David Lubinsky and Daryl Pregibon. Data analysis as search. Journal of Econometrics, 38(1–2):247–268, May/June 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900358.

Liu:1980:RCF

[LT80] Lon-Mu Liu and George C. Tiao. Random coefficient firstorder autoregressive models. Journal of Econometrics, 13(3):305– 325, August 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900822.

Lubrano:1985:BAS

[Lub85a] Michel Lubrano. Bayesian analysis of switching regression models. Journal of Econometrics, 29(1–2):69–95, July/August 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900338.

Lubrano:1985:SAP

[Lub85b] Michel Lubrano. Some aspects of prior elicitation problems in disequilibrium models. Journal of Econometrics, 29(1–2):165–172, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900375.

Lutkepohl:1981:MNN

[Lüt81] Helmut Lütkepohl. A model for non-negative and non-positive distributed lag functions. Journal of Econometrics, 16(2):211– 219, June 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901081.

Lutkepohl:1982:NCD

[Lüt82] Helmut Lütkepohl. Non-causality due to omitted variables. Journal of Econometrics, 19(2–3):367–378, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900112.

Lutkepohl:1984:LTV

[Lüt84] Helmut Lütkepohl. Linear transformations of vector ARMA processes. Journal of Econometrics, 26(3):283–293, December 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768490023X.

Lutkepohl:1988:PTS

[Lüt88] Helmut Lütkepohl. Prediction tests for structural stability. Journal of Econometrics, 39(3):267–296, November 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900590.

Lutkepohl:1989:NAD

[Lüt89] Helmut Lütkepohl. A note on the asymptotic distribution of impulse response functions of estimated var models with orthogonal residuals. Journal of Econometrics, 42(3):371–376, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407689900596.

Lien:1987:SBL

[LV87] Donald Lien and Quang H. Vuong. Selecting the best linear regression model: A classical approach. Journal of Econometrics, 35 (1):3–23, May 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900789.

Lee:1987:FEI

[LWD87] Cheng F. Lee, Chunchi Wu, and Mohamed Djarraya. A further empirical investigation of the dividend adjustment process. Journal of Econometrics, 35(2–3):267–285, July 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900285.

Lerman:1989:IAE

[LY89] Robert I. Lerman and Shlomo Yitzhaki. Improving the accuracy of estimates of Gini coefficients. Journal of Econometrics, 42(1):43–47, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900742.

Lynch:1985:SDY

[Lyn85] Lisa M. Lynch. State dependency in youth unemployment: A lost generation? Journal of Econometrics, 28(1):71–84, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900673.

Maasoumi:1980:RLM

[Maa80] Esfandiar Maasoumi. A ridge-like method for simultaneous estimation of simultaneous equations. Journal of Econometrics, 12(2):161–176, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900044.

Maasoumi:1986:RFE

[Maa86] Esfandiar Maasoumi. Reduced form estimation and prediction from uncertain structural models: A generic approach.

Journal of Econometrics, 31(1):3–29, February 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900539.

Maasoumi:1989:CDA

[Maa89] Esfandiar Maasoumi. Continuously distributed attributes and measures of multivariate inequality. Journal of Econometrics, 42(1):131–144, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768990081X.

MaCurdy:1982:UTS

[MaC82] Thomas E. MaCurdy. The use of time series processes to model the error structure of earnings in a longitudinal data analysis. Journal of Econometrics, 18(1):83–114, January 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900963.

Maddala:1981:EI

[Mad81] G. S. Maddala. Editor’s introduction. Journal of Econometrics, 16(1):1, May 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0304407681900701.

Maeshiro:1980:NES

[Mae80] Asatoshi Maeshiro. New evidence on the small properties of estimators of sur models with autocorrelated disturbances. Journal of Econometrics, 12(2):177–187, February 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900056.

Magnus:1982:MEC

[Mag82] Jan R. Magnus. Multivariate error components analysis of linear and nonlinear regression models by maximum likelihood. Journal of Econometrics, 19(2–3):239–285, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900057.

Magee:1985:EIE

[Mag85] Lonnie Magee. Efficiency of iterative estimators in the regression model with AR(1) disturbances. Journal of Econometrics, 29(3):275–287, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901563.

Magee:1987:NCO

[Mag87] Lonnie Magee. A note on Cochrane-orcutt estimation. Journal of Econometrics, 35(2–3):211–218, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900248.

Mandy:1984:MPT

[Man84] D. M. Mandy. The moments of a pre-test estimator under possible heteroscedasticity. Journal of Econometrics, 25(1–2):29–33, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900344.

Manski:1985:SAD

[Man85] Charles F. Manski. Semiparametric analysis of discrete response: Asymptotic properties of the maximum score estimator. Journal of Econometrics, 27(3):313–333, March 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900090.

Manski:1986:SAB

[Man86] Charles F. Manski. Semiparametric analysis of binary response from response-based samples. Journal of Econometrics, 31(1):31– 40, February 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900540.

Maravall:1980:EAS

[Mar80] Agustin Maravall. Effects of alternative seasonal adjustment procedures on monetary policy. Journal of Econometrics, 14(1):115– 136, September 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304407680900767. See comments

[Sto80].

Maravall:1981:NIM

[Mar81] Agustin Maravall. A note on identification of multivariate timeseries models. Journal of Econometrics, 16(2):237–247, June 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768190110X.

Marsh:1985:NLS

[Mar85] Terry A. Marsh. On non-linear serial dependencies in stock returns. Journal of Econometrics, 30(1–2):289–296, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901411.

Masson:1985:VIS

[Mas85] Edwina A. Masson. The value of imperfect sample separation information in mixtures of normal distributions. Journal of Econometrics, 29(3):341–350, September 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901605.

Mayer:1989:EDM

[May89] Walter J. Mayer. Estimating disequilibrium models with limited a priori price-adjustment information. Journal of Econometrics, 41(3):303–320, July 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768990064X.

Morrison:1981:SRL

[MB81] C. J. Morrison and E. R. Berndt. Short-run labor productivity in a dynamic model. Journal of Econometrics, 16(3):339– 365, August 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900348.

McCarthy:1981:NMP

[McC81] Michael D. McCarthy. A note on the moments of partially restricted reduced forms. Journal of Econometrics, 17(3):383–387, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900087.

McCabe:1989:MTE

[McC89] B. P. M. McCabe. Misspecification tests in econometrics based on ranks. Journal of Econometrics, 40(2):261–278, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900857.

McFadden:1987:RBS

[McF87] Daniel McFadden. Regression-based specification tests for the multinomial logit model. Journal of Econometrics, 34(1–2): 63–82, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900674.

McDonald:1981:LSE

[MD81] John McDonald and John Darroch. Large sample estimation and testing procedures for dynamic equation systems. Journal of Econometrics, 17(1):127–130, September 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900658.

McDonald:1983:CEE

[MD83] John McDonald and John Darroch. Consistent estimation of equations with composite moving average disturbance terms. Journal of Econometrics, 23(2):253–267, October 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390080O.

Manning:1987:MCE

[MDR87] W. G. Manning, N. Duan, and W. H. Rogers. Monte Carlo evidence on the choice between sample selection and two-part models. Journal of Econometrics, 35(1):59–82, May 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900819.

Meese:1980:DFD

[Mee80] Richard Meese. Dynamic factor demand schedules for labor and capital under rational expectations. Journal of Econometrics, 14(1):141–158, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900780.

McAleer:1982:TSR

[MF82] Michael McAleer and Gordon Fisher. Testing separate regression models subject to specification error. Journal of Econometrics, 19(1):125–145, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900549.

Maloney:1980:FVM

[MI80] M. T. Maloney and M. E. Ireland. Fiscal versus monetary policy: An application of transfer functions. Journal of Econometrics, 13(2):253–266, June 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900172.

Missiakoulis:1983:SDW

[Mis83]  Spyros Missiakoulis. Sargan densities which one?  Jour-

nal of Econometrics, 23(2):223–233, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390078J.

Mittelhammer:1984:RLS

[Mit84] R. C. Mittelhammer. Restricted least squares, pre-test, OLS and Stein rule estimators: Risk comparisons under model misspecification. Journal of Econometrics, 25(1–2):151–164, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900435.

Mittelhammer:1985:QRD

[Mit85] Ron C. Mittelhammer. Quadratic risk domination of restricted least squares estimators via Stein-ruled auxiliary constraints. Journal of Econometrics, 29(3):289–303, September 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901575.

Maasoumi:1988:CGO

[MJ88] Esfandiar Maasoumi and Jin-Ho Jeong. A comparison of GRF and other reduced-form estimators in simultaneous equations models. Journal of Econometrics, 37(1):115–134, January 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900784.

Mouchart:1986:NPA

[MO86] M. Mouchart and R. Orsi. A note on price adjustment models in disequilibrium econometrics. Journal of Econometrics, 31(2): 209–217, March 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900485.

Moffitt:1985:UID

[Mof85] Robert Moffitt. Unemployment insurance and the distribution of unemployment spells. Journal of Econometrics, 28(1):85–101, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900685.

Monahan:1983:FBA

[Mon83] John F. Monahan. Fully Bayesian analysis of ARMA time series models. Journal of Econometrics, 21(3):307–331, April 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900489.

Morey:1984:SIP

[Mor84] M. J. Morey. The statistical implications of preliminary specification error testing. Journal of Econometrics, 25(1–2):63–72, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768490037X.

Morrison:1986:PMN

[Mor86] Catherine J. Morrison. Productivity measurement with nonstatic expectations and varying capacity utilization: An integrated approach. Journal of Econometrics, 33(1–2):51–74, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900278.

Moulton:1986:RGE

[Mou86] Brent R. Moulton. Random group effects and the precision of regression estimates. Journal of Econometrics, 32(3):385–397, August 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900217.

Maasoumi:1982:BII

[MP82] Esfandiar Maasoumi and Peter C. B. Phillips. On the behavior of inconsistent instrumental variable estimators. Journal of Econometrics, 19(2–3):183–201, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768290001X.

Madan:1989:NEN

[MP89a] Dilip B. Madan and Ingmar R. Prucha. A note on the estimation of nonsymmetric dynamic factor demand models. Journal of Econometrics, 42(2):275–283, October 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900079.

Magnus:1989:EMP

[MP89b] Jan R. Magnus and Bahram Pesaran. The exact multi-period mean-square forecast error for the first-order autoregressive model with an intercept. Journal of Econometrics, 42(2):157–179, October 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900018.

McAleer:1986:FRS

[MPV86] Michael McAleer, Adrian Pagan, and Ignazio Visco. A further result on the sign of restricted least-squares estimates. Journal of Econometrics, 32(2):287–290, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900412.

McDonald:1981:ABG

[MR81a] James B. McDonald and Michael R. Ransom. An analysis of the bounds for the Gini coefficient. Journal of Econometrics, 17(2):177–188, November 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681900257.

Melton:1981:IAE

[MR81b] William C. Melton and V. Vance Roley. Imperfect asset elasticities and financial model building. Journal of Econometrics, 15(1):139–154, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407681901330.

Morimune:1980:DRC

[MS80] Kimio Morimune and Takamitsu Sawa. Decision rules for the choice of structural equations. Journal of Econometrics, 14(3):329–347, December 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900317.

Mitchell:1986:SFD

[MS86] Douglas W. Mitchell and Paul J. Speaker. A simple, flexible distributed lag technique: The polynomial inverse lag. Journal of Econometrics, 31(3):329–340, April 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900643.

Maddala:1981:AFN

[MT81] G. S. Maddala and R. P. Trost. Alternative formulations of the Nerlove–Press models. Journal of Econometrics, 16(1):35–49, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900749.

Morimune:1984:AED

[MT84] Kimio Morimune and Yoshihiko Tsukuda. Asymptotic expansions of the distributions of the structural variance estimators in a simultaneous equations system. Journal of Econometrics, 24(3): 279–292, March 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768490054X.

Manski:1986:OCM

[MT86] Charles F. Manski and T. Scott Thompson. Operational characteristics of maximum score estimation. Journal of Econometrics, 32(1):85–108, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407686900138.

Manski:1989:EBP

[MT89] Charles F. Manski and T. Scott Thompson. Estimation of best predictors of binary response. Journal of Econometrics, 40(1):97– 123, January 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900328.

Muellbauer:1981:EDB

[Mue81] John Muellbauer. Are employment decisions based on rational expectations? Journal of Econometrics, 16(1):156, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900865.

Mullahy:1986:STS

[Mul86] John Mullahy. Specification and testing of some modified count data models. Journal of Econometrics, 33(3):341–365, December 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900023.

Mundlak:1981:CNS

[Mun81] Yair Mundlak. On the concept of non-significant functions and its implications for regression analysis. Journal of Econometrics, 16(1):139–149, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900816.

Muthen:1983:LVS

[Mut83] Bengt Muthén. Latent variable structural equation modeling with categorical data. Journal of Econometrics, 22(1–2):43–65, May/June 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900933.

Morgenthaler:1986:CBS

[MV86] S. Morgenthaler and Y. Vardi. Choice-based samples: A nonparametric approach. Journal of Econometrics, 32(1):109–125, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768690014X.

MacKinnon:1985:SHC

[MW85] James G. MacKinnon and Halbert White. Some heteroskedasticityconsistent covariance matrix estimators with improved finite sample properties. Journal of Econometrics, 29(3):305–325, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901587.

MacKinnon:1983:TMS

[MWD83] James G. MacKinnon, Halbert White, and Russell Davidson. Tests for model specification in the presence of alternative hypotheses: Some further results. Journal of Econometrics, 21(1): 53–70, January 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901197.

Nelson:1984:ETS

[Nel84] Forrest D. Nelson. Efficiency of the two-step estimator for models with endogenous sample selection. Journal of Econometrics, 24(1–2):181–196, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900782.

Newey:1985:GMM

[New85] Whitney K. Newey. Generalized method of moments specification testing. Journal of Econometrics, 29(3):229–256, September 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768590154X.

Newey:1986:LIV

[New86] Whitney K. Newey. Linear instrumental variable estimation of limited dependent variable models with endogenous explanatory variables. Journal of Econometrics, 32(1):127–141, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900151.

Newey:1987:EEL

[New87a] Whitney K. Newey. Efficient estimation of limited dependent variable models with endogenous explanatory variables.

Journal of Econometrics, 36(3):231–250, November 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900017.

Newey:1987:STD

[New87b] Whitney K. Newey. Specification tests for distributional assumptions in the Tobit model. Journal of Econometrics, 34(1–2): 125–145, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900704.

Newey:1988:AER

[New88] Whitney K. Newey. Adaptive estimation of regression models via moment restrictions. Journal of Econometrics, 38(3):301– 339, July 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900486.

Nickelsburg:1985:SSP

[Nic85] Gerald Nickelsburg. Small-sample properties of dimensionality statistics for fitting VAR models to aggregate economic data: A Monte Carlo study. Journal of Econometrics, 28(2):183–192, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901186.

Nakamura:1985:DML

[NN85a] Alice Nakamura and Masao Nakamura. Dynamic models of the labor force behavior of married women which can be estimated using limited amounts of past information. Journal of Econometrics, 27(3):273–298, March 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900077.

Nakamura:1985:PTW

[NN85b] Alice Nakamura and Masao Nakamura. On the performance of tests by Wu and by Hausman for detecting the ordinary least squares bias problem. Journal of Econometrics, 29(3):213–227, September 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901538.

Narendranathan:1985:MPJ

[NN85c] Wiji Narendranathan and Stephen Nickell. Modelling the process of job search. Journal of Econometrics, 28(1):29–49, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768590065X.

Norman:1981:CSMa

[Nor81a] Alfred L. Norman. On the control of structural models. Journal of Econometrics, 15(1):13–24, January 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407681901251. See comment [Cho81b] and reply [Nor81b].

Norman:1981:CSMb

[Nor81b] Alfred L. Norman. On the control of structural models — reply. Journal of Econometrics, 15(1):29, January 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407681901275. See [Nor81a, Cho81b].

Nowak:1983:IDS

[Now83] Eugen Nowak. Identification of the dynamic shock-error model with autocorrelated errors. Journal of Econometrics, 23(2):211– 221, October 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390077I.

Nowak:1985:GID

[Now85] Eugen Nowak. Global identification of the dynamic shockerror model. Journal of Econometrics, 27(2):211–219, February 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900880.

Nowak:1989:ISE

[Now89] Eugen Nowak. Identification of simultaneous equation models with measurement errors based on time series structure. Journal of Econometrics, 40(2):319–325, February 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900882.

Newhouse:1980:HYC

[NPM80] Joseph P. Newhouse, Charles E. Phelps, and M. Susan Marquis. On having your cake and eating it too: Econometric problems in estimating the demand for health services. Journal of Econometrics, 13(3):365–390, August 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900858.

Nankervis:1985:TAP

[NS85] J. C. Nankervis and N. E. Savin. Testing the autoregressive parameter with the t statistic. Journal of Econometrics, 27(2):143– 161, February 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900843.

Nakervis:1987:C

[NS87] J. C. Nakervis and N. E. Savin. Corrigendum. Journal of Econometrics, 34(3):391, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900212.

Nankervis:1988:EML

[NS88a] J. C. Nankervis and N. E. Savin. The exact moments of the least-squares estimator for the autoregressive model corrections and extensions. Journal of Econometrics, 37(3):381–388, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900127.

Naylor:1988:EIN

[NS88b] J. C. Naylor and A. F. M. Smith. Econometric illustrations of novel numerical integration strategies for Bayesian inference. Journal of Econometrics, 38(1–2):103–125, May/June 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900292.

Nygaard:1989:IIM

[NS89] Fredrik Nygård and Arne Sandström. Income inequality measures based on sample surveys. Journal of Econometrics, 42(1):81–95, September 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900778.

Olsen:1985:CIS

[OF85] Randall J. Olsen and George Farkas. Conception intervals and the substitution of fertility over time. Journal of Econometrics, 28(1): 103–112, April 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900697.

Offenbacher:1980:EMA

[Off80] Edward K. Offenbacher. Economic monetary aggregatescomment. Journal of Econometrics, 14(1):55–56, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440768090072X.

Ohtani:1984:SSP

[OH84] Kazuhiro Ohtani and Yuzo Honda. Small sample properties of the mixed regression estimator. Journal of Econometrics, 26(3):375– 385, December 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900277.

Ohanian:1988:SEU

[Oha88] Lee E. Ohanian. The spurious effects of unit roots on vector autoregressions: A Monte Carlo study. Journal of Econometrics, 39(3):251–266, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900589.

Ohtani:1982:BES

[Oht82] Kazuhiro Ohtani. Bayesian estimation of the switching regression model with autocorrelated errors. Journal of Econometrics, 18(2):251–261, February 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900392.

Ohtani:1987:PDV

[Oht87] Kazuhiro Ohtani. On pooling disturbance variances when the goal is testing restrictions on regression coefficients. Journal of Econometrics, 35(2–3):219–231, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768790025X.

Oldford:1988:OOS

[Old88] R. Wayne Oldford. Object-oriented software representations for statistical data. Journal of Econometrics, 38(1–2):227–246, May/ June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900346.

Over:1980:EAM

[OS80] A. Mead Over and Kenneth R. Smith. The estimation of the ambulatory medical care technology where output is an unobservable variable. Journal of Econometrics, 13(2):225–251, June 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900160.

Olson:1980:MCS

[OSW80] Jerome A. Olson, Peter Schmidt, and Donald M. Waldman. A Monte Carlo study of estimators of stochastic frontier production functions. Journal of Econometrics, 13(1):67–82, May 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900433.

Ohtani:1980:ERC

[OT80] Kazuhiro Ohtani and Toshihisa Toyoda. Estimation of regression coefficients after a preliminary test for homoscedasticity. Journal of Econometrics, 12(2):151–159, February 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900032.

Park:1984:LBC

[PA84] Rolla Edward Park and Jan Paul Acton. Large business customer response to time-of-day electricity rates. Journal of Econometrics, 26(1–2):229–252, September/October 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900198.

Paarsch:1984:MCC

[Paa84] Harry J. Paarsch. A Monte Carlo comparison of estimators for censored regression models. Journal of Econometrics, 24(1–2):197–213, January/February 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900794.

Pagan:1980:SIE

[Pag80] Adrian Pagan. Some identification and estimation results for regression models with stochastically varying coefficients. Journal of Econometrics, 13(3):341–363, August 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900846.

Pagan:1989:RSS

[Pag89] Adrian Pagan. On the role of simulation in the statistical evaluation of econometric models. Journal of Econometrics, 40(1):125– 139, January 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768990033X.

Pal:1980:CME

[Pal80] Manoranjan Pal. Consistent moment estimators of regression coefficients in the presence of errors in variables. Journal of Econometrics, 14(3):349–364, December 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900329.

Palash:1981:AEP

[Pal81] Carl J. Palash. On the accuracy and efficiency of polynomial approximations in optimal macroeconomic policy determination. Journal of Econometrics, 15(1):49–62, January 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901299.

Parks:1980:EML

[Par80] Richard W. Parks. On the estimation of multinomial logit models from relative frequency data. Journal of Econometrics, 13(3):293– 303, August 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900810.

Park:1982:FPU

[Par82a] Soo-Bin Park. A forecasting property of the unrestricted, restricted, and partially restricted reduced-form coefficients. Journal of Econometrics, 19(2–3):385–390, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900136.

Park:1982:SSP

[Par82b] Soo-Bin Park. Some sampling properties of minimum expected loss (MELO) estimators of structural coefficients. Journal of Econometrics, 18(3):295–311, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900859.

Perryman:1981:NI

[Per81] M. Ray Perryman. News item. Journal of Econometrics, 15(2): 309, February 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901214.

Pesaran:1981:IRE

[Pes81a] M. H. Pesaran. Identification of rational expectations models. Journal of Econometrics, 16(3):375–398, August 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900361.

Pesaran:1981:PTNa

[Pes81b] M. H. Pesaran. Pitfalls of testing non-nested hypotheses by the Lagrange multiplier method. Journal of Econometrics, 16(1):158, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900889.

Pesaran:1981:PTNb

[Pes81c] M. H. Pesaran. Pitfalls of testing non-nested hypotheses by the Lagrange multiplier method. Journal of Econometrics, 17(3):323– 331, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900051.

Pesaran:1982:CMT

[Pes82] M. H. Pesaran. On the comprehensive method of testing nonnested regression models. Journal of Econometrics, 18(2):263– 274, February 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900409.

Peters:1989:EMO

[Pet89] Thomas A. Peters. The exact moments of OLS in dynamic regression models with non-normal errors. Journal of Econometrics, 40(2):279–305, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900869.

Pantula:1985:MEB

[PF85] Sastry G. Pantula and Wayne A. Fuller. Mean estimation bias in least squares estimation of autoregressive processes. Journal of Econometrics, 27(1):99–121, January 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900466.

Pagano:1981:FDL

[PH81] Marcello Pagano and Michael J. Hartley. On fitting distributed lag models subject to polynomial restrictions. Journal of Econometrics, 16(2):171–198, June 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681901068.

Park:1983:NBA

[PH83] Choon Y. Park and Russell G. Heikes. A note on Balestra’s (1980) approximate estimator for the first-order moving average process. Journal of Econometrics, 21(3):387–388, April 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900520. See [Bal80].

Phillips:1984:EDE

[Phi84a] P. C. B. Phillips. The exact distribution of exogenous variable coefficient estimators. Journal of Econometrics, 26(3):387–398, December 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900289.

Phillips:1984:EDS

[Phi84b] P. C. B. Phillips. The exact distribution of the Stein-rule estimator. Journal of Econometrics, 25(1–2):123–131, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900411.

Phillips:1986:USR

[Phi86] P. C. B. Phillips. Understanding spurious regressions in econometrics. Journal of Econometrics, 33(3):311–340, December 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900011.

Phillips:1988:CUS

[Phi88] Peter C. B. Phillips. Conditional and unconditional statistical independence. Journal of Econometrics, 38(3):341–348, July 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900498.

Picci:1989:PFA

[Pic89] Giorgio Picci. Parametrization of factor analysis models. Journal of Econometrics, 41(1):17–38, May 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900419.

Pierce:1980:DRM

[Pie80] David A. Pierce. Data revisions with moving average seasonal adjustment procedures. Journal of Econometrics, 14(1):95–114, September 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900755.

Pierce:1981:SEE

[Pie81] David A. Pierce. Sources of error in economic time series. Journal of Econometrics, 17(3):305–321, December 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768190004X.

Poirier:1981:MOM

[PK81] Dale J. Poirier and Steven Klepper. Model occurrence and model selection in panel data sets. Journal of Econometrics, 17(3):333– 350, December 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900063.

Ploberger:1989:NTS

[PKK89] Werner Ploberger, Walter Krämer, and Karl Kontrus. A new test for structural stability in the linear regression model. Journal of Econometrics, 40(2):307–318, February 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900870.

Park:1980:EAE

[PM80] Rolla Edward Park and Bridger M. Mitchell. Estimating the autocorrelated error model with trended data. Journal of Econometrics, 13(2):185–201, June 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900147.

Park:1983:CLT

[PMWA83] Rolla Edward Park, Bridger M. Mitchell, Bruce M. Wetzel, and James H. Alleman. Charging for local telephone calls: How household characteristics affect the distribution of calls in the GTE Illinois experiment *. Journal of Econometrics, 22(3):339– 364, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901094.

Palm:1982:LRU

[PN82] F. C. Palm and T. E. Nijman. Linear regression using both temporally aggregated and temporally disaggregated data. Journal of Econometrics, 19(2–3):333–343, August 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900094.

Pagan:1984:EPP

[PN84] A. R. Pagan and D. F. Nicholls. Estimating predictions, prediction errors and their standard deviations using constructed variables. Journal of Econometrics, 24(3):293–310, March 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900551.

Prucha:1986:CAM

[PN86] Ingmar R. Prucha and M. Ishaq Nadiri. A comparison of alternative methods for the estimation of dynamic factor demand models under non-static expectations. Journal of Econometrics, 33(1–2):187–211, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900333.

Poirier:1980:EUB

[Poi80a] Dale J. Poirier. Experience with using the Box–Cox transformation when forecasting economic time series: A comment. Journal of Econometrics, 14(2):277–280, October 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900986.

Poirier:1980:POB

[Poi80b] Dale J. Poirier. Partial observability in bivariate probit models. Journal of Econometrics, 12(2):209–217, February 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768090007X.

Poirier:1988:CRR

[Poi88] Dale J. Poirier. Causal relationships and replicability. Journal of Econometrics, 39(1–2):213–234, September/October 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900462.

Pollock:1984:TRF

[Pol84] D. S. G. Pollock. Two reduced-form approaches to the derivation of the maximum-likelihood estimators for simultaneousequation systems. Journal of Econometrics, 24(3):331–347, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900575.

Poli:1985:BNP

[Pol85] I. Poli. A Bayesian non-parametric estimate for multivariate regression. Journal of Econometrics, 28(2):171–182, May 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901174.

Powell:1984:LAD

[Pow84] James L. Powell. Least absolute deviations estimation for the censored regression model. Journal of Econometrics, 25(3):303– 325, July 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900046.

Powell:1986:CRQ

[Pow86] James L. Powell. Censored regression quantiles. Journal of Econometrics, 32(1):143–155, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900163.

Potscher:1986:CPA

[PP86] Benedikt M. Pötscher and Ingmar R. Prucha. A class of partially adaptive one-step m-estimators for the nonlinear regression model with dependent observations. Journal of Econometrics, 32(2):219–251, July 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900394.

Poirier:1981:AES

[PR81] Dale J. Poirier and Paul A. Ruud. On the appropriateness of endogenous switching. Journal of Econometrics, 16(2):249–256, June 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901111.

Priestley:1988:CDT

[Pri88] M. B. Priestley. Current developments in time series modelling. Journal of Econometrics, 37(1):67–86, January 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900760.

Pratt:1981:NDS

[PS81] John W. Pratt and Robert Schlaifer. On the nature and discovery of structure. Journal of Econometrics, 16(1):154, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407681900841.

Pole:1985:BAS

[PS85a] A. M. Pole and A. F. M. Smith. A Bayesian analysis of some threshold switching models. Journal of Econometrics, 29(1–2):97–119, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440768590034X.

Powell:1985:ECA

[PS85b] James L. Powell and Thomas M. Stoker. The estimation of complete aggregation structures. Journal of Econometrics, 30(1–2):317–344, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901447.

Pratt:1988:IOL

[PS88] John W. Pratt and Robert Schlaifer. On the interpretation and observation of laws. Journal of Econometrics, 39(1–2): 23–52, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407688900395.

Penm:1984:MSA

[PT84] J. H. W. Penm and R. D. Terrell. Multivariate subset autoregressive modelling with zero constraints for detecting ‘overall causality’. Journal of Econometrics, 24(3):311–330, March 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900563.

Parks:1984:MCW

[PW84] Richard W. Parks and David Weitzel. Measuring the consumer welfare effects of time-differentiated electricity prices. Journal of Econometrics, 26(1–2):35–64, September/October 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900125.

Palm:1980:LSE

[PZ80] Franz Palm and Arnold Zellner. Large sample estimation and testing procedures for dynamic equation systems. Journal of Econometrics, 12(3):251–283, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900561.

Palm:1981:LSE

[PZ81] Franz Palm and Arnold Zellner. Large sample estimation and testing procedures for dynamic equation systems. Journal of Econometrics, 17(1):131–138, September 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768190066X.

Parke:1985:SCR

[PZ85] Darrel W. Parke and Janice Zagardo. Stochastic coefficient regression estimates of the sources of shifts into MMDA deposits using cross-section data. Journal of Econometrics, 29(3):327–340, September 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901599.

Ransom:1987:CCD

[Ran87] Michael R. Ransom. A comment on consumer demand systems with binding non-negativity constraints. Journal of Econometrics, 34(3):355–359, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900182.

Rao:1982:NUS

[Rao82] U. L. Gouranga Rao. A note on the unbiasedness of Swamy’s estimator for the random coefficient regression model. Journal of Econometrics, 18(3):395–401, April 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900902.

Roehrig:1984:OCR

[Roe84] C. S. Roehrig. Optimal critical regions for pre-test estimators using a Bayes risk criterion. Journal of Econometrics, 25(1–2):3– 14, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900320.

Rogers:1986:MLM

[Rog86] Alan J. Rogers. Modified Lagrange multiplier tests for problems with one-sided alternatives. Journal of Econometrics, 31(3):341– 361, April 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407686900655.

Rossi:1985:CAF

[Ros85] Peter E. Rossi. Comparison of alternative functional forms in production. Journal of Econometrics, 30(1–2):345–361, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901459.

Rosenzweig:1987:FIH

[RS87] Mark R. Rosenzweig and T. Paul Schultz. Fertility and investments in human capital: Estimates of the consequence of imperfect fertility control in Malaysia. Journal of Econometrics, 36(1–2):163–184, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900480.

Richard:1988:BAS

[RS88] J. F. Richard and M. F. J. Steel. Bayesian analysis of systems of seemingly unrelated regression equations under a recursive extended natural conjugate prior density. Journal of Econometrics, 38(1–2):7–37, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900255.

Raj:1980:EER

[RSU80] Baldev Raj, Virender Kumar Srivastava, and Sushama Upadhyaya. The efficiency of estimating a random coefficient model. Journal of Econometrics, 12(3):285–299, April 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900573.

Raveh:1980:FCS

[RT80a] Adi Raveh and Charles S. Tapiero. Finding common seasonal patterns among time series: An MDS approach. Journal of Econometrics, 12(3):353–363, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900615.

Richard:1980:EPD

[RT80b] J.-F. Richard and H. Tompa. On the evaluation of poly-t density functions. Journal of Econometrics, 12(3):335–351, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900603.

Ruud:1986:CEL

[Ruu86] Paul A. Ruud. Consistent estimation of limited dependent variable models despite misspecification of distribution. Journal of Econometrics, 32(1):157–187, June 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900175.

Rivers:1988:LIE

[RV88] Douglas Rivers and Quang H. Vuong. Limited information estimators and exogeneity tests for simultaneous probit models. Journal of Econometrics, 39(3):347–366, November 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900632.

Rhodes:1981:SED

[RW81] George F. Rhodes and M. Daniel Westbrook. A study of estimator densities and performance under misspecification. Journal of Econometrics, 16(3):311–337, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900336.

Silver:1989:TSS

[SA89] J. Lew Silver and Mukhtar M. Ali. Testing Slutsky symmetry in systems of linear demand equations. Journal of Econometrics, 41(2):251–266, June 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900961.

Saikkonen:1989:ARE

[Sai89] Pentti Saikkonen. Asymptotic relative efficiency of the classical test statistics under misspecification. Journal of Econometrics, 42(3):351–369, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407689900584.

Salazar:1982:SCT

[Sal82] Diego Salazar. Structural changes in time series models. Journal of Econometrics, 19(1):147–163, May 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900550.

Salemi:1986:SEL

[Sal86] Michael K. Salemi. Solution and estimation of linear rational expectations models. Journal of Econometrics, 31(1):41–66, February 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900552.

Sargan:1981:IMA

[Sar81] J. D. Sargan. Identification in models with autoregressive errors. Journal of Econometrics, 16(1):160–161, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900907.

Sato:1985:IPI

[Sat85] Ryuzo Sato. The invariance principle and income-wealth conservation laws: Application of Lie groups and related transformations. Journal of Econometrics, 30(1–2):365–389, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901472.

Slottje:1989:ERB

[SBN89] D. J. Slottje, R. L. Basmann, and M. Nieswiadomy. On the empirical relationship between several well-known inequality measures. Journal of Econometrics, 42(1):49–66, September 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900754.

Schmidt:1983:NFE

[Sch83] Peter Schmidt. A note on a fixed effect model with arbitrary interpersonal covariance. Journal of Econometrics, 22(3):391– 393, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901112.

Schmidt:1988:EFE

[Sch88] Peter Schmidt. Estimation of a fixed-effect Cobb–Douglas system using panel data. Journal of Econometrics, 37(3):361–380, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900115.

Sickles:1986:ADR

[SGJ86] Robin C. Sickles, David Good, and Richard L. Johnson. Allocative distortions and the regulatory transition of the U.S. airline industry. Journal of Econometrics, 33(1–2):143–163, October/ November 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768690031X.

Hwang:1980:TDM

[sH80] Hae shin Hwang. A test of a disequilibrium model. Journal of Econometrics, 12(3):319–333, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900597.

Sharma:1985:EAA

[Sha85] Subhash C. Sharma. The effects of autocorrelation among errors on the consistency property of OLS variance estimator. Journal of Econometrics, 27(3):335–361, March 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900107.

Shaw:1988:SSR

[Sha88] Daigee Shaw. On-site samples’ regression: Problems of nonnegative integers, truncation, and endogenous stratification. Journal of Econometrics, 37(2):211–223, February 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900036.

Sheehan:1985:SOP

[She85] Dennis P. Sheehan. Second-order properties of estimators of serial correlation from regression residuals. Journal of Econometrics, 27(3):299–311, March 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900089.

Shiller:1981:ATR

[Shi81] Robert J. Shiller. Alternative tests of rational expectations models: The case of the term structure. Journal of Econometrics, 16 (1):71–87, May 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0304407681900762.

Shively:1988:ATR

[Shi88] Thomas S. Shively. An analysis of tests for regression coefficient stability. Journal of Econometrics, 39(3):367–386, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900644.

Simon:1988:COC

[SI88] Herbert A. Simon and Yumi Iwasaki. Causal ordering, comparative statics, and near decomposability. Journal of Econometrics, 39(1–2):149–173, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900437.

Sickles:1985:NME

[Sic85] Robin C. Sickles. A nonlinear multivariate error components analysis of technology and specific factor productivity growth with an application to the U.S. airlines. Journal of Econometrics, 27(1): 61–78, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900442.

Singer:1982:ANS

[Sin82] Burton Singer. Aspects of non-stationarity. Journal of Econometrics, 18(1):169–190, January 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682900987.

Singleton:1985:TSE

[Sin85] Kenneth J. Singleton. Testing specifications of economic agents’ intertemporal optimum problems in the presence of alternative models. Journal of Econometrics, 30(1–2):391–413, October/ November 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685901484.

Schmalensee:1986:EPI

[SJ86] Richard Schmalensee and Paul L. Joskow. Estimated parameters as independent variables: An application to the costs of electric generating units. Journal of Econometrics, 31(3):275– 305, April 1986. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768690062X.

Skyrms:1988:PC

[Sky88] Brian Skyrms. Probability and causation. Journal of Econometrics, 39(1–2):53–68, September/October 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900401.

Schmidt:1980:ESP

[SL80] Peter Schmidt and C. A. Knox Lovell. Estimating stochastic production and cost frontiers when technical and allocative inefficiency are correlated. Journal of Econometrics, 13(1):83–100, May 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900445.

Schmidt:1984:STA

[SL84] Peter Schmidt and Tsai-Fen Lin. Simple tests of alternative specifications in stochastic frontier models. Journal of Econometrics, 24(3):349–361, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900587.

Slade:1986:TFP

[Sla86] Margaret E. Slade. Total-factor-productivity measurement when equilibrium is temporary: A Monte Carlo assessment. Journal of Econometrics, 33(1–2):75–95, October/November 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768690028X.

Slade:1989:MSC

[Sla89] Margaret E. Slade. Modelling stochastic and cyclical components of technical change: An application of the Kalman filter. Journal of Econometrics, 41(3):363–383, July 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900675.

Slottje:1989:EIS

[Slo89] Daniel Slottje. Editor’s introduction: The state of empirical work on economic inequality. Journal of Econometrics, 42(1):1– 3, September 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900705.

Swamy:1980:EMP

[SM80] P. A. V. B. Swamy and J. S. Mehta. On the existence of moments of partially restricted reduced form coefficients. Journal of Econometrics, 14(2):183–194, October 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900895.

Swamy:1981:EMP

[SM81] P. A. V. B. Swamy and J. S. Mehta. On the existence of moments of partially restricted reduced form estimators: A comment. Journal of Econometrics, 17(3):389–392, December 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900099.

Smith:1983:MMT

[SM83a] Marlene A. Smith and G. S. Maddala. Multiple model testing for non-nested heteroskedastic censored regression models. Journal of Econometrics, 21(1):71–81, January 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901203.

Swamy:1983:RRE

[SM83b] P. A. V. B. Swamy and J. S. Mehta. Ridge regression estimation of the Rotterdam model. Journal of Econometrics, 22(3):365– 390, August 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683901100.

Smith:1987:TNA

[Smi87] Richard J. Smith. Testing the normality assumption in multivariate simultaneous limited dependent variable models. Journal of Econometrics, 34(1–2):105–123, January/February 1987. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900698.

Schankerman:1986:TSE

[SN86] Mark Schankerman and M. Ishaq Nadiri. A test of static equilibrium models and rates of return to quasi-fixed factors, with an application to the Bell system. Journal of Econometrics, 33(1–2):97–118, October/November 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900291.

Shamseldin:1984:BPR

[SP84] A. A. Shamseldin and S. James Press. Bayesian parameter and reliability estimation for a bivariate exponential distribution parallel sampling. Journal of Econometrics, 24(3):363–378, March 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900599.

Speakes:1981:ISD

[Spe81] Jeffrey K. Speakes. Inference in some disaggregated models with special covariance structure. Journal of Econometrics, 16(2):257– 274, June 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901123.

Srivastava:1983:NMC

[SS83] V. K. Srivastava and A. K. Srivastava. A note on moments of k-class estimators for negative k. Journal of Econometrics, 21(2):257–260, February 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683900180.

Srivastava:1989:BMM

[SS89] M. S. Srivastava and Balvir Singh. Bootstrapping in multiplicative models. Journal of Econometrics, 42(3):287–297, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407689900547.

Swamy:1980:LPE

[ST80] P. A. V. B. Swamy and P. A. Tinsley. Linear prediction and estimation methods for regression models with stationary stochastic coefficients. Journal of Econometrics, 12(2):103–142, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900019.

Sevestre:1985:NAE

[ST85] P. Sevestre and A. Trognon. A note on autoregressive error components models. Journal of Econometrics, 28(2):231–245, May 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685901228.

Startz:1983:TRE

[Sta83] Richard Startz. Testing rational expectations by the use of overidentifying restrictions. Journal of Econometrics, 23(3):343–351, December 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900635.

Stapleton:1984:EVD

[Sta84] David C. Stapleton. Errors-in-variables in demand systems. Journal of Econometrics, 26(3):255–270, December 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900216.

Stevenson:1980:LFG

[Ste80] Rodney E. Stevenson. Likelihood functions for generalized stochastic frontier estimation. Journal of Econometrics, 13(1): 57–66, May 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900421.

Steinberg:1989:IWP

[Ste89] Dan Steinberg. Induced work participation and the returns to experience for welfare women: Evidence from a social experiment. Journal of Econometrics, 41(3):321–340, July 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900651.

Stokes:1980:EAS

[Sto80] Houston H. Stokes. Effects of alternative seasonal adjustment procedures on monetary policy — comment. Journal of Econometrics, 14(1):137–140, September 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900779. See [Mar80].

Stroeker:1983:AEC

[Str83] R. J. Stroeker. Approximations of the eigenvalues of the covariance matrix of a first-order autoregressive process. Journal of Econometrics, 22(3):269–279, August 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683901033.

Swamy:1988:FTT

[SV88] P. A. V. B. Swamy and Peter Von Zur Muehlen. Further thoughts on testing for causality with econometric models. Journal of Econometrics, 39(1–2):105–147, September/October 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900425.

Silver:1980:LRB

[SW80] J. Lew Silver and T. Dudley Wallace. The lag relationship between wholesale and consumer prices: An application of the hatanaka-Wallace procedure. Journal of Econometrics, 12(3): 375–387, April 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900639.

Spiller:1988:ETC

[SW88] Pablo T. Spiller and Robert O. Wood. The estimation of transaction costs in arbitrage models. Journal of Econometrics, 39(3):309–326, November 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407688900619.

Schmidt:1989:PCR

[SW89a] Peter Schmidt and Ann Dryden Witte. Predicting criminal recidivism using ‘split population’ survival time models. Journal of Econometrics, 40(1):141–159, January 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900341.

Stock:1989:IEM

[SW89b] James H. Stock and Mark W. Watson. Interpreting the evidence on money-income causality. Journal of Econometrics, 40(1):161– 181, January 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900353.

Swamy:1980:CEU

[Swa80] P. A. V. B. Swamy. A comparison of estimators for undersized samples. Journal of Econometrics, 14(2):161–181, October 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900883.

Talwar:1983:DSL

[Tal83] Prem P. Talwar. Detecting a shift in location: Some robust tests. Journal of Econometrics, 23(3):353–367, December 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900647.

Tauchen:1985:DTE

[Tau85] George Tauchen. Diagnostic testing and evaluation of maximum likelihood models. Journal of Econometrics, 30(1–2): 415–443, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901496.

Taylor:1980:SSC

[Tay80] William E. Taylor. Small sample considerations in estimation from panel data. Journal of Econometrics, 13(2):203–223, June 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900159.

Taylor:1981:ECO

[Tay81] William E. Taylor. On the efficiency of the Cochrane-orcutt estimator. Journal of Econometrics, 17(1):67–82, September 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681900592.

Turkington:1988:III

[TB88] D. A. Turkington and R. J. Bowden. Identification information and instruments in linear econometric models with rational expectations. Journal of Econometrics, 38(3):361–373, July 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900516.

Terasvirta:1982:UMS

[Ter82] Timo Teräsvirta. Underestimation of mean square error matrix in misspecified linear models. Journal of Econometrics, 18(2):281– 284, February 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900422.

Terasvirta:1987:UPV

[Ter87a] Timo Teräsvirta. Usefulness of proxy variables in linear models with stochastic regressors. Journal of Econometrics, 36(3):377– 382, November 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900091.

Terza:1987:ELM

[Ter87b] Joseph V. Terza. Estimating linear models with ordinal qualitative regressors. Journal of Econometrics, 34(3):275–291, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407687900133.

Tiao:1980:FCA

[TG80] G. C. Tiao and Irwin Guttman. Forecasting contemporal aggregates of multiple time series. Journal of Econometrics, 12(2):219– 230, February 1980. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407680900081.

Tinsley:1981:EDD

[TGF81] P. A. Tinsley, Bonnie Garrett, and Monica Friar. An exposé of disguised deposits. Journal of Econometrics, 15(1):117–137, January 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407681901329.

Theil:1989:DII

[The89] Henri Theil. The development of international inequality 1960– 1985. Journal of Econometrics, 42(1):145–155, September 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900821.

Thomson:1982:SRS

[Tho82] Michael Thomson. Some results on the statistical properties of an inequality constrained least squares estimator in a linear model with two regressors. Journal of Econometrics, 19(2–3):215– 231, August 1982. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407682900033.

Thornton:1987:NEC

[Tho87] Daniel L. Thornton. A note on the efficiency of the Cochrane– Orcutt estimator of the AR(1) regression model. Journal of Econometrics, 36(3):369–376, November 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440768790008X.

Thursby:1987:OGP

[Thu87] Jerry G. Thursby. OLS or GLS in the presence of specification error?: An expected loss approach. Journal of Econometrics, 35(2–3):359–374, July 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900339.

Tishler:1983:ICD

[Tis83] Asher Tishler. The industrial and commercial demand for electricity under time-of-use pricing. Journal of Econometrics, 23(3):369– 384, December 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900659.

Tjostheim:1981:GCM

[Tjø81] Dag Tjøstheim. Granger-causality in multiple time series. Journal of Econometrics, 17(2):157–176, November 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900245.

Tanaka:1984:SDP

[TM84] Katsuto Tanaka and Koichi Maekawa. The sampling distributions of the predictor for an autoregressive model under misspecifications. Journal of Econometrics, 25(3):327–351, July 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900058.

Toyoda:1986:TEB

[TO86] Toshihisa Toyoda and Kazuhiro Ohtani. Testing equality between sets of coefficients after a preliminary test for equality of disturbance variances in two linear regressions. Journal of Econometrics, 31(1):67–80, February 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900564.

Trenkler:1984:PBE

[Tre84] G. Trenkler. On the performance of biased estimators in the linear regression model with correlated or heteroscedastic errors. Journal of Econometrics, 25(1–2):179–190, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900459.

Trivedi:1980:SSC

[Tri80] P. K. Trivedi. Small samples and collateral information: An application of the hyperparameter model. Journal of Econometrics, 12(3):301–318, April 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407680900585.

Trivedi:1987:EI

[Tri87] Pravin K. Trivedi. Editor’s introduction. Journal of Econometrics, 36(1–2):1–6, September/October 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407687900406.

Tsurumi:1982:BAR

[TS82] Hiroki Tsurumi and Tsunemasa Shiba. A Bayesian analysis of a random coefficient model in a simple Keynesian system. Journal of Econometrics, 18(2):239–249, February 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900380.

Tsurumi:1985:STC

[TS85] Hiroki Tsurumi and Neil Sheflin. Some tests for the constancy of regressions under heteroscedasticity. Journal of Econometrics, 27(2):221–234, February 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685900892.

Tse:1982:EAF

[Tse82] Y. K. Tse. Edgeworth approximations in first-order stochastic difference equations with exogenous variables. Journal of Econometrics, 20(2):175–195, November 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900185.

Tse:1987:NSD

[Tse87] Y. K. Tse. A note on sargan densities. Journal of Econometrics, 34(3):349–354, March 1987. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407687900170.

Tinsley:1980:IFA

[TSF80] P. A. Tinsley, P. A. Spindt, and M. E. Friar. Indicator and filter attributes of monetary aggregates: A nit-picking case for disaggregation. Journal of Econometrics, 14(1):61–91, September 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900743.

Tsurumi:1982:BML

[Tsu82] Hiroki Tsurumi. A Bayesian and maximum likelihood analysis of a gradual switching regression in a simultaneous equation framework. Journal of Econometrics, 19(1):165–182, May 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900562.

Tsurumi:1983:UJA

[TT83] Hiroki Tsurumi and Yoshi Tsurumi. U.S.–Japan automobile trade: A Bayesian test of a product life cycle. Journal of Econometrics, 23(2):193–210, October 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390076H.

Turkington:1989:CTC

[Tur89] Darrell A. Turkington. Classical tests for contemporaneously uncorrelated disturbances in the linear simultaneous equations model. Journal of Econometrics, 42(3):299–317, November 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407689900559.

Tinsley:1981:MPA

[TV81] P. Tinsley and P. Von Zur Muehlen. A maximum probability approach to short-run policy. Journal of Econometrics, 15(1):31– 48, January 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901287.

Tsurumi:1986:GSM

[TWI86] Hiroki Tsurumi, Hajime Wago, and Pekka Ilmakunnas. Gradual switching multivariate regression models with stochastic crossequational constraints and an application to the klem translog production model. Journal of Econometrics, 31(3):235–253, April 1986. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407686900606.

Ullah:1984:SDS

[UCS84] A. Ullah, R. A. L. Carter, and V. K. Srivastava. The sampling distribution of shrinkage estimators and their F -ratios in the regression model. Journal of Econometrics, 25(1–2):109–122, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768490040X.

Ullah:1983:PSE

[USC83] A. Ullah, V. K. Srivastava, and R. Chandra. Properties of shrinkage estimators in linear regression when disturbances are not normal. Journal of Econometrics, 21(3):389–402, April 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900532.

Villasenor:1989:ELC

[VA89] José A. Villaseñor and Barry C. Arnold. Elliptical Lorenz curves. Journal of Econometrics, 40(2):327–338, February 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407689900894. See corrigendum [Kra13].

Vandaele:1981:REA

[Van81] Walter Vandaele. Robust estimation of ARIMA models. Journal of Econometrics, 16(1):163, May 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407681900920.

VanDijk:1985:EI

[Van85]  Herman K. Van Dijk.  Editor’s introduction.  Journal

of Econometrics, 29(1–2):1–2, July/August 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900296.

Varian:1985:NPA

[Var85] Hal R. Varian. Non-parametric analysis of optimizing behavior with measurement error. Journal of Econometrics, 30(1–2): 445–458, October/November 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407685901502.

vandenBroek:1980:EDS

[vdBFHM80] Julien van den Broek, Finn R. Førsund, Lennart Hjalmarsson, and Wim Meeusen. On the estimation of deterministic and stochastic frontier production functions: A comparison. Journal of Econometrics, 13(1):117–138, May 1980. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900469.

vanDijk:1980:FEB

[vDK80] H. K. van Dijk and T. Kloek. Further experience in Bayesian analysis using Monte Carlo integration. Journal of Econometrics, 14(3):307–328, December 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407680900305.

Vinod:1980:ISR

[Vin80] H. D. Vinod. Improved stein-rule estimator for regression problems. Journal of Econometrics, 12(2):143–150, February 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900020.

Vinod:1981:ISR

[Vin81] H. D. Vinod. Improved Stein-rule estimator for regression problems. Journal of Econometrics, 17(1):125, September 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900646.

Vinod:1982:MEM

[Vin82] H. D. Vinod. Maximum entropy measurement error estimates of singular covariance matrices in undersized samples. Journal of Econometrics, 20(2):163–174, November 1982. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407682900173.

VanPraag:1988:LSP

[VKD88] B. M. S. Van Praag, T. Kloek, and J. De Leeuw. Large-sample properties of method of moment estimators under different datagenerating processes. Journal of Econometrics, 37(1):157–169, January 1988. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407688900802.

VanDijk:1985:PMC

[VKGB85] Herman K. Van Dijk, Teun Kloek, C. Guus, and E. Boender. Posterior moments computed by mixed integration. Journal of Econometrics, 29(1–2):3–18, July/August 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685900302.

VanMontfort:1989:ERC

[VMD89] Kees Van Montfort, Ab Mooijaart, and Jan De Leeuw. Estimation of regression coefficients with the help of characteristic functions. Journal of Econometrics, 41(2):267–278, June 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900973.

Velu:1987:RRR

[VR87] Raja P. Velu and Gregory C. Reinsel. Reduced rank regression with autoregressive errors. Journal of Econometrics, 35(2–3): 317–335, July 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900315.

VandeVen:1981:DDP

[VV81] Wynand P. M. M. Van de Ven and Bernard M. S. Van Praag. The demand for deductibles in private health insurance: A probit model with sample selection. Journal of Econometrics, 17(2):229– 252, November 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681900282. See corrigendum

[Ano83c].

VanPraag:1989:EMA

[VW89] Bernard M. S. Van Praag and Bertram M. Wesselman. Elliptical multivariate analysis. Journal of Econometrics, 41(2):189– 203, June 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900936.

Veloce:1985:EED

[VZ85] William Veloce and Arnold Zellner. Entry and empirical demand and supply analysis for competitive industries. Journal of Econometrics, 30(1–2):459–471, October/November 1985. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407685901514.

Waldman:1982:SPS

[Wal82] Donald M. Waldman. A stationary point for the stochastic frontier likelihood. Journal of Econometrics, 18(2):275–279, February 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900410.

Waldman:1984:PTE

[Wal84] Donald M. Waldman. Properties of technical efficiency estimators in the stochastic frontier model. Journal of Econometrics, 25(3): 353–364, July 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440768490006X.

Waldman:1985:CDM

[Wal85] Donald M. Waldman. Computation in duration models with heterogeneity. Journal of Econometrics, 28(1):127–134, April 1985. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407685900715.

Wampler:1980:TPT

[Wam80] Roy H. Wampler. Test procedures and test problems for least squares algorithms. Journal of Econometrics, 12(1):3–22, January 1980. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407680900494.

Watson:1989:RSM

[Wat89] Mark W. Watson. Recursive solution methods for dynamic linear rational expectations models. Journal of Econometrics, 41(1): 65–89, May 1989. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407689900432.

Watson:1983:AAE

[WE83] Mark W. Watson and Robert F. Engle. Alternative algorithms for the estimation of dynamic factor, mimic and varying coefficient regression models. Journal of Econometrics, 23(3):385– 400, December 1983. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407683900660.

Webb:1985:BSP

[Web85] Robert I. Webb. The behavior of speculative prices and the consistency of economic models. Journal of Econometrics, 27(1):123– 130, January 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900478.

Weiss:1984:SST

[Wei84] Andrew A. Weiss. Systematic sampling and temporal aggregation in time series models. Journal of Econometrics, 26(3):271– 281, December 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900228.

West:1986:FVL

[Wes86] Kenneth D. West. Full-versus limited-information estimation of a rational-expectations model: Some numerical comparisons. Journal of Econometrics, 33(3):367–385, December 1986. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407686900035.

Wegge:1983:CE

[WF83a] Leon L. Wegge and Mark Feldman. Comment to the editor. Journal of Econometrics, 21(2):255–256, February 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900179.

Wegge:1983:ICM

[WF83b] Leon L. Wegge and Mark Feldman. Identifiability criteria for muth-rational expectations models. Journal of Econometrics, 21(2):245–254, February 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683900167.

White:1982:EI

[Whi82a] Halbert White. Editor’s introduction. Journal of Econometrics, 20(1):1–2, October 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407682901002.

White:1982:RCC

[Whi82b] Halbert White. Regularity conditions for Cox’s test of non-nested hypotheses. Journal of Econometrics, 19(2–3):301–318, August 1982. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407682900070.

White:1983:EI

[Whi83] Halbert White. Editor’s introduction. Journal of Econometrics, 21(1):1–3, January 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407683901161.

Wills:1987:NST

[Wil87] Hugh Wills. A note on specification tests for the multinomial logit model. Journal of Econometrics, 34(1–2):263–274, January/February 1987. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407687900753.

Wansbeek:1989:EEC

[WK89] Tom Wansbeek and Arie Kapteyn. Estimation of the errorcomponents model with incomplete panels. Journal of Econometrics, 41(3):341–361, July 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407689900663.

White:1981:CDE

[WO81] Halbert White and Lawrence Olson. Conditional distributions of earnings, wages and hours for blacks and whites. Journal of Econometrics, 17(3):263–285, December 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900026.

Wolak:1989:TIC

[Wol89] Frank A. Wolak. Testing inequality constraints in linear econometric models. Journal of Econometrics, 41(2):205–235, June 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900948.

Wales:1983:ECD

[WW83] T. J. Wales and A. D. Woodland. Estimation of consumer demand systems with binding non-negativity constraints. Journal of Econometrics, 21(3):263–285, April 1983. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900465.

Xekalaki:1984:LRY

[Xek84] Evdokia Xekalaki. Linear regression and the Yule distribution. Journal of Econometrics, 24(3):397–403, March 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900617.

Xie:1988:SWC

[Xie88] Wen Zhi Xie. A simple way of computing the inverse moments of a non-central chi-square random variable. Journal of Econometrics, 37(3):389–393, March 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407688900139. See corrigendum [Xie11].

Xie:2011:CSW

[Xie11] Wen Zhi Xie. Corrigendum to “A simple way of computing the inverse moments of a non-central chi-square random variable” [J. Econom. 37 (1988) 389–393]. Journal of Econometrics, 161(2): 338, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002241. See [Xie88].

Yancey:1984:SIE

[YJM84] T. A. Yancey, G. G. Judge, and S. Miyazaki. Some improved estimators in the case of possible heteroscedasticity. Journal of Econometrics, 25(1–2):133–150, May/June 1984. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407684900423.

Yahav:1981:IMA

[YN81] J. Yahav and G. Nathan. International meeting on analysis of sample survey data and sequential analysis. Journal of Econometrics, 16(3):399, August 1981. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407681900373.

Zaman:1981:EMC

[Zam81] Asad Zaman. Estimators without moments: The case of the reciprocal of a normal mean. Journal of Econometrics, 15(2):289– 298, February 1981. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407681901184.

Zaman:1984:AMS

[Zam84] A. Zaman. Avoiding model selection by the use of shrinkage techniques. Journal of Econometrics, 25(1–2):73–85, May/June 1984. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407684900381.

Zellner:1988:BSA

[ZBV88] Arnold Zellner, Luc Bauwens, and Herman K. Van Dijk. Bayesian specification analysis and estimation of simultaneous equation models using Monte Carlo methods. Journal of Econometrics, 38(1–2):39–72, May/June 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900267.

Zellner:1981:POR

[Zel81] Arnold Zellner. Posterior odds ratios for regression hypotheses: General considerations and some specific results. Journal of Econometrics, 16(1):151–152, May 1981. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0304407681900828.

Zellner:1988:BAE

[Zel88a] Arnold Zellner. Bayesian analysis in econometrics. Journal of Econometrics, 37(1):27–50, January 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900723.

Zellner:1988:CCL

[Zel88b] Arnold Zellner. Causality and causal laws in economics. Journal of Econometrics, 39(1–2):7–21, September/October 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407688900383.

Zellner:1988:CME

[ZH88] Arnold Zellner and Richard A. Highfield. Calculation of maximum entropy distributions and approximation of marginalposterior distributions. Journal of Econometrics, 37(2):195–209, February 1988. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407688900024.

Zellner:1989:FIG

[ZH89] Arnold Zellner and Chansik Hong. Forecasting international growth rates using Bayesian shrinkage and other procedures. Journal of Econometrics, 40(1):183–202, January 1989. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407689900365.

Zieschang:1983:NDC

[Zie83] Kimberly D. Zieschang. A note on the decomposition of cost efficiency into technical and allocative components. Journal of Econometrics, 23(3):401–405, December 1983. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407683900672.

Zellner:1985:BRD

[ZM85] Arnold Zellner and Brent R. Moulton. Bayesian regression diagnostics with applications to international consumption and income data. Journal of Econometrics, 29(1–2):187–211, July/ August 1985. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407685900399.

Zellner:1984:BAD

[ZR84] Arnold Zellner and Peter E. Rossi. Bayesian analysis of dichotomous quantal response models. Journal of Econometrics, 25(3): 365–393, July 1984. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407684900071.